A Complete Bibliography of Publications in Archive for History of Exact Sciences

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

20 June 2024 Version 1.60

Title word cross-reference

c [Gol77]. ^{e} [Rās78]. ^{èm} [Ras83]. ^{th} [Ari72, BS84, Bie85, Bie87, Car86, Dem82, Fra89, Hof65, Høy88, Jah87, JS94, Lüt95, Mon72, Pet61, Rās78, Ras83, Vig00, You76]. A_{7} [Hel22]. α [Hei68]. β [Hei68]. G [Duc11]. ≫ [BS77]. H [Dia94]. ∈ [Pal85]. ≪ [BS77]. n^{ième} [Rās78]. Φ [Shn17, BB20]. π [Jam88, Sme70, Twe91a]. πλασματικóν [Ace09]. R = 150 [MB03]. S_{6} [Hel22]. ∑_{n=0}^{∞}(−1)^{n}/(2n + 1) [Kri68]. ∑_{n=0}^{∞} 1/n^{2} [Kri68]. y^{2} − k = x^{3} [GL19].

-theorem [Dia94].

1 [Pou03]. 1050 [Bri01, HA98]. 1166 [Moz19]. 12 [DF20]. 1260s [Moz18b]. 1270s [Moz18b]. 12th [Gru17]. 14 [HY86]. 1400 [Gol03]. 149a7 [Ace00]. 149a7-c3 [Ace00]. 1570s [MS15b]. 1637-ca.1750 [Bos84]. 1685 [LD01]. 170 [BS22]. 170/624 [BS22]. 17th [SS05, Lam13]. 1800 [Ull84]. 1856 [Sch75].

1858 [Sch82a]. 1866 [Bie81a]. 1871 [Bie81b]. 1878 [Tum92]. 1884 [Bie81c]. 1887-1888 [BYS^{+}21a]. 1893 [Bur72]. 18th [Ham76, Zab88]. 1900 [SS82]. 1908 [Bur71]. 1912 [Sau08]. 1916 [HB86]. 1916/17 [Rei92]. 1919/20 [SS06]. 1922 [HY86, Rei92]. 1922/23 [Rei92]. 1930s [Cen10, Kra92]. 1950s [Cen10]. 1960s [Bro16]. 1989 [Spa91a]. 19th [Epp98]. 19th-Century [Epp98]. 1h [Gru17].

2 [Dob01, Nau11]. 20th [Cou04, Fer99, Hen97]. 23 [Rei92]. 248-day [Jon83]. 28 [Pou01]. 2nd [Scr61].

3 [Dob01]. 322 [BPS11]. 3405 [Ste00b]. 39 [Dob01].

501 [MP22]. 52 [Kau85]. 5th [Bec67].

6 [Nau14]. 624 [BS22]. 65156 [Ste20].

76829 [FHR21, Ste22].

8 [Ace03a, Mar12].

ādh [Gol77]. Āryabhata [vdW88a].

A-type [SdJ23]. A. [Bur71, Pet61, She86, She89b, She89a]. A.D. [Gol03]. abacus [Uli15]. Abbaci [Han11]. Abel [GS94]. Abraham [Sch68, FG23, Gol70, GC18, Sch68]. Abschnitt [Scr61]. Absolute [Tru79, Cog22]. abstract [Kop71]. Absurdum [Lov01]. acceleration [Kus89, Twe92, Wes86]. According [HA98]. account [Cha15]. Accuracy [Aok96, Aok92]. achievement [Høy88]. Achromatic [Sha79]. acoustic [Dar07]. acoustics [Ull84]. Acronychal [Swe99]. ACT [HA98, Bri01]. Action [Dia99, Sau05, AB22, Wis81]. active [Bra85, PB06]. Actives [PB06]. ad [Car20, Lov01, Moz19]. adaequare [Bre94a]. Adaptation [Shi03]. additions [Aok96, Tru92]. additive [Sha78]. Adiabatic [NP04, Pér09a, Pér09b, NP06]. Adolf [OS14, Osw17]. adoption [Gua15]. Adrain [Dut90]. Adriaan [Boc09]. Adrian [FS05]. adventures [De 12]. advocacy [Not17]. Aeginetan [Amb12]. aerodynamics [Fer14]. aether [Ros69]. affine [Rei92]. affiner [Rei92]. Affinitätstheorie [Car86]. affinity [Car86]. after [Arn83a, Sch75]. Again [Hah02]. Against [Dob99a]. age [Bul09, Han77, Tru60]. ages [Eva76, GC23]. Agnesi [Tru89, Tru92]. air [Aub17, Cha15, Mid69, Web65]. aircraft [Fer14]. aire [GP09]. Akustik [Ull84]. Al [Ras70, Ses77, Bre94b, BN84, Jen72, Kau85, Moz14, Rās74, AH15, BHM15, Kau85, Man98b, Man04, MB03, Yaz11a, Yaz11b, Yaz18, Bre94b, CG94, CG15, Mae98, Moz14, Moz19, Moz24b, Ras72, Rās74, Ras80, Sab82, Van06]. Al-Bīrunī [Yaz18]. al-Baghdadi [Jen72]. al-Battani [Mae98]. Al-Bitūjī [Man04]. al-Din [Moz14]. al-Fahhad [Moz19]. al-Haggāg [Bre94b, Bre94b]. Al-Hasib [Yaz18]. al-Haytham [Ras80, Sab82]. al-Kāshī [Van06]. al-Kammâd [CG94]. Al-Kammad [Man98b, CG15]. al-Karaji [Ras72]. Al-Kashi [AH15]. Al-Khwārizmī [Kau85, MB03]. al-Khwārizmīs [Kau85]. Al-Khwarizmi [Yaz11a, Yaz11b]. al-Maghribi [Moz14]. al-Muqtabis [CG94]. al-Tūsī [Rās74]. al-Tusi [Rās74]. al-Zîj [CG94]. al-Zarqalluh [Moz24b]. Albert [Bac89, HY86, Bac89, Blu22, FS05, HY86, Mal16, Rou80, Sch88]. Albrecht [Fie97]. Alexander [Cha62, GH00]. Alexandria [Mas15]. Alexandroff [Arb79]. Alexandrov [Arb81]. Alfonso [Mid73, Cha04]. Alfred [Lê17]. Algebra [Goe68, Kau85, Ste00a, Blå16, Fre77, Haw08b, Kop71, Lam86, Nab08, Oak09, Oak18, Rās74, Fis99, Gil91]. Algebraic [Abr82, Mit19b, Cou10, DG20, Fra89, Hof65, Hol09, Mit19a, Sch82a, Sch86, Shk72, vdW71, vdW72b]. algebraicist [Loh66]. Algebraiker [Loh66]. algebraische [Hof65, Sch82a]. Algebraischen [Sch86, vdW72b]. algebraischer [Hof65]. Algebras [FS07, Cic15, Haw82, Haw88, Par85, Fen98, FS05]. algèbre [Gil91, Rās74]. Algorithms [Bie04, Fri14, IM21, Sar12]. Alhazen [Sab82]. aliquot [Ras83]. aliquotes [Ras83]. Aliza [Con15, Con18]. alla [BF93, Mai90]. Almagest [Jon99, Rid78]. Almanac [Ste20]. Almanach [Man92]. Almanacs [GS08]. alternative [Ste22, Vit04]. Alternatives [Vit04]. Altertum [Fre91]. altitude [Not23]. am [HY86]. America [DS12, Par88]. amiables [Ras83]. Ampère [Vol89, Vol89]. Analogien [Bie83a]. Analogies [Bie83a]. analogy [Wis81]. Analyse [Del11, BS77, BS85, Gis83, Jaë10]. Analysis [GC17, Hüt00, Jah87, SS98, BS77, CEF17, Cha16, Con18, Cou10, Dar07, Dom23, Dug73, FT13, Fra89, Gis83, GG70, Gra84, HH79, Lau89, Mah68, Man75, Mat78, Moe74, Moz16b, Moz18a, RKH^{+}21, SS82, Smi86, dC16, Ace11a, Del11]. analyst [Tay87]. analytic [Bil82, Lor16, Spe67, Wim12]. Analytical [Dia99, BC12, Cou14, FP12, Ull89]. analytischen [Ull89]. Anastácio [Dom23, dO88]. Anatolia [YT18]. Anatolius [McC96]. Ancient [Bel88, Bel98, Duk02a, Duk05a, Duk07, Wri00, AA72, BC20, Gil74, Jon83, LS84, LS86, Sei84, Sme70, Liu05]. Andalò [RGM19]. Andalusian [CG94]. Andean [AA72]. Andersson [Lüt94]. André [vdW71]. Anfänge [SS82, Sza60]. angle [Mai90]. angled [LS84]. angles [Duk10]. angolo [Mai90]. Angular [Cap02]. Annäherung [Spa02]. anniversary [Tru96]. annual [Moz24b]. annular [Yaz11a, Yaz11b]. Anomalies [Bri07]. anomalous [Nau15]. Anomaly [Duk07, Bri09]. Anonymous [Gol03]. anschauliche [Row24]. antecedents [Rab74]. anthropic [Kra10]. Anthyphairetic [Fow81]. anticipations [Dea88, VT18a, VT18b]. Antikythera [CE14, EC19, IJ19a, IJ19b]. Antique [Hüt00]. antiquity [Ace11a, Fre91, Lin68, Mal10]. Antonio [SV24]. Antoon [Kox88, Kox13, Sch84]. Anu [Tum92]. Anwendung [Sch74a]. Any [Fra98]. Apercu [LM98, Bus19]. Apollonius [DF20, Cio23b, Hog86]. Apologia [Mai90]. appeal [Pye82]. appearance [Ses85]. appearances [Smi81, Smi82]. application [Nau10, Roq92, Sch74a]. applications [Dal82, Del22b, Arn83i]. appraisal [Kyr21]. Approach [Bor07, Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, Dah80, Fra85, Man75, Spa02, Spe67, Tup22, Zam10]. Approaches [Kno99, Dai21]. approaching [Roq11]. approche [Dah80]. Approximation [Man98a, Wri00, BN93, Fow91]. Approximative [Gal01]. Apsides [Gro98]. Arab [Vin05]. Arabic [BS07, Bre94b, De 12, Hog84, Hog86, Oak09]. arabischen [Bre94b]. Arabo [Vin05]. Arbeiten [Sch82a]. Arc [Mer02, Hen97, Raw82]. Arc-en-Ciel [Mer02]. archaelogy [Eis91]. Archimedean [Ehr06, Fie97, Key22, Kno78, Pal08, SFS08]. Archimedes [Ber76, Del70, Kno76, Kno78, Kno86, Mar93, Net99]. architecture [Ray20]. archive [Tru96, BG08]. Archytas [Mas16]. arcs [Sma11]. Area [Dav03, Gio21, Sei72]. Arénaire [Del70]. Argument [Pou03, CG19, Wat90]. Arguments [Bos84, Hof65, PS10]. Aristarchus [BS07, Car14, Cio23a]. Aristote [Mer02]. Aristotele [Mar80]. Aristotelicum [Tót66]. Aristotle [Kno81, Kno82, Mar80, Men07, Mer02, Mil17, Smi78, Vit02]. arithmetic [AM14, Cor13, FS07, Neu97, GGHPMR21]. Arithmetica [BC12, Bus98, Tho05a, Ace09, Bus98, Ste01, Tho05b]. Arithmetical [Mit19a, Mit19b]. Arithmetics [Fen98]. Arithmetik [Neu97]. arithmétique [Bou13, MS15a]. arithmetisation [Jah87]. Arithmetisierung [Jah87]. arithmetization [Mat78]. Arnold [Sch85]. arquimediana [Mar93]. Ars [Con15]. art [Lam94]. Artaxerxes [Ste20]. article [Kno82, Kyr21, Lam89, Wei80]. artifacts [BC20]. Artin [BS84, DS12, Kie71, vdW72a]. as-Samaw’al [Ras72, Wat78]. ascendancy [Dea92]. Aspect [Fer99]. Aspects [Ari72, Ber03, BN93, BS84, Car20, Dho86, Fer07, Lau89]. Associations [Duk02a]. Astonishing [Gea10]. astrolabii [RGM19]. Astrologer [Fie84]. astrologers [Too85]. astrological [Ste22]. Astronomentreffen [Her70]. astronomer [Her70]. astronomers [Too85]. Astronomiae [Dav03, Del16a]. Astronomical [Man92, Moz18b, Shi03, Ste00b, Cha91, CG15, FHR21, GC17, GC18, MZ24, Moz16b, NTP04b, NTP04a, She73b, YT18, ZH12, vdW84a, vdW84b, vdW84c, vdW85a, vdW85b, FHR21]. astronomique [NTP04b, NTP04a]. Astronomy [Cha04, Duk07, Man98a, Str81, Bel88, BG96, BB20, CG94, Duk08, GB91, Gra15, GS08, GS09, Jon83, Moz16b, Not17, Not18, Not23, She84a, She93b, vdW78b, vdW82, FHR21]. Asymmetry [HG05a]. asymptotic [Sch77]. atelier [AB19]. atmosphärischen [Sch74a]. atmosphere [Gol77, Sch73]. Atmospheric [Bru81, Sch74a]. atom [Hei68, MR10, Pet14]. Atomic [Hoy73, McC66, Ros71]. Atomtheorie [Hoy73]. Atronomice [Cha04]. attempt [Bie83b, Tum92]. attempts [Bie87, Lüt79, Cao91]. attitudes [GG85]. attorno [Mai90]. Attractions [Gro98]. attributed [McC96, Yaz18]. attribution [Zab88]. Auerbach [RC23]. Augenzeugen [Sch86]. Auguste [Pap17]. Auseinandersetzung [Hof65]. Ausgleichspunkt [vdW60]. author [Art85]. auto [GL12]. auto-entretenues [GL12]. automorphic [AG09]. Autor [Art85]. autrement [AB22]. Autres [Vit04]. aux [Ras83, Sma11]. avoided [Art85]. awakening [Del22b]. axiom [Big83, Sza60, Tru79, Bre84]. axiomatic [BN78]. axiomatically [Sei75]. axiomatics [De 16, Hol17]. axiomatisation [Jaë20]. axiomatization [Cor97a]. Axiomatizations [Lac00]. Axiomensystems [Sza60]. axioms [Seg94].

B [vdW72b, Bre94b, Vit02, BSS88]. B-splines [BSS88]. B. [vdW72b]. B.C. [Gul87]. Bīrunī [Yaz18]. Babylon [Ste06, SdJ23]. Babylonian [FHR21, BB20, BJ00, Bri07, Bri09, Bri10, Duk10, EC19, GS08, GS09, HA98, Jon04, Oss18, RSW04, Shn17, SM21, Ste24a, Swe99, dJ19a, dJ19b, dJH20, dJ21, vdW68, vdW80a, vdW80b]. Babylonians [Goe68]. Babylonier [Goe68]. Back [IJ19b, Roc21, IJ19a]. Background [She03, Bie85, Dau71, Edw75, Edw77, Frö94, Sha79, Wil82]. Bacon [Lin70a]. Bad [Ses77]. Baghdadi [Jen72]. Baire [Dug76, Mon72]. Baisiu [Man89]. Baker [AB12]. Bakr [Ses77]. Balance [Mar06, SC80, vdW60]. balistique [Eis96]. ballistic [Eis96]. Ballistics [Aub17]. Banach [Jaë20]. bar [FG23, Jao74]. Barbaro [Fie97]. Barbier [SPJ01]. barre [Jao74]. Barrow [Lüt94]. barycentric [Rec94]. based [BS22]. Basis [Duk05a, Bie85, Bie92, Gis83, Sch82b, Spa91b]. Battani [Mae98]. Baye [Dal82]. Bayes [Dal82, Dal86, Hal90]. be [WH14]. Bearbeitung [Kau85]. became [Car86, Eck24]. Becquerel [dAM97]. Bedeutung [Hof66, Sch82c, Hof60]. been [Fre77]. Before [Van21, Bik78, Gou18, Joh84, Pet11, Sch74b]. beginning [Ste24a]. beginnings [Arb79, Arb81, Kle61, SS82, Sza60]. Begriff [Pet63]. Begriffes [Ras71]. begriffliche [Car86]. Behavior [KC00]. beiden [Sch84]. Being [Pou04]. Beitrag [Sch75, Vol87]. Beiträge [Neu73a]. Beltrami [PT96, Taz93]. Bemerkungen [ENP82, Lau65]. Ben [Rab70, Cha91, GC17, Man98a]. Beobachtbarkeit [ST94]. Bericht [Sch86]. Bernard [Joh77]. Bernoulli [Sha78, Spe96]. Bertrand [SPJ01, She94a]. Bessel [Dut95]. Bestimmung [Kri68]. beta [Dut81, ED09, Hal90]. betrayed [Wei78]. betriebenen [Bie87]. bets [FGP81]. Betti [Bot77b, Bot77b, Wei79, Wei80]. Between [Boc09, Buc01, Dar02, Dia99, Duk02a, Ste01, AT14, BN88, Bot77a, Bre94b, Bur71, Bur72, Car18, DF12, Cor99, Gil75, Gou14, GS08, Lam87, LM98, Lüt95, Mai90, Mai94b, Mil83, Moe74, Mon72, Neu81, Neu97, Osw17, Pap17, Pye82, Scr63, Spa91b, Tan09, Twe91b, Ver20, Vie21, Voe08, vdW72b]. Beweis [SS86, Vol89]. Beyond [Bie04, FS07, Cor08, Del22a]. Beziehung [Bie82]. Beziehungen [Sch88]. Bianchini [Cha16, GC04, Van18, Van21]. Biao [Gua10]. big [Duc11]. biggest [Blu22]. Binocular [Gou18]. binomial [BHM15]. biography [Cou14, Fre74]. biology [BG08, She80]. biquadratic [Col77]. Birefringence [Buc07]. Birkhoff [Roq11]. Birth [SS05, Wei80, FGP81, Haw91, Wei79]. Births [Dar02]. bis [Bie92, HB86, Neu97, Rei73, vdW72a]. Bisecting [Oss18]. Bitūjī [Man04].

Biton [Key22]. Bitter [Gea10]. Björling [Brå07]. black [Eis91]. blocks [Eis87]. Blondel [GL12, Ray20]. blunder [Blu22]. BM [FHR21, Ste20, Ste22]. Bodies [KC00, BS84, Kje08, Wil87]. Bodleian [Kau85, Kau85]. Body [Dob98, McG66, Pou04, BG94, Pou92a, Ste22]. Bohr [Hoy73]. Bohrschen [Hoy73]. Bois [Fis81]. Boltzmann [Bie92, Hoy80, Bac90, Bie81a, Bie92, Hoy80, vP91]. Boltzmanns [Bie81a]. Bolyai [Tan09]. Bolzano [Spa91b, Fre71, GG70, Joh77, Lau65, Spa91a, Spa91b, vR63]. Bolzanos [Lau65, Spa91a]. Bombelli [RS98, Wag10]. bones [BC20]. Bonfils [GC17]. Bonjorn [Cha91]. Book [Dau08a, Dau08b, Mar12, Nau11, Abo08, Jao74, Ace03a, Bec67, Ber76, Cor13, Dob01, FHR21, Fow80, Fow82, Pou09, Rus98, Sei75]. books [DF20, GGHPMR21, Neu73b, Neu74]. boostered [RKH^{+}21]. Borel [BN78, Mon72]. Borelli [DF20, Mid73]. Born [Im96]. Boscovich [Pro88, She73c, Spe67]. Boscovich’s [DF18]. Boyle [Cha15, Web65]. Bradwardine [Bus98, Bus98, Gra65, Mol78]. Brahe [Car20, Mae98, Ros81]. Brahmagupta [Pot74]. Brahman [vdW88a]. braids [Fri20]. Breaking [Pal08]. Brechungsgesetz [Hen01]. Brevity [Bra03]. Brianchon [Del22a, Del22b]. Briefbüchern [Mes65]. Briefwechsel [Bur71, Bur72, Sch84, vdW72b]. Britton [Jon10, Shn17]. Broglie [Kub70, Kub70]. Brook [Fei85]. brothers [OS14]. Brouillon [AB17, AB22]. Brouncker [Dut82]. Brown [Bru68]. Brownian [Bru68, Yea11]. Brudzewo [Mal16]. bruslans [AB22]. Buch [Bec67]. Bücher [Neu73b, Neu74]. Buffon [Lov01, Zab88]. Building [Ray14]. Bulkeley [Gou14]. Buniakovsky [She91b]. Bunte [ENP82]. Buono [Mid69]. Buoyancy [Pal05]. Burbury [Dia94]. Bürgi [WH19]. burning [Ace11a]. Burnside [AG09, Ald09]. Bury [Dav86, Dav90]. Byzantine [Vin05]. Byzantins [Vin05].

C [Gru17]. C. [BF93, Dav90, Gar89b, Gar91, Mai82, She79, She94b]. c3 [Ace00]. ca.1750 [Bos84]. Cady [Kat08]. Caerleon [MZ24]. Caianello [CD13]. Calabria [Mai94b]. Calcolo [FGP81, RS98]. calcul [Gil91, Jao74]. Calculating [Bro05, WH19]. calculation [Car09, Car17, Jao74, Pep68, RS98, ZH24, vdW60]. calculations [Sch74a]. Calculi [Car20]. Calculus [Fra94, Bos74, CEF17, Cog22, Dom23, FGP81, Fra85, Fra89, Gil91, GXB21, Kop71, Lüt79, Nau10, Pet87, RA20, Rei89, Roc21]. Calendar [Bie07, GS09, vdW84b, vdW84c, Yaz20]. calendars [vdW84a, vdW84b, vdW84c, vdW85a, vdW85b]. calidum [Bus65b]. called [Wil70]. Callippos [vdW84b]. caloric [Bie81b, Bie81c, Bie82]. Cambridge [Not18]. came [Dea92]. camera [Thr94]. Campbell [AB12, Mil83]. can [Fra98]. Canonical [FN23]. Cantor [Dau71, Mes65, Spa91b]. Cantors [Mes65]. capabilities [Gea85]. Capella [Car17]. Capillarity [Bik78]. carbon [Kra10]. carbon-12 [Kra10]. Cardano [Con15, Con18]. Carl [DF12, Net20]. Carnot [Del22b, DPC95, Hoy74, Hoy76, Ler72, Tru79].

Carnotschen [Hoy74, Hoy76]. Carré [Vin05]. carrés [Ses98]. Cartan [Cog11, Haw05, Kat85]. Cartesian [Hal60, Mac23]. cartography [Pap17, Raw82]. Case [Dal01, Fis99, Net99, Gou14, Smi82, ZH17]. case-study [Smi82]. cases [AT14]. Casimir [CG86]. Casorati [Neu78, Bot77a, Bot77b, Neu78, Ros15]. Castelnuovo [Gar91]. casts [Nab08]. casus [Con15]. Catacaustics [SS05]. catalog [Mar21]. Catalogues [Duk02a]. catalysis [Lai86]. Cauchy [Spa02, Bou13, Brå07, Dah80, Fre71, GG70, Jah87, Rei03, Smi86, Spa02]. Cauchys [Jah87, Rei03, Spa02, Spa02]. Cavalieri [And85, Pal09]. celestial [Gra61]. Center [Gro98]. centers [BS22]. central [BSS88]. Centrifugal [Mel06]. centrum [Rec94]. centuries [Dem82, Mon72, Not24, Rās78, Ras83]. Century [Bie85, Car86, Cha04, Cou04, Dar03, Epp98, Fer99, Pet61, SS05, Wei00, Ari72, Bel91, Bie87, BN84, Dea88, Dos75, Eva76, Fra89, GG70, Gul87, Ham76, Hen97, Hof65, Hol17, Høy88, Jah87, JS94, Lin68, Lin70b, Lüt95, MZ24, Not18, Sha73, Sil79, Tay74, Ung74, Vig00, Web65, Whi61, You76, Zab88, vB17, Lam13, SPJ01]. Certainty [Lov01]. Cesàro [PT96]. Ceva [BN84]. challenges [Uli15]. chamber [FT13]. Chandrasekhar [Dob99a]. change [Cra13]. changed [Rit15]. changing [Duc11, Fra85]. chapter [Cou10, Gra84]. chapters [Lam94]. character [AM14]. characters [Haw71, Haw74]. charge [Gal17]. Charles [Dav86]. Chasles [Bus19]. Chebyshev [She94c]. Chemical [KL84, Lai85, Lai88, Zam10, Dai21]. chemistry [Dav86, Lai85]. chez [Ras70]. ch’i [Lam79]. ch’i-meng [Lam79]. Chicago [vD09]. Chieh [Yon69, Lam79]. China [CM15, Che15, LS84, LS86, LD01, LM98, Liu05]. Chine [LD01, LM98]. Chinese [BC20, Jam88, Lam87, Lam88, Lam89, She88, Shi03]. chinoise [Jam88]. Chladni [Bel91, Ull84]. choice [dAM97]. choices [Mal10]. chords [vdW88c]. Chr [Hof66]. Chr. [Hof66]. Christian [Kra92]. Christophorum [Mai90]. chronological [Kno78]. chu [Gua15, Lam79]. Ciel [Mer02]. circa [Mai90]. circle [Del22a, DS12, Hof66, Kno76, Kno86, Men23, Pal08, Sei72, Sei81]. circles [Kno92, Lor95, Wes64]. circular [Arn83f]. circulation [SS82]. city [RC23]. city-size [RC23]. Civita [CT24, IR16, Rei92]. Civitasche [Rei92]. Clairaut [Bik78, Del11]. Class [FS07, Pal85]. classe [Pal85]. classic [Lam97]. Classical [KC00, Ron60, Bro16, CRT10, FT13, Hoy76, She71b, Smi81, Smi82]. Classification [Jae98]. clause [Bie81a, Bie81b, Bie82, Bie83b, Bie87]. Clausiu [Sch75, Bie81b, Sch74a]. Clausius [Bie81b, Bie83a, Sch74a, Sch75]. Clausiussche [Bie83a]. Clavasio [Bus65a]. Clavio [Mai82, Mai90]. Clavium [Mai90]. Clavius [Bie07, Lit22, Mai90, Pal01]. Clebsch [Lê17]. Clifford [Cog15]. clos [BS84]. closed [BS84]. closing [Arn83e]. Clothing [Lê17]. cloud [FT13, Man92]. codicum [Ray14]. coefficients [BHM15, CT24]. Cognitive [Bor07, Pal05]. coin [Amb12]. coinage [Big19]. coincidences [Tak80]. cold [Mal91a]. Colin [Mil83, Mil85, Twe91b]. Collaboration [FS05, DS12, Ill89]. Collection [Ber03, Cio22]. College [Lai88]. collision [Im96]. collisions [Kra92]. color [PR23, RP23]. colors [Loh61]. colour [Bec73]. coloured [Wes64]. colouring [Big83]. column [BB20, Shn17]. combined [ZH17]. Comets [Ruf00]. coming [Han77, WH14]. coming-to-be [WH14]. Commandino [Cio22, Cio23a, Cio23b]. commensurability [Gra61]. Commentaire [NB84, NB89]. Commentariolum [Mal16]. Commentariolus [CFH^{+}70, CFH^{+}70]. Commentary [Dau08a, Dau08b, Wei00, NB89]. comments [Ewa69]. Commins [Fra98]. Commitment [Top71]. Common [Mit19b, Mit19a, Tup22]. common-sense [Tup22]. Communication [Sau05, Fra98]. communicator [Lüt94]. comparison [GS08]. compatibility [Bro81]. competition [BG94]. compiled [GC17]. Complementarity [Pet11, Sar12]. Complete [Ace00]. completeness [ST94]. Complex [OS14, Man75, Mil83]. compositione [RGM19]. compounds [dAM97]. Computation [DEHY76, Oss18]. computational [Nau94]. Computations [Dau08a, Dau08b]. Computed [HA98]. Computer [Liu05, Bul09]. computers [Cor08]. concentric [Moz22]. Concept [Har03, Neu06, AM14, Bie83a, BG96, Bra85, Cha15, Cho10, Dah80, Dho86, Frö94, Kje08, Mon72, Sch17, Sid18a, Spa91b, Van92, Vie21, Web65, You76, dO88, vB76]. Conception [Kje01, Ari72, Smi86]. conceptions [Lam13]. concepts [Bie85, Del22c, Nem62, vB17]. Conceptual [Dia99, BS84, Car86, DPC95, Lam86, Sch71, Tan09]. conceptualistic [Kyr21]. conceptuelles [BS84]. concerning [RD94, Sal74]. Condition [Dia94]. conditions [Hoy73]. configuration [dJH20]. conic [DF18, Del20, Sma11, Twe91b]. Conical [Hüt00, BS22]. conics [Del16a, DF20, Cio23b]. coniques [Sma11]. conjectural [Del20]. conjecture [Abo08, JW22]. Conjunction [Moz19]. connected [Dho86]. connection [Hoy74, Rei92]. connexes [Kay23]. consequence [Tru79]. consequences [Bac89, Vol89]. consequentia [Ace19]. conservation [Cra13, Sar12]. Considerations [Mai90]. Considerazioni [Mai90]. consistent [Hei70]. Constant [Ear01, Kra03, Vol89]. constraint [Mac23]. constraints [Im96]. constructableness [Pet63]. Construction [Ses03, SS09, Bos84, Che15, Sid18b]. Constructions [Har03, Hog84, Shk72, vdW78a]. contact [Cou10, Kay23, Mai90, Hon94]. Contactu [Mai90]. Contained [Rus98, BS22, Mai94b, VT18a, VT18b]. contatto [Mai90]. contemporary [DG20]. content [Lit22, NTP04b, NTP04a]. contents [Hof66]. contenu [NTP04b, NTP04a]. contenuti [Mai94b]. context [Høy88, RKH^{+}21, Sch74a, Seg94]. Contexts [Kje01]. contiguous [Wes64]. Continuation [Lac00]. Continued [Haw08a, Dut82, OS14]. Continuity [Cra13, BN88, Duc11]. Continuum [Bor07, CR07, Spa02, Rit15]. contraction [War91]. contribution [CR07, RW12, Sch75, She89b, Vol87, Yaz18]. contributions [Arn84, Ayo80, Ayo84, Dav86, Fra83, Row24, Neu73a]. controverse [Ver20]. Controversy [Bie07, Hon94, SS06, Gar89b, Kus89, Mai90, Mal91a, Mil83, Ren97, Ver20, dC16, vB17]. Conventions [Han11]. convergence [Twe92, Vie21]. convergent [dO88]. conversion [Don91, Van21]. convex [Kje08]. Cooling [Ruf63]. Cooperative [Nis05]. Coordinate [Duk02b]. coordinates [Mac23, Mar12, Van18, Van21]. Copernican [Car18, Gea85]. Copernicanae [Dav03]. Copernicu [CFH^{+}70]. Copernicus [Car18, CFH^{+}70, GC04, Mae98, Mal16, Moe74]. Córdoba [Cha04]. Corporum [Kal77]. corps [BS84]. Corpus [Tót66, Kno78, Tót66]. Corpuscularity [NP04]. Correction [BYS^{+}21a, Ehb23, IJ19b, Lam89, Mit19a, VT18a, Kno82]. Corrections [Aok96, She83, Tru92]. Correspondence [Pet14, Bot77a, DF12, Gar91, PT96, Ruf13, Twe91b, DJ07a, DJ07b]. Corrigenda [Car20]. corrispondenza [Bot77a, Gar91, PT96]. cosmologie [Ker86]. cosmology [Ker86, PR11]. could [Hol17]. Coulomb [Mar06]. Countable [Ull00]. Counterexamples [Pou01]. counting [Sei62]. counts [Fra98]. courbes [Del11]. Court [Cha04]. courtly [Fra13]. Cousin [Cho10]. Covariant [Nor04, Nor92]. Covering [MT13]. Cracking [BC20]. Creation [Hes00, Haw72, Haw74, Shn17, Smi77, Zan14]. Creative [Fis99]. crescent [Yaz18]. criteria [Bec67]. critical [Nis11, RGM19]. Criticism [Not17, Net20]. critique [For69]. Crofton [SPJ01]. cross [AB19]. crossing [Fri20]. crunching [Cor08]. crystallization [Arn83g, Sch74b]. crystals [Ewa69, For69, Kat08]. cube [Gui16, Mas16, RS98]. Cubic [Net99, Col77]. Cubiche [RS98]. culmination [McL11]. cultures [Asc88]. Cunha [Dom23, dO88]. current [Bro67]. currents [RD94]. curvature [Ger82]. curve [Gug77]. curves [Bos81, Del11, Hof65, She95a]. Cyclotomie [Bou13].

D [She78]. D. [BS84]. d’Abraham [BS85]. d’Abu [Ses77]. d’Albert [Rou80]. D’Alembert [Wil87]. dall’enunciato [BF93]. Damianos [Eas70]. Damis [Key22]. d’analyse [Dug73]. Daniele [Fie97]. Dans [Vit04, AB17, AB19, AB22, BS77, Bou13, Del70, Lam13, MS15a, Nab08, Ses77, Vit02]. daraus [Car86]. Darboux [Del11]. d’Archimède [Bac63, Del70]. d’Aristote [Vit02]. dark [Kra12]. d’Artin [BS85]. data [Fra98, GS08]. date [Tum92, Yaz20, vdW68]. dated [GB91]. dating [Mas15, Ste20]. d’Augustin [Bou13]. d’Augustin-Louis [Bou13]. David [Cha91, Cor97a, Cor99, Gio21, Toe86]. day [Jon83, vdW88a]. death [Fis66]. debate [Che97]. Debeaune [Scr61]. Debeauneschen [Scr61]. Débuts [Arb79, Arb81]. decay [Bra85]. December [HY86, Tum92]. decimal [Rās78]. Décimales [Rās78]. decline [Bos84]. Dedekind [VT18a, ENP82, Sch82a]. Dedekinds [Sch82a, VT18b, ENP82]. Dedò [Fri20]. Deducing [Dar20]. deduction [Neu78]. deductive [Ace11b]. Defence [vdW76, Blå16]. defense [Ros81]. deferents [Gru17]. Defining [Tho05b, Tho05a]. Definite [Lau89]. Definition [Fer99, HG05b, She98, GP09, Roq11, GP09]. Definitions [Men07, Rus98]. deformations [Cic15]. Dehn [Gug77]. Del [Gar89b, CD13, FGP81, Mid69]. Delambre [Wil80a, Wil80b]. Delicate [FS05]. della [Fie97]. dell’Equazione [CG80]. demise [Nor92]. demonstrated [BF93]. demonstrations [Vit04, Vit04]. Dénes [Fra97]. Denis [Arn84, Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e]. Densities [Coh98]. Density [She95a, Duc11, Zan13, Zan14]. deren [Art85]. derivation [Don21]. derivations [Eas70, Mil85]. derivative [Bos74]. Desargues [AB17, AB19, AB22, Bri21, Del22c]. Descarte [Sch71]. Descartes [Scr61, BF93, Boc09, Bos81, Gal80, Sak82, Sch71, Scr61]. descriptions [Nau14]. descriptive [Coo93]. design [Amb12]. Detector [Hes00]. Determination [Coh98, Kri68, Mae98, Pro88, Swe80]. Determining [Buc07]. detonations [RKH^{+}21]. dÉuclide [Vit04, RDV01]. develop [Sei75]. developing [Bie85]. Development [Kje02, Pal05, Ari72, Ber66, Bil82, Bol72, Bru73, Bru74, BY16, Car86, CGM87, Dai21, De 16, Dea81, Dea82, GG85, Ger82, HB86, Hol09, Hoy73, JP13, Jon83, Kie71, Kje08, Lai86, Loh68, Mit65, Nem62, Neu97, PR23, Pet87, Sch77, Sch82c, She88, Tuc73, Ull84, Vie21, War91]. Developments [SPJ01, Tay74]. Dezember [HY86]. d’Hipparque [NB84, NB89]. diagrammatic [BHM15]. diagrams [Ray14]. dialogue [Scr63]. diameter [AB19]. diameters [Fow82]. diamètre [AB19]. diaries [Mes65]. Dickson [Fen98]. Did [Fre71, Sei75, Gil74, Raw82]. difference [Car18, Tan09]. differences [Ros15]. Different [Kje01, Kje02, Kno99, BS84, BPS11]. differentiability [Vol89]. differentiable [Bil82]. Differential [Kat85, Cog22, Dem82, Dem83, IM21, Lüt84b, Lüt95, McH71, RA20, Rei73, Rei89, Sch77, Tay74]. Differentialgeometrie [Rei73, Rei89]. Differentials [Bos74]. différentielles [Bac63]. différents [BS84]. Differenzierbarkeit [Vol89]. diffraction [BY16, Ewa69, For69]. diffusion [Zam10]. digamma [Fer07]. dilemma [Bel91]. dimension [Joh77, Joh79, Joh81, Kno86]. dimensional [dC16]. Dimensions [Arb81, dC16]. dimostrazione [BF93]. Din [Moz14, Rās74]. d’intégration [Sma11]. d’involution [AB17]. Dionysios [vdW84c]. Diophantine [GL19]. Diophantos [Tho05a, Tho05b]. Diophantus [Ace09, BC12, GGHPMR21]. Dioptrik [Loh68]. Dirac [Ehb23, Ehb22, Kra81, MR10]. direction [RD94]. Dirichlet [AM14, Osw17]. Disagreement [Buc01]. Disappear [HG05a]. Disappointment [Gea10]. Discontinuity [NP04]. Discovered [Bel98, BS22, BYS^{+}21a, BYS^{+}21b, Dal86]. discoveries [Del16b, Del16c]. Discovery [Cap02, Kat03, Sau99, Ste01, And94, Bel85, Bra85, Bur67, DPC95, Dut96a, Ewa69, For69, Hen97, Hen01, Kat08, Kub70, Moz24b, Sch74b, Was71, Wat72, Web65]. Discrepant [Buc06]. Discussion [Sch71]. Discussions [KC00, Mon72]. disgregation [Bie83a]. disgregations [Bie85]. Disgregationskonzept [Bie83a]. Disgregationskonzepte [Bie85]. Diskussion [Sch71]. dispersion [Bec73, Sha79]. Displaced [CG13, Hüt00]. displacement [Bro67]. disposition [Rid80]. Disquiet [Arn83d]. Distance [Bro05, Vig00, Car09, Wis81]. Distances [BS07, Cio23a, Car14, Too74]. distant [Sal74]. distribution [Bac90]. distributions [HB86, RC23]. Divergent [Fer99, Dut96b, Tuc73]. diversity [Sch82b]. Division [FS05, Fri14]. divisions [Abo08]. divorced [Eck24]. dix [BS84, Lam13, Vig00]. dix-huitième [Vig00]. dix-septième [BS84]. do [CM15]. Doctrine [Ras70]. documented [Dar93]. documents [Dug76, Hen01, Loh66]. Doetsch [Dea82]. Doing [Fra04]. Dokumente [Hen01, Loh66]. Dominicus [Bus65a]. don’t [Kno92]. double [Buc80a, CG19, Går89a]. doubling [Mas16]. Dragging [Ped00]. Dragon [MP22]. Dragon-Head [MP22]. dream [Car86]. Drowning [Ace03a]. Drude [Ren97]. dualism [Bac89]. Dualismus [Bac89]. Duan [LA84]. Duhem [Dai21]. Duillier [Zeh83]. Duke [CD13]. d’une [GP09, Jao74, Ker86]. duplication [Sei66]. duration [Hal88]. durch [Scr61, Was71]. Dürer [Fie97]. during [Che97, Not24, SFS08, SS82]. Durnheimer [Bus98]. dynamical [Dea88, RW12]. Dynamics [Dob98, Bel91, Cou14, Gio24, Hal60, LP15, Mac23, Nau94, Nem62]. Dynamique [Mil73]. Dynkin [AB12].

E. [BS84, Bot77b, Brå07, Bur71, Bur72, Fra98, SPJ01]. earlier [Eva76]. earliest [CM15, Raw82]. Early [Buc06, Cha04, Dos75, FS05, FS07, KC00, Lac00, Rab74, Sen98, She77a, Tru75, AB12, Bil82, CM15, Dal82, Dar93, Dut84, Dut91, Dut95, Fer96, Fra83, Gal17, GG70, Hen97, Hof65, Hol09, Hol17, Kal77, Kau85, Kra92, Kra23, Nau94, Nau10, Neu73a, Nor92, OS14, PR11, Rec94, Roq92, Row24, Sau07, Sch82b, Scr61, She93b, Shk72, SdJ23, Sza64, Tay82, Top71, Twe92, Ver13a, Ver13b, Wat80, Wim12, dC16, vB17, vdW78a, vdW82]. earth [Duc11, Ber91, Bro05, Car09, Coh98, Dut93, Gre96]. Easter [Bie04]. Easy [Bra03]. eccentricity [Car21]. eccentrics [Gea85]. échanges [LM98]. échelles [Sma11]. eclectic [Lit22]. Eclipse [Gua15, Liu05, Ste00c, CE14, IJ19a, IJ19b, Yaz11a, Yaz11b]. eclipses [Moz24a]. École [Arb79]. edition [DF20, Cio22, Cio23a, Cio23b, RGM19]. editor [Tru96, Wei78]. Editorial [BB97]. Editors [BG09]. education [Zan14]. Effect [Kat03, Hen97, Hof66, Knu76, Str92]. effects [Neu81]. efflux [Eck24]. Egidius [Man89]. Egyptian [Abr84b, Gil74]. Egyptians [vdW85a]. Ehrenfest [Jae98, NP04, NP06, Pér09a, Pér09b]. Eigenartige [Loh68]. eighteenth [Bel91, vB17]. eighteenth-century [Bel91, vB17]. Eindringen [Rei89]. Einfluß [Bot77b, Loh68]. Einführung [Sch75]. Einleitung [Ull89]. Einstein [Bac89, Sch88, Bac89, BN88, Blu22, CT24, EG78, HY86, Hen94, HG05a, Ill89, Mal91b, Nav98, Nor92, Nor04, PS10, Pye77, Ren97, Sal74, Sau05, Sau07, SS21, Sch88, ST94, vD04, vD09]. Einsteins [HY86, ST94]. Elasticity [Dar02, CRT10, Mid69, SV24, Taz93, Web65, Arn83g]. Electric [Gal17, BYS^{+}21a, BYS^{+}21b, Hen97]. electricity [Hom67, Wis81]. electrodynamic [Dar93]. Electrodynamics [Nor04, Bro16, Roq92]. Electromagnetic [Cor99]. electron [CG86, Gol70, Pye79, Mil73]. electron-theory [Pye79]. electronic [Dav90, War91, Yea11]. Electrons [vD04, Kra81]. element [PR23]. Elementa [Car20]. elementary [Kra23]. Elemente [Art88, Neu73b, Neu74, Art85]. Elements [Art88, Bri07, Dug73, De 12, Dug73, Fow80, Fow82, Har03, Kno78, Neu73b, Neu74, Rus98, Sei75, Vit04, Jaë10, RDV01, Vit04, Ace03a, Art85, Cor13, RDV01, Sid18b]. Eliakim [SS98]. elimination [Kno74]. Eliminations [Kno74]. Eliminations- [Kno74]. Elisabeth [Cha04]. ellipses [Gea85]. elliptic [Ayo84, Cog14, Cog15]. Embedding [Hol14]. Emergence [Ehr06, FT13, AG09, Arn84, Cho10, GZ79, Im96, Lüt84b, Man75, Pal08, Rod15, Sch86, Vol87]. Emil [Sch84, Sch85, vdW72a, DS12, Sch82c, Sch84, Sch85, vdW72a]. Émile [Ver20]. EMLR [Gil75]. Emmy [McL11]. Emphasis [Ace03a]. Empire [Sen98]. Empirical [Bri07, Wil70]. end [Van18, Van21]. Energy [Hes00, Kra12]. engagement [Ald09]. Engine [Key92]. Engineer [Dar03]. engineers [Yea11]. England [Big19, MZ24]. English [Dau08a, Dau08b, RGM19]. enigma [BC20, Yaz20]. Enjeux [Mer02]. Enlil [Tum92]. Enoch [FHR21]. entasis [Ray20]. Entdeckung [Bur67, Kub70, Mit65, Sch74b, Was71]. Entdeckungspfades [Hen01]. Entities [Rus98]. entre [LM98, Ver20, Voe08]. entretenues [GL12]. Entropie [Bie85]. Entropie- [Bie85]. entropy [Bie85, DPC95]. Entstehung [Sch86, Vol87]. Entstehungshintergrund [Bie85]. Entwicklung [Bol72, Car86, Ger82, HB86, Hoy73, Loh68, Neu97, Sch82c, Ull84]. Enuma [Tum92]. enunciated [BF93]. Envelopes [SS05]. envoyés [LD01]. Epicycles [Gea85]. epicyclic [Abr84a]. episode [HH79, Not18]. epistemological [Sch17]. Epistemology [Dal01]. Epitome [Dav03]. epoch [CE14]. épreuve [Eis96]. Equality [Gra65]. Equant [Duk05a]. Equation [Dar02, Net99, Ber66, CG80, GL19, GZKB^{+}12, Lê17, MZ24, Moz24b, Zam10, Dut97]. Equations [Gal01, Sau05, AT14, Bos84, CPFA07, Dem82, Dem83, Dho86, Fri14, GS94, Gil91, Ham76, IM21, Lüt84b, MR10, McH71, Oak09, Rās74, Sch77, Ses77, Ses77, Dho86, Gil91, Rās74]. equilibrium [Ber76, Lüt84a]. Equimultiples [Pal01]. Equinoxes [Dob98, Dob99a, Wil87]. equivalence [Nau14]. Equivalent [Pou04, Hoy78, Shk72, Sme70]. Era [Buc06]. Eratosthenes [Dut93, Gul87, Raw82]. erfolgreich [Bie87]. Erforschung [Sch73]. ergänzender [Hen01]. ergodic [vP91]. Ernesto [PT96]. Ernst [Bel91]. Errata [Ano81, She89b]. Erratum [Dau08a, NTP04a, Pér09b, Tho05a, Yaz11b, NTP04a]. Error [Hüt00, BG94, Rab74, Van18, Van21]. Errors [Dob99b, Pro88, She77b, She79, She95b, She95a]. erste [Sch74a]. ersten [Jah87, Neu73b]. eruption [Nab08]. Erwin [Han77, Hen94]. Esbozo [Mar93]. Essay [Wri00, Høy88, She73a, She73b]. Essays [Loh79]. est [Bus65b]. estate [Osw17]. estimation [Mag22]. Ether [Ped00, Taz93, Buc80b, Kox88]. Etherealization [Mit19b, Mit19a]. ethnomathematics [Asc88]. étiage [Eis86]. Étoiles [Vig00, NB89]. Étude [Mer02, Dah80, Jao74]. Euclid [Abo08, Ace03a, Ace08, Bec67, Bre94b, Bro81, Cor13, De 12, Fow80, Fow82, Har03, Kno92, Neu73b, Neu74, Pal01, RDV01, Rus98, Sei75, Sid18b, Vit04, Wei78]. Euclidean [Eva76, Amb12, Art85, Art88, BGTZ14, De 16, Fre91, Sza60, Tan09, Tho92]. Euclideo [Mai82]. Euctemon [vdW84a]. Eudoxan [Fow80, Fow81, Fow82, Men07, Rid79]. Eudoxos [Art88]. Eudoxus [Mue70, Sch23, Yav98, Yav01]. Eugene [Fen98, PT96]. Eugenio [PT96]. Euklid [Bre94b]. euklidischen [Sza60]. Euklids [Bec67, Neu73b, Neu74]. Euler [Bel10, Bis22, Cal76, CEF17, Dea81, Dea85, Dia99, Dut96b, Eck02, ED09, Fra94, Hof60, Mal00, MT97, Mil85, She72, Ver13a, Ver13b, Wil87]. Eulers [Hof60]. Europe [Not24]. Evaluations [Hal90]. evidence [GC18, Sha79]. Evolution [Cen10, Jae98, Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, Van92, Voe08, Wim12, Dho86, Voe08]. Exact [BG08, Mac23, Sch82c]. exactly [Dia94]. exakten [Sch82c]. examination [Dal82, Mol78]. examining [Pal98]. exchange [GCWL23, Sch84]. exchanges [LM98]. exercises [Scr61]. Existence [Har03, RA20]. existing [Yon69]. expanded [LA84]. expands [BG08]. expansions [Mal93]. Experiment [Hes00, Mar06, Fer14, Gol70, Gou14, Raw82, Seg80, vD09]. Experimental [Buc80a, Buc06, Fra98, Im96, Duc11, Sak82, Sha79, Ull84]. experimentellen [Ull84]. Experiments [Hah02, Bel91]. Explaining [Bro16]. explanation [Kno74]. explicating [Che15]. Explikationstheorie [Kno74]. Exploring [Lam87]. expositions [Dar93]. expression [Cal76]. extant [Kno86]. extends [DS12]. extinction [Bru81]. Extract [Wei78]. extracting [Wat78]. extraction [AH15, Rās78]. Extreme [Mae06, Sch71]. Extremwertmethode [Sch71]. eyewitness [Sch86].

F [BF93, Bot77b, Bur72, She79, She94b]. F. [Bot77b, Bur72]. Fa [Yon69]. factorial [Dut91, Fer07]. factors [Wil70]. Fagnano [Ayo84]. Fahhad [Moz19]. failure [Bie85, Høy88]. Fantappiè [Cic15]. Faraday [Knu76, RD94, Spe67]. Fassung [Hen01]. Fast [HJB85]. Fater [Mai94b]. Fatio [Zeh83]. Federico [Cio22, Cio23a, Cio23b]. Fedorow [Bur71, Bur72, Bur71, Bur72]. Fehlinterpretation [Bac89]. Felice [Ros15]. Felix [Hel22, Kay23, Row24]. Fermat [Bre94a, Del08, Edw75, Hof60, Sch71, Str68]. Fermats [Hof60]. Fermatschen [Sch71]. Ferromagnetic [Nis05]. Fibonacci [Han11, MS15a, RS98]. Fiction [EQ21]. fictions [EQ21]. Field [FS07, Cao91, CT24, Ehb22, Ehb23, Rei03, Sal74]. fifteenth [MZ24, Not18]. fifteenth-century [MZ24, Not18]. fifth [Tru96, Ace03a]. figured [Ras83]. figures [Abo08, Lüt84a, Ras83]. fi’l [Ses77]. fi’l-Hisab [Ses77]. filiations [BS84]. Filled [Ped00]. Filosofia [Mai94a]. final [Cic15]. Fine [Kra03, DG20]. Fine-Structure [Kra03]. Finite [She73a, Ros15, Sil79]. Finlay [Hen94]. Finsler [vdW72b, vdW72b]. First [Cou04, Dar02, Fer99, Bis22, Car18, Cho11, Dem82, Jah87, Mal91b, McL11, Neu73b, Par88, Sch74a, Wat90, Sau99, Sau05]. first-phase [McL11]. FitzGerald [War91]. Five [Dar02]. Fixed [Man04]. Flamsteed [Ruf13]. flawed [Van21]. flights [Dut85]. flow [Eck21, Wis81]. FLT [Cor08]. fluid [Aub17, Cra13, Lüt84a, Nem62]. fluidity [Cha15, FG23]. Fluids [Dob99b]. Folgen [Bac89, Vol89]. fonction [Dho86]. fonctionnelle [Jaë10]. fonctionnelles [Dho86]. fondamental [Gil91, Voe08]. fondements [Gis83, MS15a]. Fontana [ZH24]. fool [Fra98]. force [Nau10, Nau14, vB17]. Forces [Mel06, Muk74]. forefathers [Zan13]. forgotten [Dar20, Høy08, Moz16a, Not18]. form [Lam86]. formal [Im96]. formalism [McL11]. Forman [Ewa69]. formes [Bou13]. forms [Ace11b, Kat85]. Formula [Mil85]. formulae [Abr82]. formulation [BN78, GXB21]. found [Osw17]. foundation [Bie81a, Bie81b, Bie81c, Bie83b, Bie87, vdW71, vdW72b]. Foundational [Pal01, Bus19, Fer07]. Foundations [ABZ02, Bru67, Nav98, Sau99, Sau05, Bie82, Boo85, Fra85, Gio21, Lau89, Smi81, Toe86]. Fountains [Eck02]. four [Neu73b]. Fourier [HJB85, Arn83d, HH79]. fourteenth [Lin70b]. foyerz [AB22]. fra [Mai90, Mai94b]. fraction [Dut82]. Fractions [Rās78, Haw08a, MS15a, OS14]. fragment [FHR21]. fragmentation [Ace11a]. fragments [Yon69]. Framework [Tho05b, BC12, Tho05a]. France [AT14, Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, Del22b, Gis83, LM98]. Francesca [Fie97]. Francesco [ZH24]. Francis [Hom67]. Francois [Ray20, BS77, Oak18, Ste08]. französischen [Neu81]. Fréchet [Arb79, Arb81, Jaë10, Tay82, Tay85, Tay87]. Fred [Kra10]. Frege [Boo85]. French [AB17, AB19, Arb79, Arb81, BS77, BS84, Dah80, Del11, Del70, Dho86, Dug73, Dug76, Eis82, Eis86, Eis87, Eis91, Eis96, GP09, Gil91, GL12, Gis83, Jam88, Jao74, Ker86, Lam13, LD01, LM98, MT13, Mer02, Nab08, NB89, NTP04b, NTP04a, Neu81, PB06, Ras72, Rās74, Rās78, Ras80, Ras83, RDV01, Rou80, Ses77, Ses98, Sma11, Ver20, Vig00, Vin05, Vit02, Vit04, Voe08]. Frenkel [Fre74]. frequency [Kat08]. Fresnel [Arn83f, Ped00]. Freundlich [Hen94]. Fricke [AG09]. Friedrich [DF12]. Friendly [Ras83]. Frigyes [Rod15]. Frobenius [Haw74, Haw05, Haw08a, Haw08b]. frühe [Kau85, Sch82b]. Frühgeschichte [Neu73a, Scr61]. frühgriechische [Sza64]. frühgriechischen [vdW78a]. Frühzeit [Hof65]. frustum [BS22]. function [Ber66, Dho86, Dut81, Dut84, Dut91, Mon72, Wim12, You76]. functional [Dho86, SS82, Zan13, Zan14]. functionals [Tay82]. functions [AG09, Bil82, Bot77a, Cog14, Dut95, Dut96b, Fer07, FP12, Osw17, Ull89, Vol87, Vol89]. Fundamental [Rus98, Gil91, Voe08]. Funktionalanalysis [SS82]. Funktionen [Ull89, Vol87, Vol89]. funzioni [Bot77a]. Further [De 12, Amb12, Con18, Kno81].

G [Brå07, Lüt94, Pet61, Pro88]. G. [Gar91, Mai82]. Gâvre [Aub17]. Gaetana [Tru89, Tru92]. Galen [Too85]. Galilean [Nor04]. Galilei [BS22]. Galileo [Bas14, BS22, Gra65, Hah02, Kno99, Pal98, Pal01, Pal05, Pal06, Pou20, Seg80, Wis74, ZH17]. Galluppi [Mai94a]. Galois [Neu97, vdW72a, Edw90, Gal01, Kie71, Neu97, Neu06, Ste24b, vdW72a]. gas [PS10]. Gase [Sch75]. gases [Bru73, Bru74, Han77, Sch75, Tru75]. Gastvortrag [HY86]. gateway [Eck21]. gauche [GP09]. gauches [Del11]. Gauß [Rei73, Rei03, BF93, Bre84, Bul09, DF12, Dut96a, Fer07, HJB85, Mag22, Rei73, Rei03, She79, She94b, Wat90, Bie04, Rei03]. Gearing [LP15]. Geminorum [Sau08]. Geminos [vdW85b]. Geminus [BG96]. General [Arb81, SS98, Gio24, Rod15, Tay85, Tru79, EG78, Eis86, Ill89, Kox88]. générale [Eis86, Arb81]. Generality [Tho05b, Tho05a]. generalized [Cog22]. generated [Bel91]. generation [Pot74]. Genesis [NP06, Cog11, Edw80, Kno86, Kra81, Lam88, Lam89, Ruf63, Str81, BS77]. genetics [Wim12]. genuine [Jaë20]. geodesy [Bre84]. geodetic [She94b]. Geographic [Not24, Not23]. Geographike [Mar12]. Geography [GM17]. geomagnetic [Cha62]. Geometram [BN84, BN84]. Geometria [BN84, Gar91, Lit22]. Geometriae [Bus65a]. Geometric [Fri14, Rus98, SPJ01, Van06, Ray14, Vit02]. Geometrical [Lê17, Shk72, SS09, Blå16, Joh77, Kje08, Mae88, Mah68, Mal10, Smi81]. Geometrie [Boo85, Fre91, Neu73a, Toe86, AB22, Del11, Nab08, Voe08, Har08, Sch82b, vdW72b, vdW78a, Gal80, Bos81, Row24]. geometries [Lor16]. Géometrique [Ras70, Vit02]. Geometry [Art88, ABZ02, CEF17, Cor13, Dom23, HG05b, Ace11a, Amb12, Boo85, BN84, Bus19, Del22b, Del11, Eva76, Fre91, Gar91, Lüt95, Mol78, Nab08, Neu73a, PR23, Rec94, Rei73, Rei89, SS21, Sch82b, Sei61, Sei75, Wim12, vdW71, vdW72b, vdW78a, Gal80, Toe86, Voe08]. geophysics [Sch82c]. Geophysik [Sch82c]. Georg [Mes65, Dau71]. George [DS12, Mes65, Mil83]. Georgii [Mal16]. Germain [DF12, Del08, Sam90]. German [Art85, Art88, Bac89, Bec67, Bie81a, Bie81b, Bie81c, Bie82, Bie83b, Bie83a, Bie85, Bie87, Bie92, Bol72, Bot77b, Bre94b, Bur67, Bur71, Bur72, Bus65b, Bus98, Car86, CFH^{+}70, Con76, Dar93, ENP82, Fre91, Ger69, Ger82, Goe68, HY86, Hen01, Her63, Her70, HB86, Hof60, Hof65, Hof66, Hoy73, Hoy74, Hoy76, Hoy78, Hoy80, Jah87, Joh84, Kau85, Kno74, Kri68, Kub70, Lau65, Loh66, Loh68, Mes65, Mit65, Neu73a, Neu73b, Neu74, Neu78, Neu81, Neu97, Pai71, Pai74, Pai77, Pai82, Pet61, Pet63, Ras71, Rei73, Rei89, Rei92, Rei03, Sch82a, Sch86, Sch68, Sch71, Sch74a, Sch75, Sch82b, Sch73, Sch82c, Sch84, Sch85, Sch88, ST94, Sch74b, Scr61, SS82, SS86, Spa91a, Spa91b, Spa02, Str92]. German [Sza60, Sza64, Tót66, Ull84, Ull89, Vol87, Vol89, Was71, Zeh83, vdW60, vdW72b, vdW72a, vdW78a]. Germany [Cal76, Dar93, Pye82]. Gershon [Rab70]. Gerson [Man98a]. Geschichte [Bur67, Ras71, Rei73, Vol87, Ger69]. Geschichtliche [Bol72]. Gesellschaft [Sch88]. Gesichtspunkte [Bec67]. gestation [PR23]. Gesualdo [Mai94b]. Ghetaldi [Mai89]. Giant [Cen10]. Gibbs’ [Dai21, HH79, Nav98, Roc21]. Giochi [FGP81]. Giovanni [BN84, Cha16, Mid73, Van21]. Girard [AB17, AB19, AB22]. Girolamo [BN84]. Giulio [Zel09]. given [Sid18a]. Givens [Ace11b]. glass [Gou14]. glasses [Wes64]. gleichverteilung [HB86]. glimpse [Ren97]. Global [Cho11]. Glories [Ste00a]. Gmunden [GC24]. goal [GS08, GS09]. goal-year [GS08, GS09]. Goals [Kje02]. good [Fra98, VT18a, VT18b]. Gordon [CG80]. Gotha [Her70]. Göttingen [Im96, Pye79, Sch82a, Sch88]. Göttinger [Sch82a]. Gottlob [Boo85]. grains [CM15]. Grand [YT18]. Graphical [Mal10]. Graphs [Asc88]. Grasping [Tan09]. gravitation [Aok92, Aok96, Duc11, Eis91, Nor92, Ros69, Wil70, Mal91b, Ros65]. Gravitational [Zeh83, Blu22, Car21, Sau08]. gravitational-wave [Blu22]. gravitatione [Ruf12]. Gravitationstheorie [Zeh83]. gravity [BS22, Wes86]. great [McL11, Wil85, Moz19, vdW78b]. greca [Vit82]. Greco [BG96]. Greek [Ace11a, Ace11b, Ace19, BHM15, BS07, BJ00, Duk10, GB91, Hog84, Mah68, Neu73a, Sid18a, Sza64, Thu92, Ung75, Vit82, Wri00, vdW78b, vdW78a, vdW82, vdW84a, vdW84b, vdW84c, vdW85a, vdW85b, vdW88c]. Greeks [Was71]. Green [Lüt94]. Gregorian [Bie07]. Gregorie [Mal93]. Griechen [Was71]. griechischen [Neu73a]. Größen [Was71]. Größenlehre [Lau65]. Grosseteste [Eas70]. Group [Neu06, Dah80, Haw71, Haw72, Haw74, Hol17, Ste24b]. groupe [Dah80]. groups [Bur67, Gra80, Haw72, Hel22, Hol14, Sil79, Haw91]. growth [Joh79, Joh81]. Grundlagen [Bie85, Boo85, Sch82b, Spa91b, Toe86]. Grundlegung [Bie81a, Bie81b, Bie81c, Bie83b, Bie87, Bie92, vdW72b]. Grundlegungen [Bie82]. Grundzüge [ENP82]. Gu [LA84]. guest [HY86].

H [Par85, She71a]. H. [She91a]. ha-Meshihah [FG23]. Habash [Yaz18]. Haggāg [Bre94b, Bre94b]. Half [Cou04, Hof65, Jah87, Lam13]. Hälfte [Hof65, Jah87]. Halley [Bro05]. Hamilton [Mat78]. Hansen [WH14]. Hariot [Kal77]. harmonic [Dar07]. Harmony [Dia99, Pye82]. Harriot [Loh66, Big19, Gou14, Loh66, Loh79, Pep68, Ste00a, Ste08]. Hartree [Zan13]. Hasib [Yaz18]. Hasse [FS05, FS07]. Hastings [SS98]. Hauksbee [Hom67]. Hauptsatzes [Bie81a, Bie81b, Bie82, Bie83b, Bie87]. Hausdorff [AB12]. having [Hof66]. Haytham [Ras70, Ras80, Sab82]. Head [MP22]. heart [dC16]. Heat [Gea10, Arn83d, Hoy78, Ler72, Bie81a]. heated [Mal91a]. heavens [Ber91]. Heaviside [Lüt79, Pet87, Roc21]. Hebrew [Gol03]. height [Gol77]. Heinrich [BYS^{+}21a, vdW72a, BYS^{+}21b, HD95, vdW72a]. Heisenberg [Bel85]. heliocentrism [Car18]. Heliostat [Mid73]. Helmert [She95b]. Helmholtz [Bie81c, Bie83a, Buc01, PR23, RP23]. Helmholtzens [Bie81c]. Helmholtz’s [PR23]. Helmholtzschen [Bie83a]. Helmut [FS05, FS07]. Hendrik [Kox88, Kox13, Sch84]. Henri [Mil73]. Henrik [GS94]. heptagon [Hog84]. her [Del08]. Heritage [Høy08]. Hermann [Bie81c, Bie81c, Cor97b, Pye77]. Hermite [Bot77a]. Hero [Mas15, Men23]. Heron [Key92]. Hertz [Buc01, BYS^{+}21a, BYS^{+}21b, Don91, HD95]. Hertzian [Don91]. Hesses [Haw88]. Heuristic [Man98a, Ace11a, CGM87]. heuristical [Pot74]. heuristics [CG20]. hexagram [Del20]. Hibbur [FG23]. Hidden [MT97]. higher [Bos74]. higher-order [Bos74]. highlight [CM15]. Hilbert [SS86, BS84, Con76, Cor97a, Cor99, EG78, Gio21, Lac00, Sau99, Sau05, SS86, Toe86]. Hill [GZKB^{+}12]. hindered [GG85]. Hindu [MB03, vdW78b]. Hipparchus [Ace03b, Duk02b, Mae98, Moz24a, NB89, Swe80, Too74]. Hippopede [Yav01]. Hire [Bri21, SS05]. Hisab [Ses77]. Histoire [Eis82, RDV01, Del11, Dho86, Gil91, Jam88]. Historical [Frö94, JP13, Kje02, She88, Van92, Dut81, McH71, Nem62, Pot74, RKH^{+}21, Seg94, She73a, She73b, YT18, ZH17, Bol72]. historigraphical [Spa02]. historiografische [Spa02]. historique [Bus19]. History [Alv99, BG08, Dal01, Går89a, Kra03, Nis05, Nis09, Nis11, Ped00, WL23, ZH17, BN78, Ber68, BGTZ14, Bru68, Bur67, BSS88, Cou10, Dar20, Del11, Dem83, Dho86, Dut84, Dut91, Dut95, EG78, Fer96, Fre77, Gal17, GG85, Gra84, Har08, HJB85, Jam88, Kra23, Lam88, Lam89, Muk74, Neu73a, Ras71, RW12, Rei73, RDV01, Ros71, SV24, Sch71, Scr61, She77a, She80, She82, She84a, She84b, She85, She93a, Ung75, Vol87, Zam10, BG09, Eis82, Ger69]. Hiyya [FG23]. Hobbes [Don21]. Hochatmosphäre [Sch73]. Hoher [Joh84]. Holbroke [Not18]. holes [Eis91]. Homocentric [Yav98]. homoclinic [And94]. homogène [Jao74]. Homogeneity [Mue70]. homogeneous [Jao74]. Homologiebegriffs [Bol72]. homology [Bol72]. homotopic [Van92]. Hopkins [Par88]. horizon [Bru81]. Horse [Dar03]. Hoyer [Joh84]. Hoyle [Kra10]. Hsiang [Yon69]. hsüeh [Lam79]. Hui [Yon69]. huitième [Vig00]. Humboldt [Cha62]. Hurwitz [OS14, Osw17]. Huygens [Hof66, Buc80a, Buc07, Hof66, Nau15, Sak82]. hydraulics [Eck24]. Hydrodynamics [Dar02, Dar03, Eck24, Fer14]. Hydrogen [Gea10, MR10]. hydrostatics [Cha15]. hyperbolic [Gal17]. Hypercomplex [Haw72]. Hyperelliptic [Ull00]. hypergeometric [Dut84, Fer07]. Hyphegesis [Mar12]. Hypothese [Was71]. Hypothesen [Bre94b]. Hypotheses [Duk09, Bre94b]. hypothesis [KL84, Mae90, Pér09b, War91, Was71, vP91, NP06, Rit15].

I. [Pet63]. Iberian [CG15]. Ibn [GC23, Moe74, Jao74, CG94, CG15, Gol77, Kin73, Man98b, Moz19, Moz24b, Ras70, Ras80, Sab82]. idea [Sau08]. ideal [Edw80, FP12, Han77, PS10]. ideas [Nem62, Ros71]. Ideengeschichte [Sch71]. ihr [Bie85, Hof66, SS82]. ihre [Bie85, Hof60, Hof66]. ihren [Hof66]. ihrer [Sza64]. II [Arn83b, Asc88, BG94, Bri09, Cor13, Dea82, Del16c, Dia94, Fow80, Fow82, GS09, Joh81, NB89, Pai74, Pou09, RA20, Tay85, Wil80b, dJ19b, vdW80b, vdW84b]. III [Arn83c, BC12, Bri10, Pai77, Ste20, Tay87, Vit02, dJ21, vdW84c]. Il’ich [Fre74]. im [Bie87, Car86, Fre91, Her70, Pet61, Rei03, Sch74a, SS82, Sza64, Tót66]. image [Gou18]. images [Lin68, Lin70a, Lin70b, Man89, Man92, Str81]. imagination [Gou14]. Immanuel [GC17]. Impact [Dai21, Kal77]. Impasses [Eis87]. imperial [CM15]. implications [FHR21, Mas16]. import [DPC95]. Importance [Pou04]. important [FHR21]. impossibility [EQ21]. inception [Ehb22, Ehb23]. Inclined [Hah02]. include [BG08]. incommensurabilità [Mar80]. Incommensurability [Bor07, Gra61, Kno81, Mar80, Kno82]. incommensurable [Was71]. incomplete [Dut81]. inconsistent [BY16]. increased [GGHPMR21]. increments [Fei85]. Indecision [Fis99]. independent [Mil85]. indeterminacy [Bel85]. indeterminate [BS77, Ses77]. indéterminée [BS77]. Indéterminées [Ses77]. India [Duk05a, WH19]. Indian [CG15, Duk05a, Duk07, Duk08, Thu92, vdW60]. indische [vdW60]. Individual [Pal85]. Individuo [Pal85]. indivisibles [And85]. induced [RD94]. Induction [Ace00, BHM15, Kyr21, Rab70, Ras72]. Inequalities [Kje02]. inequality [Wil85]. Inferior [Van06]. infinite [Ber68, Fis81, Lau89, Sak82, Bus65b]. infinitely [Bil82]. infinitesimal [Bas14, Fis81, Hof65, Lau89]. Infinitesimalmethoden [Hof65]. infinitesimals [Art13, RA20]. infinitorum [Ste01]. Infinity [Kno99, Del70, Fra97]. influence [Bel85, Bot77b, Bra85, BN93, Car21, Eis91, Fer14, Loh68, Ros15, Wil82]. influenced [vB17]. Influences [Sen98]. Inhalt [Hof66]. Inherent [Mel06]. initial [CGM87, Ehb22, Ehb23, GG85, RD94]. inkommensurablen [Was71]. inner [Gru17]. innovator [Høy88]. Inscription [IJ19a, IJ19b]. inspiration [BHM15, Del22c]. instance [Raw82]. instrumental [Pro88]. Integrability [Pou01]. Integral [AT14, Ber66, CEF17, Dea85, ED09, Gil91, Hal90, Gil91]. Integrals [Ull00, Ayo84]. integration [Sma11]. intellectual [Høy88]. Intelligibility [Mit19b, Mit19a]. Interactions [Lüt95]. interna [Mar93]. internal [Mar93]. Interpretation [Dob99a, KSFW90, Pou06, Bec73, Ber91, Dav90, Gua10, Kno76, Sme70, Ste22, Zai94]. interval [Ste24a]. Introduction [Jae98, Bri10, GB91, Lam79, Sch75, Ull89]. intuitionism [McL11]. Invariance [NP04, Joh79, Joh81]. Invention [Rās78, Dea85, Ker86, Mid73, vdW68]. Inverse [Wei00, Aok92, Aok96, Gui16, Nau10, Scr61, Scr63, Twe91a]. inverse-cube [Gui16]. inverse-square [Aok92, Aok96]. inverse-tangent [Scr61, Twe91a]. inversen [Scr61]. Investigations [Mar12, Nor04, Buc80a, Joh77]. involution [AB17, Del22c]. involutions [SS21]. irreducibilis [Con15]. irrelevance [Nis09]. irruption [Nab08]. Isaac [Gre96, Osa13, Wes64]. Ishiwara [HY86]. Ishiwaras [HY86]. Ising [Nis05, Nis09, Nis11]. Islamic [Moz16b, Ses91, Ses03, Shi03]. isolation [Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84]. isomorphicism [Sch74b]. isomorphism [RW12]. Isomorphismus [Sch74b]. Isospin [Ras71]. issues [Mer02]. Istanbul [MS15b]. Italia [RS98]. Italian [BF93, Bot77a, BN84, CG80, FGP81, Gal80, Gar91, Gue84, Mai89, Mai90, Mai94b, Mai94a, Mar80, Pal85, PT96, RS98, Vit82]. italiana [BN84]. Italy [AT14, ABZ02, RS98]. itinerary [Tup22]. IV [Arn83d, vdW85a]. IX [Arn83e].

J [Bie92, GL19, SPJ01, She73c, Top71]. J. [Bie92, Fra83, Fra85, Lüt94, Rec94, SPJ01, She71a, Top71]. J.-É. [SPJ01]. Jacob [GC17, Cha91]. Jacobi [Cog14, FN23, Haw91]. Jacques [GGHPMR21]. Jahen [Her63]. Jahre [Bie81a, Bie81b, Bie81c, Her70, Joh84]. Jahren [HB86]. Jahrhundert [Bie85, Bie87, Car86, Pet61]. Jahrhunderts [Jah87, Hof65]. Jakob [Lor16]. Jalali [Yaz20]. Jamblique [Vin05]. James [Fre74, Mal93, Par88, Twe92]. Jamshīd [Van06]. Janet [IM21]. János [Tan09]. Jeffreys [GCWL23, WL23]. Jensen [Str92, Str92]. Jensen-Transformation [Str92]. Jerusalem [Zel09]. Jesse [Sch73, Sch73]. Jewish [Bel88]. Jing [Gua15]. Jiu [Lam94]. Jiushao’s [Yiw20]. Johann [Spe96]. Johannes [Fie84, Fie97]. John [GC24, Jon10, Kje01, McC66, Not18, RKH^{+}21, Ste01, Twe91a, Vig00]. Johns [Par88]. Joint [BG09, Del22a]. Jordan [Bel85, Gug77, Lac00]. Jordanus [Høy88]. José [CPFA07]. Joseph [Par88, GC23, Ham76, Lüt84a, Par85, War91]. Jost [WH19]. Journals [BG09]. Journey [Kje01]. Julius [Lor16, OS14]. Jun [HY86]. Jupiter [Abr82, Car21, Coh98, Eis96, Wil85]. Jupiter’s [Oss18]. justification [RA20].

K. [Arb81, Lüt94]. Kāshī [Van06]. Kaestner [Pet61, Pet61]. Kähler [Cog11]. Kammâd [CG94]. Kammad [CG15, Man98b]. Kant [Pet63]. Karaji [Ras72, Ses77]. Karl [Dug73, Vie21]. Kashi [AH15]. Kepler [CR19, Dav03, Del16a, Fie84, Fie97, Gou18, Mae90, Spe96, Str81, Wil70, Wil85, Dut97]. Keplerian [Mae88]. Kepler’s [CG20]. Kerala [RR86]. Keralese [RR78]. Khazini [Moz22]. Khwārizmī [Kau85, MB03]. Khwārizmīs [Kau85]. Khwarizmi [Yaz11a, Yaz11b]. Killing [Haw82]. kinds [EQ21]. Kinematic [Sha73]. kinematics [CFH^{+}70]. Kinematik [CFH^{+}70]. kinetic [Bru73, Bru74, Tru75]. kinetics [KL84, Lai85, Lai88, Zam10]. Kirchhoff [BY16, Dil92]. klassischen [Hoy76]. Klein [Bur72, Bur72, CG80, Hel22, Kay23, Zam10]. Klein’s [Row24]. km [Sei76]. Knight [Ses03]. Knowledge [Buc06, Fis66]. known [CM15]. Kohn [Zan14]. Kolmogorov [BN78]. König [Fra97]. Konstruierbarkeit [Pet63]. Konstruktionen [vdW78a]. Kontinuum [Spa02]. Korean [Shi03]. Kramers [Zam10]. Kreismessung [Hof66]. Kronecker [ENP82, Ver20]. Kroneckers [ENP82]. Krümmungsbegriffs [Ger82]. Kudurru [KSFW90]. Kummer [Edw77, Cor08, Edw75]. Kuo [Gua10]. Kuratowski [Arb81]. Kurven [Hof65]. Kyoto [HY86, HY86].

L [Bie92, Fra83, Fra85, Wil82, vdW72b]. L. [Bie92, Dav90, She72]. Laboratories [Lai88]. laboratory [Duc11, HD95]. Lacaille [Wil80a, Wil80b]. Lage [Har08]. Lagrange [FP12, Fra83, Fra85, Ham76, Kie71]. Lagrangian [LP15]. Laguerre [MR10]. L’algebra [Wag10]. lálgèbre [Nab08]. Lambda [Ear01]. Lambert [Sha78, She71a]. Landen [Sma11]. language [Ace11b, Ehb22, Ehb23]. Laodicea [McC96]. Laplace [Bik78, Dal82, Dea81, Dea82, Dea92, Pou20, She76, She77b, Wil85, Zab88]. Laplacian [Arn83b]. L’archéologie [Eis91]. large [McH71]. Larmor [War91]. laser [Bro16]. last [Edw75, ST94, Del08]. lasting [Hof66]. Late [Mit19b, RSW04, Bel91, BN84, MZ24, Mit19a, Pal08, Pye82, FHR21, Fra13]. late-twentieth [Bel91]. lateinische [Kau85]. lateinischen [Hen01]. later [Whi61]. Latin [BG96, CG13, Cio22, Cio23a, Cio23b, Hen01, Kau85, Not17, Not23, Not24, Ung74]. Latitude [Van06, Moz16b, Not23]. Latitudes [GC04, Moz16b, Rid78]. Law [Dav03, Ruf63, Aok92, Aok96, CG20, Dar20, Don21, Eas70, Loh61, Pou20, RC23, Sha79, Hen01, Pou06, Ros65, Smi82, Web65]. laws [CR19, Col77, Wil70]. leading [JW22]. learned [Yea11]. Least [Dia99, Dut90, Dut96a, Mag22, She93a, Wat90]. least-squares [Mag22]. Lebesgue [GP09, Mon72, Net20]. L’Échelle [Vig00]. lecteur [Sma11]. lecture [HY86, ST94]. Lectures [Ull89, She94c, Vie21]. legacy [RP23]. Legendre [HG05b, Sma11]. Lehre [Spa91a]. Leibniz [Kno74, Ait72, Art13, Dut82, EQ21, Kno74, Kno99, RA20, SS05, Scr63]. Leibnizian [Bos74]. lemma [Fra97, DF20, Pou01]. Lemmas [Pou98, MT97, Sab82, Dob01]. lemmatic [Ace11a]. lemniscate [Ayo84]. Length [Abr84a, Bel98, Not18, Ste22, Swe80]. lens [ZH24]. lensing [Sau08]. Lenz [Nis05, Nis09, Nis11]. Leonard [Fen98]. Leonardo [Fie97, Thr94]. Leonhard [Mal00, Mil85, Ver13a, Ver13b]. Leopold [Ver20]. letter [Wei78]. lettere [PT96]. Letters [AG09, Cal76, Ren97, Arb79, Sal74, Sch84, Bur71, Bur72, PT96, vdW72b]. Lettres [Arb79]. letzter [ST94]. level [Eis86, Kra10]. Levi [Rab70, CT24, IR16, Man98a, Rei92]. Levi-Civita [CT24, IR16, Rei92]. Levi-Civitasche [Rei92]. Lewis [MZ24]. L’Extraction [Rās78]. Li [LA84, Gua15]. Liber [Han11]. Library [Kau85]. Lichtstreuung [Sch74a]. Lie [Cic15, Haw72, Haw82, Haw88, Haw91, Kay23]. liés [Dho86]. Lieu [Vit02]. life [Dug76, FGP81, Fra13, Ste22]. Light [Dob98, Bro67, Bro16, Che97, Eis91, Haw74, Mal91a, Sak82, Sch74a, SFS08, Sha73]. like [Kno92]. likelihood [Mag22]. Likely [Jon99]. limitations [Smi82]. Lindley [GCWL23, WL23]. L’induction [Ras72]. line [PR23]. linéaires [Jaë10]. Linear [Kje02, Bro81, Dem83, IM21, Lüt84b, McH71, SV24, Sch77, Shk72]. Linearum [Mai90]. Lines [SS05, Ste01]. L’infini [Del70]. link [Pap17]. Linkages [Lam87]. Liouville [Lüt84a, Lüt84b]. livre [Jao74]. Lobachevskii [Tan09]. Local [Cho11]. locali [Wis74]. location [Gou18]. locus [Høy08]. L’OEuvre [Rou80]. logarithmic [Pep68]. Logic [MP22, Fer96, Mar93, Zai94]. logica [Mar93]. logicism [McL11]. long [Del16b, Del16c, Duc11, Nau15]. long-standing [Nau15]. long-term [Duc11]. longitude [Not24]. longitudes [Duk08, Thu92]. Longomontanus [GC24]. Look [Pal06, Brå07, Key92, Kno92, Mah68]. L’optique [Eis96]. Lorentz [Sch84, Ill89, Kox88, Kox13, Nor92, Sch84, War91]. Lorentz-covariant [Nor92]. Lost [Gui16, Cou10, Del20, GXB21, GGHPMR21, Hog86]. Lösung [Scr61]. Lotka [RC23]. Louis [Bou13, GL19, Kub70, LD01]. Love [Ren97]. low [Eis86, GG85]. low-water [Eis86]. Luca [Fie97]. lumière [Eis91]. Lunar [Bri07, Duk07, Mae06, Ste06, Bri09, Bri10, CG19, Gua15, Jen72, Jon83, McC96, Moz14, MS15b, Moz24a, Shn17, Ver13a, Ver13b, Yaz18]. Lutheran [Fie84]. Lyell [Kau85].

M [Par85]. M. [Arb81, SPJ01]. Machin [Twe91a]. MacLaurin [Mil83, Mil85, Twe91b]. Macroscopic [KC00]. Made [HG05a, Ste08]. Maghribi [Moz14]. Magic [Kra03, Ses03, Vin05]. magics [Ses98]. Magique [Vin05]. magiques [Ses98]. magistrales [Cha16]. Magna [Con15]. magnetic [dC16]. magnetism [Wis81]. Magnification [ZH12]. magnitude [Vig00]. Magnitudes [Ehr06, Vig00, Bas14]. main [Bie81a, Bie81b, Bie82, Bie83b, Bie87, Con18, Nem62]. Mainly [Tay85]. Making [Bie04, Dia99, Fis99, Mar21]. Malus [Buc80a]. management [CM15]. Manifolds [Ull00]. Mannigfaltigkeitsbegriff [Sch82b]. Manuel [Ses98]. manuscript [BS22, Blu22, GGHPMR21, Gug77, Hen01, RKH^{+}21, RD94]. manuscripts [Del08]. Manuskriptes [Hen01]. map [Big83, Raw82]. maps [Ber91]. Maragha [Moz14, Moz16b, Moz18b]. Maria [Tru89, Tru92]. Mariano [CPFA07]. Marino [Mai89]. Markov [She89b, She89a]. Mars [Abr84b]. Martianus [Car17]. Martin [Jah87, Jah87]. Maslama [Lor95]. mass [Cra13, Lüt84a]. Masses [Coh98]. masterly [DF18]. Masters [Uli15]. Matar [Bre94b, Bre94b]. matematica [Vit82]. Matematiche [Mai94a]. matematici [Mai94b]. Materials [Fre74]. Materiewellen [Kub70]. mathematic [Mai94b]. Mathematical [Duk05a, Hah02, Kje01, LS86, Pal06, Pou98, She73b, Arn84, BHM15, Bel88, Bos84, CM15, DS12, EQ21, Fis66, Fra89, Fri19, Gar89b, Gil74, Høy88, Høy08, Jaë20, Lam79, Lam94, Lam97, LM98, Loh61, Pai71, Pai74, Pai77, Pai82, Par88, Pot74, Rab70, RKH^{+}21, Ron60, Rou80, Sei84, She72, Sme70, Smi78, Spa91b, Vit82, Vol87, Whi61, Wil82, Ras72]. Mathematician [Dar03, Arn83c, CD13, Jah87, Sch68]. mathematicians [LD01]. Mathématiciens [LD01]. Mathematics [Dav03, Ehr06, Fer99, Fra13, Mit19b, Ses03, Ace11b, Ace19, Bus98, CGM87, DG20, Fri14, Går89a, Jah87, Kai11, LA84, Mat85, McL11, Mit19a, Pal09, Pye79, Pye82, RR78, RR86, Rid79, Sch71, Sch17, Sei78, Ses85, Sid18a, SS82, Ung75, vdW80a, vdW80b, Mai94a]. Mathematik [Bus98, Jah87, SS82, Sch71]. Mathematiker [Sch68]. Mathematikers [Jah87]. mathématique [Arn83e, Ras72, Rou80]. Mathématiques [Ray20, LM98, Bri21]. Mathematische [Pai71, Pai74, Pai77, Pai82]. mathematischen [Spa91b, Vol87]. Mathesis [Boc09]. matrices [Ber68, Haw77, Haw08b]. matrix [Cao91, CG19]. Matter [Epp98, Arn83g, Kub70, War91]. mature [Cal76]. Maurice [Arb79, IM21, Tay82, Tay85, Tay87]. Max [Gug77, Kle61]. maxima [Str68]. Maximum [Net99, Mag22]. maximum-likelihood [Mag22]. Maxwell [Bro67, Dar93, Dia94, Hei70]. Mean [Duk08, Duk09, Mer07, BG96, Duc11, Kus89, Mae88]. meaning [Ace09, Hof60, Hof66, Pal85]. measurable [Spa91a]. Measurement [Mit19b, Dut93, GB91, Hof66, Kno76, Mit19a]. Measurements [Buc06, Not23, Moz14]. measuring [Kje08]. mécanique [Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, Arn83h]. Mechanical [Cor99, Bie81a, Bie81b, Bie81c, Bie82, Bie83b, Bie87, Bie92, Cha15, Hoy78, Moz22, Bie85]. Mechanics [Dia99, Lac00, Nav98, Aub17, Bru67, CR07, Cra13, Fra83, Ger69, GXB21, Im96, Lüt95, Pal08, Ren97, Tru60, Wil87, vB17]. mechanische [Bie81a, Bie81b]. mechanischen [Bie82, Bie83b, Bie85, Bie87, Bie92]. mechanischer [Bie81c]. mechanisches [Hoy78]. mechanism [CE14, EC19, Top71, IJ19a, IJ19b]. mechanistic [Bec73]. mediaeval [Ses85]. medical [She82]. Medicine [BG09]. Medieval [Syl71, Bus98, CG19, Cor13, Moz16b, Moz18a, Not17, RR78, RR86]. medium [Ait72]. meeting [Her70]. Melacrinò [Mai94b]. members [Sch88]. Memoir [Neu06, BG94, Del22a]. memoirs [Gue84]. Memorie [Gue84]. Men [Ros71]. Menelaos [NTP04b, NTP04a, NTP04b, NTP04a]. meng [Lam79]. Mengoli [ED09]. mensuration [Pot74]. Mercury [Rid78, dJ21, vdW82]. meridional [Pep68]. Merton [Syl71]. Meshihah [FG23]. meson [Muk74]. Mesopotamia [Ste00c]. Mesopotamian [Fri14, Fri19]. meßbaren [Spa91a]. Metaphysical [Eas70]. meteorologies [Vit02]. Météorologiques [Vit02]. Meteorology [Dal01, She84b]. meteoroscope [GJZ^{+}23]. Method [Bro05, And85, Dut90, Dut96a, Fei85, FP12, Hof60, Mai94b, Nau94, Net20, Sch71, Scr63, She80, She84a, She84b, She85, Wat78, WH14, vdW60]. Methode [vdW60]. Methoden [Hof60, Sch75]. méthodes [Bac63, Sma11]. methodological [Duc11]. methodological-experimental [Duc11]. Methodologie [Vol87]. Methodology [vD04, Smi82, Vol87]. Methods [Buc07, Tho05b, Fra83, GS09, Hof65, Sch75, Str68, WH19, Sma11, Tho05a]. metodo [Mai94b]. metrics [RP23]. Michel [Bus19]. Michell [Vig00]. Michelson [vD09]. mid [Bel91, Eva76]. mid-nineteenth [Bel91]. mid-twelfth [Eva76]. middle [Eva76, You76, GC23]. might [Hol14]. mind [Rit15]. minima [Str68]. Minimax [Kje01]. minimum [Rec94]. Minkowski [Cor97b, Kje08, Pye77]. Minnesota [DJ07a, DJ07b]. mirabilis} [Ace19]. mirror [Sza64]. mirrors [Ace11a]. Miscellaneous [Ste06]. Mischkristallisation [Sch74b]. Misconception [Ehr06]. Mises [SS06]. Misfortunes [Ste00a]. misinterpretation [Bac89]. missing [Sau05]. mistake [Blu22, RD94]. misunderstand [Hol17]. Mitchell [Vig00]. Mitgliedern [Sch88]. mittelalterliche [Bus98]. mixture [Sch74b]. mobile [NB84]. Model [BJ00, Van06, Bec73, CG86, Car20, Don21, Moz16a, Moz22, Nis09, Nis11, Sch17, Nis05]. Modeling [Bri07]. Models [Dal01, Duk05a, Pou04, CRT10, CR19, Gea85, Not17]. Modern [Buc06, Fer99, Ald09, BG08, Dea88, Joh79, Joh81]. modernity [VT18a, VT18b]. Modes [Mit19b, Hei70, Mit19a]. Modesto [Fri20]. modular [Osw17]. module [Sma11]. modules [Sma11]. Möglichkeit [Mit65]. moitié [Lam13]. Moivre [Sch68, Hal88, Sch68]. molecular [CRT10]. molecules [Mac23]. Møller [Kra92, Roq92]. moment [Jao74]. Moments [Cap02]. monatomic [PS10]. Monge [Bik78]. Monocycle [Bie83a]. Monozykel [Bie83a]. month [Ste24a]. months [EG78]. Moon [Aok96, Car14, Ste24a, Too74, Abr82, Abr86, Aok92, BS07, Car17, CG19, GC24, Kus89, Mer07, Moz24b, Cio23a]. Moon-test [Aok96, Aok92]. Moore [SS98, Wil82]. Moral [Men07]. Mordell [GL19]. Morgan [Big83]. Morley [vD09]. Mortalità [FGP81]. mortality [FGP81]. Moschopoulos [Ses98]. mosques [YT18]. Motion [Bru68, Dob99b, Mal00, Pou04, Abr80, Abr84b, Abr85, Abr86, Ait72, BG96, CG19, CFH^{+}70, Gal17, GC23, Jon83, Kus89, Nau14, Oss18, Wis74, vdW82, vdW85b]. Motions [Duk09, Man04, Van06, Duk08, Gra61, Mae88, Yea11]. motivation [Bos84, Mal16]. Motivations [Kje02, JW20]. Motive [Jah87]. Motives [Jah87]. Mott [FT13]. motu [Wis74]. Move [Ses03]. movement [Gue84]. Movimento [Gue84]. Mr [Gou14]. MS [Kau85, Kau85]. Mu [Gol77]. Much [Fra04, GH00]. Muhyi [Moz14]. multicolored [ENP82]. Multiple [Bot77a]. Multiple-periodic [Bot77a]. Multiples [Ace03a]. multipli [Bot77a]. Mundi [McG66]. Muqtabis [CG94, CG15]. Music [Bor07, Dos75]. Mu‘tabar [Moz22]. mutual [Fer14]. my [Kno82, She83, Wei80]. mysteries [Bre94a]. mysterious [Yaz18]. mystic [Del20]. myth [Ewa69, Kra10]. myths [Ewa69, For69].

n [Rās78]. nach [Sch75]. Nachlaß [Neu78]. Nachwirken [Hof66]. narration [Bre94b]. Nascita [CG80, FGP81]. Nathan [Blu22]. Natural [Epp98]. nature [Arn83g, Sar12, WL23, Wil82]. natures [Wag10]. Navier [Dar02]. Neapolitan [CD13]. Near [Jon04, Bru81]. Necessity [NP04]. need [Ung75]. negative [Ses85]. neglected [DF18, Roq92]. Negro [RGM19]. Nekrasov [She03]. nel [BN84, Pal85]. nella [Bot77a, Gal80, Gar91, PT96, Vit82]. Nemore [Høy88]. neolithic [Mat85]. neue [KSFW90, Was71, Bec67]. Neumann [Kje01, Lac00, Net20, RW12, RKH^{+}21]. Newly [Bel98, BS22, Dal86]. Newly-Discovered [Bel98, Dal86]. Newton [Aok96, Aok92, Bec73, Bra03, Car86, DF20, Coh98, Dob98, Dob99a, Dob99b, Dob01, Gre96, Gro98, Gui16, Loh61, McG66, Mel06, Mil17, Nau94, Nau14, Osa13, Pou92a, Pou98, Pou01, Pou03, Pou04, Pou06, Pou09, Pou20, Ros65, Ros69, Ruf63, Ruf00, Ruf12, Ruf13, Scr63, Sha79, She71b, Shk72, Spe96, Wei00, Wes64, Wes86]. Newtonian [Eis91, Eis96, Rid80]. newtonienne [Eis91, Eis96]. Newtons [Car86, Dar20]. Nicholson [McC66]. Nichteuklidische [Fre91]. Nicolas [Zeh83, Pap17, Ray20, Zeh83]. Nicolas-Auguste [Pap17]. Nicole [Gra61]. Nicomedes [Sei66]. Niels [GS94]. Nile [Raw82]. nine [Del22a, Lam94]. nine-point [Del22a]. Nineteenth [Dar03, Dea88, Bel91, GG70, Sil79, SPJ01, Tay74]. Nineteenth-Century [Dar03, SPJ01]. No [HA98, Ace19]. Nobody [Hol17]. Noether [McL11]. noirs [Eis91]. noise [Yea11]. nombre [Jam88]. Nombres [Ras83]. Nombrils [AB22]. Non [Sha78, BS85, Bil82, Ker86, Kra97, Moz16b, PS10, SV24, Shk72, Tan09, Fre91, Ehr06]. Non-additive [Sha78]. non-analytic [Bil82]. non-Archimedean [Ehr06]. non-Euclidean [Tan09, Fre91]. non-linear [SV24, Shk72]. non-problem [Kra97]. non-Ptolemaic [Moz16b]. non-standard [BS85]. non-statique [Ker86]. non-statistical [PS10]. Noncommutative [Cou04]. nonnull [Mai89]. Nonnullae [Mai89]. nonstatic [Ker86]. Nordström [Nor92]. Norm [MB03]. Normal [GS08, Jon04, RSW04, Ste20]. notate [Fri20]. notation [Fri20, Pal85]. Note [Gal01, Lüt94, Sau99, Vit02, Wat78, Dav90, Dut97, Edw90]. Notes [Arb79, Dug76, Ste08, HD95, Jaë10, RGM19, Sch82b, She83, Dug76]. Nothing [Fra04]. notion [AB17, Jao74]. Notions [Epp98, Tan09]. Notizen [Sch82b]. Nova [Sau08]. novas [Mal16]. Novel [Van06]. nuclear [Muk74]. Number [Cor08, Kra03, Bel10, Bul09, Hof60, Jam88, Sch86, Ses91, Spa91b]. number-theoretical [Bul09, Hof60]. Numbers [Ace05, AA72, Dau08a, Dau08b, BC20, Car09, Duk10, Fri19, Haw72, Ras83, Sam90, Spa91a, VT18a, VT18b, Wag10, vR63]. numeral [Lam86, Lam87, Lam88, Lam89]. numerals [DEHY76]. Numerical [CPFA07, Gal01, BC20, Rās74, Del70]. numérique [Del70]. numériques [Rās74]. numerorum [Bus98, Bus98].

O [Joh84]. O. [Joh84]. Obituary [Jon10]. object [Kje08]. obra [Mar93]. Obscurity [Pal01]. observability [ST94]. Observation [Jon99, Pro88, Wil80a, Wil80b]. Observations [BSS88, Jon04, Fra89, GB91, MS15b, Moz18b, She72, She73b, She93b, She94b, dJH20]. observatory [Moz14, Moz18b]. ODEs [DG20]. oeuvre [Bou13, Dug76, MS15a]. office [Pai71, Pai74, Pai77]. Ohm [Jah87]. old [LA84, Pér09a, Pér09b]. Olinde [Gra80]. Omitted [Bra03]. One [Pou04, DJ07a, Pou92a]. One-Body [Pou04, Pou92a]. operational [Lüt79]. operations [Kop71, Jaë10]. Operator [GXB21]. optica [Del16a]. Optical [Buc07, Str81, BY16, Che97, Kno92]. Optics [Buc80b, Bro81, Eis96, Gou14, Ron60, Sha73, Smi81, ZH17]. Optique [Ras70]. oral [Mat85]. Orbit [Mae98]. Orbitae [Car20]. Orbital [Nau14]. Orbits [Wei00]. order [Bos74, Dem82, Sil79, Ste24b]. ordering [Kno78]. ordinary [Lüt84b, McH71, Sch77]. Ordinates’ [Osa13]. Oresme [Gra61]. orientation [YT18]. Origin [CG80, Abo08, Bil82, CG15, Dal82, Del22c, GZKB^{+}12, GM17, Gul87, Jao74, Kra97, Mat85, Pal85, Ren97, Sau08, Sei61, Sei62, Sei78, SC80, Sei81]. original [Mal16, Van21]. origine [Jao74, Pal85]. Origins [Fra94, Ber66, Col77, Con18, Dar07, Fra97, Haw71, Haw08a, Lai85, Lam86, Oss18, Rab70, Sar12, Sch17, Smi78, Toe86]. Oscar [BG94]. oscillating [Gru17]. oscillations [GL12]. other [Vit04]. Otto [Sch73, Sch73]. our [Lam86, Lam87, Lam88, Lam89]. Outer [Duk05b, dJ19a]. Outline [Mar93]. Outlines [ENP82]. outlook [Cal76]. Ovals [Pou01]. overview [Lam94, Lam97, Neu81]. Oxford [Kau85, Lai88]. Ozanam [GGHPMR21].

P [Joh84, Jon10, vdW72b]. P. [Arb81, Ewa69, Gar89b, Joh84, Mai94b, Mai94a, She76, vdW72b]. Pacioli [Fie97]. page [Sau05]. pairs [Moz24a]. Pancasiddhāntikā [Abr85, Abr82]. Paolis [ABZ02]. Paolo [Ayo80, Mid69]. paper [Ewa69, Gra80, Mat78, GH00]. papers [Lüt94, She83, Shk72]. Pappus [Ber03, Cio22, Kno92, Ung74]. papyrus [Gil74]. Parabola [Mai89]. paradox [Dut88, GCWL23, WL23]. parallel [IR16, Pet61]. Parallelenforschung [Pet61]. Parallelenproblem [Pet61, Tót66]. parallelism [Cog15, Sal74, Tan09]. Parallelogram [Mil17]. Parallels [Tót66, Bre84]. Parallelverschiebung [Rei92]. Parameter [Rid80, McH71, Moz16b]. Parameters [Buc07]. parapegma [vdW84a, vdW85a, vdW85b]. Parmenides [Ace00]. pars [Del16a]. Part [Fow82, Str81, Ano81, Bri10, Joh79, Joh81, Wil80a, Wil80b, Wim12, Wis81, BN78, Bri07, DJ07a, DJ07b, Gua15]. Partial [Kra03, Dem82, IM21]. particle [Bac89, Kra23]. particles [Hei68]. particular [Ber66, Mon72]. parties [Ras83]. parts [Pep68, Ras83, Ste22]. Pascal [Del20]. Pasch [Con76, Con76]. Paschal [McC96]. paschali [McC96]. Pascual [Bel85]. Pasquale [CD13]. path [Hen01]. pathological [Vol87]. pathologischen [Vol87]. paths [GM17, Van92]. Patterns [Whi61]. Paul [Arb79, NP04, NP06]. Pavia [Neu78]. Peacock [Fis99]. Peano [GP09, Gue84, Pal85, Seg94, Zai94]. Peirce [Kyr21, MP22]. Peletier [Mai90]. Pendulum [Hah02, Ari72]. penetration [Rei89]. Peninsula [CG15]. pentagonal [Bel10]. Perfect [Ace05]. period [Hof65]. periodi [Bot77a]. periodic [Dav90, Bot77a]. perpetual [vdW85a]. Perrin [Bru68]. Perron [Haw08a]. Perser [vdW60]. Persian [vdW60, vdW78b]. perspective [Arn84, Bri21, BPS11, Bro81]. Perturbations [Wil80a, Wil80b]. Petersburg [Dut88]. Petrović [DG20]. Pezzo [CD13, Gar89b]. Pfaff [Haw05]. pharmacology [GZKB^{+}12]. Phase [Jae98, McL11]. Phenomena [Nis05, Hon94, Nis11]. phenomenon [HH79]. Philolaus [Gra15]. Philosophical [Mer02, Spa91b]. Philosophiques [Mer02]. philosophischen [Spa91b]. Philosophy [Epp98, Mai94a]. physical [Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, Bra85, Dav86, Knu76, Lai85]. physicist [Arn83c]. physicists [Sch84]. Physics [Pye79, Arn84, Cor97a, GG85, Kra23, Pye82, Sch75, Seg80, She85, Zel09, Arn83a, Dos75, Sau99, Sau05]. Physik [Sch75]. Physiker [Sch84]. physique [Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84]. Picard [Ver20]. piece [Sch71]. Pieri [ABZ02]. Piero [Fie97]. Pietro [ED09]. Piezoelectric [Kat03]. Pincherle [Ros15]. pinhole [Lin70a, Man89, Man92, Str81, Thr94, Lin68, Lin70b]. Piola [CR07]. pioneers [RC23]. Pipe [Eck21]. Pisa [Gru17]. place [SS82, Vit02]. places [GM17]. plagiarize [Fre71]. Planck [Gio24, Hoy80, Kle61, Pet14]. Planck-type [Pet14]. plane [Gio21, Shk72]. Planes [Hah02, Ber76]. planet [dJ19b, dJ21, vdW60]. Planetarum [Man92]. Planetary [BJ00, Duk05a, Duk09, GC04, Moz16b, Swe99, Abr84a, Abr84b, CR19, Car20, CFH^{+}70, Gea85, GS08, Rid80, SM21, Thu92, dJ19a, dJ19b, dJH20, dJ21, vdW68]. Planetenbewegung [CFH^{+}70]. Planetenrechnung [vdW60]. Planets [Duk05b, Jon04, Van06, Gru17, dJ19a, CG13]. planning [Sar12]. Plate [IJ19a, IJ19b]. Plato [Ace00, Ace05]. Platz [SS82]. play [Hal88]. Plimpton [BPS11]. Plücker [Lor16]. pneumatics [Cha15]. Poîvre [Bri21]. Poincaré [And94, BG94, Dea82, JW20, JW22, Lüt94, Mil73, PB06, She91a, Ull00, Roq11]. point [Del22a, Kra12, Tay82, Vit82, vdW60, vdW76]. points [And94, McH71]. Poisson [Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, She78]. polarity [Che97]. pole [Hen97, dC16]. polemica [Mai90]. poles [Gue84]. Poli [Gue84]. Pólya [SS06]. Pólya-Controversy [SS06]. Polygons [DG20]. Polyhedra [Fie97]. Polynomials [Oak09, MR10]. Poncelet [Del16b, Del16c]. Poncelet’s [Del22a]. Popolazioni [FGP81]. population [FGP81]. porism [Del16b, Del16c]. Portable [Wri00]. position [Pai82, NB89]. possibility [EQ21]. possible [Bra85, Del22c, WH19]. possibly [Hol17]. Posthumous [Ste00a]. Postscript [Edw77, Wei80]. Postulate [vdW78a, Cor97b]. Postulates [vdW78a]. Postulato [Mai82]. Potential [Buc01, Net20]. Poussin [BN93]. power [RC23]. power-law [RC23]. Practica [Bus65a, Lit22]. practice [AT14, Gio24, Pal09, ZH24, vB17]. practices [MZ24]. Pre [Amb12, Ped00, Abr84a, Art85, Art88, Cou10, Fow80, Fow82, Rid80, Tho92, vdW80a, vdW80b, Car18]. pre-Babylonian [vdW80a, vdW80b]. pre-epicyclic [Abr84a]. Pre-Euclidean [Amb12, Art85, Art88, Tho92]. pre-Eudoxan [Fow80, Fow82]. Pre-History [Ped00, Cou10]. pre-Newtonian [Rid80]. Precession [Dob98, Dob99a, Not17, Swe80, Wil87]. precise [GB91]. precision [Hen97]. précurseur [Vin05]. precursor [Vin05]. predecessor [BY16]. prediction [GS09, Ste00c]. predictive [Gea85]. predictor [CE14]. preestablished [Pye82]. prehistory [Ger82, Hof66, She74, vB76]. Prejudicial [GG85]. Preliminary [Gol03, Pou98]. Prelude [Joh77]. Preludes [Kra12]. prepare [Gil74]. pressure [Cha15]. preuve [Voe08]. Preuves [Vit04]. Price [Hal90, Zab88]. primes [AM14, Edw75]. Primitivity [Neu06]. princess [Cal76]. Principal [FS05]. principe [Ker86]. Principia [Aok96, Pou92b, Aok92, Dob01, Nau11, Pou98, Pou03, Wei00, Pou09, DF20, Nau14, Pou20]. principle [Bel85, Cas92, Haw88, Pet14, Rei92, She93a, DJ07a, DJ07b, JS94, Ker86]. Principles [Dia99, KC00, Che15, Dar20, Fra83, Gio24, Loh61]. Prior [Nor04]. priority [Dut96a]. private [Fra98]. prize [BG94]. Probabilistic [Lov01, JS94, Sch75]. Probabilistico [FGP81]. probabilities [Sha78]. Probability [Sen98, SPJ01, She03, SS06, Bac90, BN78, Bra85, FGP81, GZ79, Hal90, Kyr21, Mit65, Sch74a, She71a, She71b, She73c, She74, She76, She77a, She78, She89b, She89a, She91a, She91b, She94a, She94c, vB76, She98]. Problem [Cen10, Sch74b, Tót66, BG94, BS84, BGTZ14, Dut85, Eck24, FT13, Gal80, GP09, Gre96, Hal88, Joh79, Joh81, Kal77, Kra97, Nau10, Pet61, Pou92a, RW12, Sab82, Scr61, Tak80, Thr94, Tup22, Vit02, ZH17, AG09, Alv99, Haw05, Spe96]. Problema [Gal80]. Probleme [Jah87, Vit02, BS84, GP09]. Problems [Dob98, Scr61, Car14, Cho10, Fri14, Han11, Jah87, LS86, Ses91, Sid18b]. Procedures [Man98a]. process [Don91, SS82]. processes [Bru68, Im96]. product [Dut82]. Production [Fra98]. profile [Dut81]. Program [Lac00, Arn83b, Cao91, Tru60]. programme [Bus19]. progression [AM14, Jaë20]. proiettiva [BN84]. Project [AB17, AB22]. projection [Lor95]. projections [Ace11a]. Projective [ABZ02, AB22, BN84, Hel22, SS21, Voe08]. Proof [Ace00, Dav03, SS86, Edw75, Edw77, Loh61, Sak82, Voe08, Vol89]. Proofs [Pal06, AB12, Sau05, Vit04, Wat80]. Prop. [Ace03a]. propagation [Eis91]. Proportion [Men07, Fow81, Mue70, Sza64]. Proportionen [Art85]. Proportionenlehre [Sza64]. proportions [Art85, Lam13, Tho92, Pal01]. Propos [Vit04]. Proposal [Kno78]. Proposition [Dia94, Dob01, Nau11, Pou03, Pou09]. Propositiones [Mai89]. Propositions [Ruf00, Mai89, Nau14]. proto [SV24, Ste20]. proto-history [SV24]. proto-Normal [Ste20]. Protons [vD04]. proves [Dia94]. Pseudaria [Ace08]. Pseudo [Mil17]. Pseudo-Aristotle [Mil17]. Ptolemaic [CR19, Car21, Moz16b]. Ptolemy [Ber91, Car09, Duk05b, Duk09, GC04, GM17, GJZ^{+}23, Jon99, Lor95, Mae98, Mar12, Moe74, Smi82, Van06]. Ptolemy’s [Mar21]. public [Pai71, Pai74, Pai77]. Publication [Har08, Bus19]. publications [Lor16, Vie21]. punctiform [Buc80b]. Punnett [Wim12]. punto [Vit82]. Purbachii [Mal16]. pure [Nab08]. purity [FP12]. pursuits [Duc11]. puzzle [Nau15]. Pythagorean [Pot74, Vin05]. Pythagoricien [Vin05].

qarastūn [Jao74]. Qin [Yiw20]. Qiujian [Lam97, Lam97]. Quadrate [RS98]. quadratic [Fri14]. quadratiques [Bou13]. Quadratur [Hof65]. Quadrature [Osa13, Hof65]. Quadratures [Bra03]. quadrilaterals [Pot74]. quaero [BN84, BN84]. qualitative [Roq11]. qualities [Syl71]. Quanta [NP04]. Quanten [ST94]. Quanten- [ST94]. quanti [Bas14]. quantifications [Syl71]. quantitative [GZKB^{+}12]. Quantities [Mer07, Fis81]. Quantization [NP04]. Quantum [Gea10, KC00, Lac00, Bro16, Cao91, Ehb22, Ehb23, FT13, GXB21, Han77, Im96, Kle61, Kox13, Pér09a, Pér09b, PS10, Roq92, ST94]. quasiconformal [Pap17]. Quellenstudien [Neu81]. Quelques [Dho86, Lam13]. quest [PR11]. questions [Uli15]. Quetelet [She86]. Quiescence [Gro98]. quintic [Ayo80, Lê17]. Quinto [Mai82]. quipus [AA72]. Quotient [Cou04]. Qurra [Jao74, Moe74].

R [Dav90, MP22]. R. [BF93, Pro88, She73c, Wil82]. Rabbi [Rab70]. Racah [Zel09]. Racine [Rās78]. Radially [Gro98]. Radially-Symmetric [Gro98]. Radiation [Hes00, BN88, Ehb22, Ehb23, KL84]. Radical [Pou92b]. Radici [RS98]. radii [GC24]. radio [Bra85]. radio-active [Bra85]. radioactivity [Kra97]. radix [Yaz20]. Rahmen [Sch74a]. Rainbow [Mer02]. random [Bru68, Dut85, She73a]. randomness [Yea11]. Rapporto [Mai94b]. rapprochement [Wil80a, Wil80b]. rate [Swe80]. ratio [Fow80, Fow81, Fow82]. rational [Tru60]. ratione [McC96]. Raumgeometrie [Art88]. Raumgruppen [Bur67]. rays [Ewa69, For69]. Re [Pal98]. Re-examining [Pal98]. Readable [Liu05]. Reading [Bul09, Ste01, Mas16, Gui16]. real [BS84, vR63]. Realität [ST94]. reality [ST94]. reason [Tru60]. reasoning [BHM15, Man98a]. Reassessment [Hah02, Ruf12]. reception [Bul09, Ehb22, Ehb23, Fis81, Rei03]. Recherches [Arb81, BS77]. reciprocal [Neu81, Str92]. reciprocals [MT97]. reciprocity [Col77]. reckoning [Kin73]. reconciling [Cao91]. reconsideration [Lin70a]. reconsidered [Dut93]. Reconstruction [Hüt00, vdW88c, BS22, Hen01, Str68]. Records [Liu05, SM21]. recouvrement [MT13]. recovered [GJZ^{+}23]. recreational [Han11]. recto [Gil75, Gil74]. Recurrences [BN88]. recursive [Fri14]. redaction [De 12]. Rediscovering [Fie97, Tru60]. rediscovery [SFS08]. reduced [GGHPMR21]. Reductio [Lov01]. Reductionism [Cor99]. réels [BS84]. Reference [Ace05, Ber66, Pot74]. Reflection [Jaë20]. reflections [Kno81]. Reflexione [Kal77]. reformulation [Wis81]. refraction [Bru81, Buc80a, CG20, Don21, Eas70, Går89a, Hen01, Smi82]. Refuses [Ear01]. regard [Mon72, RW12]. reggeization [Cao91]. Reggio [Mai94b]. Regine [Cha04]. Regiomontanus [Van18]. regola [BF93]. Regula [Con15, Con18]. Regular [Liu05, Edw75, Hog84, Wat72]. reign [Ste20]. Reihen [Bus65b]. Reinterpretation [Ruf63, DJ07a, DJ07b]. rejection [Mal91b]. rejoinder [Sak82]. Rekonstruktion [Hen01]. Relation [She98, Gil75, Osw17, Spe67]. relations [AA72, BS84, Sch88]. relationship [Bie82, Cha15, Hoy76, Sch84, Mai94b]. relationships [Hof60]. Relativistic [Kra92, Gio24, PR11, Kra81]. Relativitätstheorie [ST94]. relativité [Eis86, PB06]. Relativity [Mal00, Sau99, Cic15, Cor97b, Dar20, EG78, Eis86, Hen94, HG05a, Ill89, Kox88, PB06, Pye77, Pye82, ST94]. relevance [Nis09]. remainder [She88]. remark [ENP82]. Remarks [Ace03a, Bri01, Sei66, Amb12, vB76, Lau65, RDV01]. Remarques [RDV01]. Renaissance [Fra13, Mal10, Pal08, SFS08]. René [Dug76]. renewed [Del16b, Del16c]. Replication [Mar06]. Report [Gol03, Jon99, Sch86]. Representation [Cap02, Bos81, Gra84, Haw72, Haw88, Hei70]. representations [Dut86, Hel22]. representing [Cra13]. research [Arb81, BS77, BN88, Duc11, Mal16, Par88, Pet61]. researches [RD94, Spe67]. reserves [Sar12]. resilience [BY16]. resistance [Aub17]. resisting [Ait72]. Resolution [Gar89b, Bel91, Rās74, Rās74]. resonance [Bis22, Kat08, Kra10]. respect [Arn83d]. Response [BG09]. Rest [Dia99]. restoration [Del20]. Result [Net99, Dal86]. Results [Fra98, Twe91b]. retrograde [GC23]. revalidation [Loh66]. Revalidierung [Loh66]. revaluation [Del08]. review [BPS11, Ruf12]. revised [Kno78]. revisited [IJ19a, IJ19b, Moz24a, Nau11, Ruf13]. Revisiting [BHM15]. Revolution [HG05b, Arn83a, Che97, Dar93, Oak18]. rewrite [Ung75]. rewriting [Lam89, Lam88]. Reymond [Fis81]. Rezeption [Rei03]. Rham [Kat85]. Rhind [Gil74]. Ribaucour [Rou80]. Ricci [CT24, Mal91b]. Richard [Sch82a, Sch82a]. Riemann [Neu81, Rei73, Bot77b, Neu81, Rei73, Sch82b, Wei79, Wei80]. Riemanns [Bot77b, Sch82b]. Riesz [Gra84, Rod15]. Right [LS84]. Right-angled [LS84]. Rigid [Dob98, Mac23, Wil87]. rigidification [Cas92]. rigorise [Lüt79]. rigour [Smi86]. Rings [Cou04]. Rise [Ehr06, Fer99, Bos84, Kop71]. Risings [Swe99, Sch23]. ritual [Sei61, Sei62, SC80, Sei81]. RMP [Gil75]. Robd́ [Ste00a]. Robert [AG09, Cha15, Dut90, Twe91a, Twe91b]. Robinson [BS85]. rod [Lam87]. Rodrigues [Gra80]. rods [Bel91, Dil92]. Roger [Lin70a]. Role [Har03, SS09, Yav01, vD09, CGM87, Hoy73, IR16, Nis11, RA20, Seg80, War91]. Rolle [Hoy73]. Romaka [Abr86, vdW85b, vdW88b]. Roman [DEHY76, Wri00]. Rome [De 12]. Roomen [Boc09]. Root [AH15, Rās78, Sei76]. Roots [Ehr06, Dut86, Mil83, RS98, Roc21, Wat78]. Rosen [Blu22]. rotating [Lüt84a]. rotation [CT24]. Rotational [Dob99b]. Rotundorum [Kal77]. Rounding [Car09]. Rowan [Mat78]. Rudolph [Sch75]. Ruffini [Ayo80]. Rugeley [Dav86]. Ruggero [BN84]. rule [BF93, Mal93, Mil17]. ruled [Del11, GP09]. rules [Pot74, ZH24]. Russian [Sen98]. Rutherford [Hei68, Hes00].

S [Arb79, Bur71, Bur72, She76, Cao91]. S-matrix [Cao91]. S. [She78]. sa [Voe08]. Saccheri [BN84, Mai82]. Saint [Man92]. Sake [Bra03]. Salvatore [Ros15]. Samarqand [Moz16b]. Samaw’al [BHM15, Ras72, Wat78]. sambuke [Key22]. Sanssouci [Eck02]. Saraf [Rās74, Rās74]. Saraf-al-din [Rās74, Rās74]. satellites [Eis96]. satisfaction [Mac23]. Saturn [Abr82, Car20, Car21, Coh98, SdJ23, Wil85]. Saturni [Car20]. Saving [Smi81, Smi82]. say [Dia94]. scalar [Nor92]. scale [Vig00]. scales [Sma11]. scattering [Hei68, Roq92, Sch74a]. Scheme [Mer07, IJ19a, IJ19b, SdJ23]. schemes [Jon83, Ste22]. Schmidt [SS86, SS86]. Schmidt-Theorems [SS86]. Schoenflies [Bur71, Bur71]. School [Syl71, Neu81, Par88, Arb79, Art88]. schools [Uli15]. Schreier [BS85, Joh84]. Schrödinger [Han77, MR10]. Schrödinger’s [RP23]. Schule [Art88, Neu81]. Schwarzschild [Eis82, Eis87]. Science [BG09, Mal00, Cha62, Mit65, Ron60, Sch82c, Wis74, ZH17]. Sciences [BG08]. Scientific [Sch88, Cal76, Fre74, Mer02, Zab88]. Scientifique [Mer02]. Scommesse [FGP81]. scope [BG08]. screen [Arn83f]. scribe [Gil74]. Search [Pal06, Bec73, Bie92, Smi82]. Searches [Sch17]. Second [Duk07, Neu06, Bie81a, Bie81b, Bie82, Bie83b, Bie87, Blu22, Dar20, Hof65, Lam13, Pou20, Pou06]. second-biggest [Blu22]. seconde [Lam13]. section [Scr61]. sections [DF18, Del20, LA84, Sma11, Twe91b]. secular [Kus89]. Seeberg [Her70, Her70]. seem [Hol14]. segment [Men23]. Segner [Bik78]. segni [BF93]. Segre [Gar89b, Gar91]. seicento [BN84]. Seidenberg [Mat85]. seine [Sch82c, Vol89]. seinen [Sch88]. seiner [Bie82, Sch82a]. seines [Hen01]. selection [Sar12]. Selectivity [Fra98]. selenelion [Moz24a]. self [GL12]. self-sustained [GL12]. semi [Mac23, Sei72]. semi-circle [Sei72]. semi-rigid [Mac23]. semigroups [Hol09, Hol14]. seminar [Pye79]. Semivectors [vD04]. Sense [Mit19b, Mit19a, Tup22]. sent [LD01]. separation [Big83]. septième [BS84, Lam13]. sequences [BC20]. Series [Fer99, Bus65b, Coo93, Dut82, Dut96b, Mal93, Osw17, RD94, Tuc73, Twe91a, dO88]. Servois [Del22b]. ses [BS77]. set [Coo93, Tay82]. sets [Dau71, Fer96]. settings [Sch23]. Seven [CG13]. seventeenth [Dos75, Sha73, Web65, Whi61]. Severi [vdW71]. sexagesimal [Fri19]. shadow [GC24, Pye79]. SHAKE [Mac23]. shape [Gre96]. shaping [Gra84]. sharp [Kat08]. sheaf [Cho10]. Shen [Gua10]. Shih [Lam79]. Shih-chieh [Lam79]. ship [Fer14]. shock [RKH^{+}21]. Shocking [vdW76]. short [She83]. Shoulders [Ace03b]. Shu [Dau08a, Dau08b, Dau08a, Dau08b]. Sicht [Jah87]. Siddhānta [Abr80, Abr86, vdW85b, vdW88b]. Sides [Fow82]. Siècle [Vig00, Lam13]. Siècles [Rās78, Ras83]. Sierpinski [Arb81]. significance [Duc11, Sch82c]. significato [Pal85]. Signorini [SV24]. signs [BF93]. simbolismo [Pal85]. Siméon [Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84]. similarities [Lam87]. Simple [Sil79, Hol14, Nis11, Ste24b]. simplicity [Ruf13]. simplifies [CT24]. Simson [Twe91a, Twe91b]. simultaneous [Sch23]. Sine [MB03, Don21, Loh61]. sines [WH19]. Singolarità [Gar91]. Singularités [Eis82]. singularities [Eis82, Gar89b]. Singularity [Gar91]. six [GGHPMR21]. Sixteenth [Cha04]. sixth [BN84]. sixty [Sei65]. size [Car17, Lau65, PR11, RC23]. Sizes [BS07, Cio23a, Car14, Was71]. small [FHR21]. snare [Ruf13]. Snellius [Hen01, Hen01]. so-called [Wil70]. Society [Sch88]. sociology [Fis66]. Solar [Bri07, Liu05, MS15b, EC19, Moe74, Moz16a, Moz18a, Moz22, Wil80a, Wil80b, Yaz11a, Yaz11b]. Solid [HG05b]. solids [BS22, SFS08, Wat72]. Solution [Eis82, Eis87, Gal01, Nau15, Kal77, Pou92a, Scr61, Thr94]. solutions [BN84, Hal88, Sch77, Ses85]. soluzioni [BN84]. solvable [GS94, Kra97]. Solving [Cen10, BC20, CPFA07, Sab82, ZH17]. Some [Kno81, McG66, Twe91b, vB76, Amb12, Dut96b, Arn84, Dho86, Fra89, Lam13]. someone [Dia94]. Sommerfeld [Sch85, Sch85]. son [Dah80]. Sophie [Sam90, DF12, Del08]. Sophistic [Ber03]. sophistication [Høy08]. Sophus [Kay23]. sopra [Gar91]. Source [BJ00, Jon99, Neu81, RR78]. sources [Del22c, Lit22, McG66]. Souslin [Alv99]. Soviet [Arb79]. Soviétique [Arb79]. sowie [Hen01]. Space [Epp98, Bur67, Cog15, PR23]. spaces [Ber66, Jaë20]. Spanish [Mar93]. Spatial [Art88]. Special [Ace03a, Pot74, RW12, Pye77]. Specific [Gea10]. specious [GGHPMR21]. spectrometry [Hen97]. speculations [Kra12]. speculativa [Bus98, Bus98]. sphere [Sei88, NB84]. Spheres [Yav98, Rid79]. spherical [Raw82]. spherics [NTP04b, NTP04a, SS09, Mal10]. Sphériques [NTP04b, NTP04a]. Spiegel [Sza64]. Spirited [Dar03]. Spitzer [Dea81]. splines [BSS88]. splittings [Con18]. spring [Cha15]. Spurious [Ber76]. spyglass [ZH12]. Square [Wei00, Aok92, Aok96, Dut86, Nau10, RS98, Sei81, Vin05]. Squares [Ses03, Dut90, Dut96a, Mag22, MT97, Ses98, She93a, Wat90, Wim12]. St. [Dut88]. Staatsstelle [Pai71, Pai74, Pai77]. Stabilitätsbetrachtungen [Hoy73]. Stability [Roq11, Hoy73]. Stable [Hes00]. stades [Raw82]. stadion [Gul87, Tup22]. Standard [Mer07, BS85]. standards [Kat08]. standing [Nau15]. Star [Duk02a, GS08, RSW04, Ste20, Mar21]. Stars [Jon04, Man04, NB89, Vig00, Cen10]. start [Yea11]. state [CM15]. stated [JW20]. statement [BB97]. static [Jao74]. Stating [Net99]. stations [SM21]. statique [Jao74, Ker86]. Statistica [FGP81]. Statistical [Nav98, Bru67, PS10, Ren97, She80, She84a, She84b, She85, vB76]. statistician [She86]. Statistics [Sen98, She98, Ald09, FGP81, Han77, PS10, She82]. Staudt [Har08, Nab08]. Steam [Key92]. Stech [Str92, Str92]. Steiner [BGTZ14, Lor16]. Stellar [Cen10, Van18, Van21]. stemma [Ray14]. Step [Men07]. Stephen [Gru17]. Steps [Ruf00]. stereographic [Lor95]. stereometric [Neu74]. stereometrischen [Neu74]. stetiger [Vol89]. Stevin [AH15]. Stirling [Kri68, Twe92]. Stirlings [Kri68]. Stokes [Dar02]. Stonehenge [Kai11]. story [Del16b, Del16c]. Strabo [Raw82]. strange [Loh68]. strings [Fri20]. Structure [Kra03, Haw82, Ler72, Par85, Wim12]. structurels [Jaë10]. structures [Cho10, Kai11]. struggle [Kox13]. Stück [Sch71]. Studien [Kno74]. Studies [Bri07, Bri09, Bri10, GS08, GS09, Pal01, Kno74, Lam79, Mer02, Neu81]. Study [Dah80, Dal01, Jao74, Jon04, Asc88, Dem82, Dia94, Fis66, Hol17, Mae88, Mil73, Pot74, Sch73, Sha73, Smi82, SM21, Tay82, Tay85, Tay87, Wis74, dJ19a, dJ19b, dJ21]. Sturm [Lüt84b]. Suan [Yon69, Dau08a, Dau08b, Lam79]. Suan-hsüeh [Lam79]. Suanjing [Lam97]. suanshu [Lam94]. sub [Eva76]. sub-Euclidean [Eva76]. subgroup [Joh84]. subsequent [BY16]. Substitutions [Dah80, Vit04]. Successes [Gea10]. Successful [Dav03, Bie87]. suggested [Bec73]. sul [Gue84]. sulla [FGP81]. Sum [Fer99, Brå07, Spa02]. Sumer [Sei65]. summable [Tuc73]. summary [Bel10]. Summation [Mil85, Dut96b, MT97]. Summensatz [Spa02]. sums [Lau89, She73a]. Sun [BS07, Cio23a, Abr80, Car14, Kin73, Too74, Abr85, Bro05, Car09, Coh98, GC24, Mae98, vdW82, vdW85b]. Sundial [Hüt00]. Sundials [Wri00]. Sunrise [Lov01]. super [Mal16]. superconductivity [CGM87, GG85, Sau07]. superficie [Gar91]. Superposition [Pal09]. supplementing [Hen01]. supply [Raw82]. surface [GP09, Gar91]. surfaces [Gar89b]. surrounding [Kus89, Rei03]. survey [McH71]. surveying [LS86]. surviving [Raw82]. suspension [Nau15]. sustained [GL12]. SWAC [Cor08]. Swings [Hah02]. syllogistic [Smi78]. Sylow [Wat80]. Sylvester [Par88, SPJ01]. symbiosis [Gol70]. Symbole [KSFW90]. symbolic [Lam86, Pet87]. symbolical [Haw08b]. Symmetric [Gro98]. Symmetry [HG05a, HG05b]. syncategorematic [Art13, RA20]. synthesis [Arn83e]. Synthetic [Lor16, Del22b]. System [Bri01, Duk02b, Shn17, Del16a, Lam86, Lam87, Lam88, Lam89, Sei65, SdJ23, Sza60, Yiw20, HA98]. systemate [McG66]. Systems [Ehr06, Kje02, Dea88, FT13, IM21, RW12, ZH24]. Szilys [Bie83b, Bie83b].

Tābit [Jao74]. Tūsī [Rās74]. Tabelle [FGP81]. Table [Liu05, Man98b, MB03, BHM15, Dav90, Yaz18, vdW88c]. Tables [GC04, GC24, MB03, Shi03, Ste06, Abr84b, Cha91, CG13, CG19, Che15, FGP81, GC17, GC18, Kin73, McC96, Wil80a, Wil80b, vdW85a, CG13, Not18]. tablets [Tum92]. Tabulae [Cha16, Her63]. tabular [MZ24]. tabulation [Van18]. Tabule [Cha04]. Tachyon [Frö94]. take [Fra98]. Taking [Van06, Fra98]. Tale [Men07]. tangent [Scr61, Twe91a]. Tangentenaufgaben [Scr61]. Tangenti [Gal80]. tangents [Gal80, Mal93, Pep68, Scr63, Str68]. tardo [BN84]. Taylor [Fei85, Fow91, Mal93]. teacher [Sma11]. teaches [Ill89]. teachings [Lau65, Sza64]. technical [Car20]. technics [Fra13]. technique [Fow91]. Technology [BG09]. Teichmüller [JP13]. Teilchen [Bac89]. telescope [ZH12, ZH17]. Telescopes [Ped00]. témoins [Vin05]. temperature [GG85]. temperatures [Tru79]. ten [Sei76]. tensor [Mal91b, Roc21]. tensorial [Mal91b]. tentative [Tum92]. Teoria [FGP81]. term [Duc11, Ger82, Pet63, Ras71, Sch82b]. terminologia [Vit82]. terminological [Kra23]. Terminologie [Sza64]. terminology [Sei84, Sza64, Vit82]. terms [Bol72]. test [Aok92, Aok96]. tested [Eis96]. Testing [Duc11, Hen94]. Text [Bri01, HA98, HY86, RSW04, Ste00b, Kno86, RDV01, Tum92, Bre94b]. Texte [Bus98, RDV01]. Texts [BS07, GS08, Bus98, Fri19, Sme70]. textual [FG23, Vit02]. Textuelle [Vit02]. Textzeugen [Bre94b]. Thabit [Moe74]. Theaetet [Pai82]. Theaetetsstelle [Pai82]. Theaetetus [Ace05]. their [Ber66, Bie85, Cha15, Che15, Del22c, Dut86, Fis81, Hof66, Ill89, Nem62, RA20, Seg94, Sza64, vdW85a]. theme [Cog15]. Theodosius [Mal10, SS09]. Theorem [FS05, Gil91, MT13, Voe08, AB12, AM14, Bel10, Brå07, Dia94, DPC95, Edw75, Gra84, Gug77, Haw08a, Joh84, Kno92, She88, Spa02, Wat80, Bie82, Dal82, Del08, Kje01, Ras80, SS86, Ull00]. théorème [Gil91, MT13, Ras80, Voe08]. theorems [Ber76, Sid18b, SS86]. Theoretical [Sha79, Bul09, Bus98, Hof60, Zel09]. theoretischen [Bus98]. Theorica [Bus98, Bus98]. Theoricas [Mal16]. Theorie [HB86, Hoy74, Hoy76, Neu97, Ull89, vdW72a, Arb81, BS84, BS85, Eis91, Gil91, Lam13, Nab08]. Theories [Abr85, Cen10, CT24, Jen72, Lai86, Moz18a, Nor92, Rid80, Tru75, Ver13a, Ver13b]. theorist [Ehb22, Ehb23]. Theory [Arb81, Bri07, Dar02, FS07, Gea10, Hoy73, Kje02, Kox88, Kra81, Man04, Men07, Neu06, Pye77, Swe99, vdW72a, Abr84a, Ald09, AT14, Arn83a, Arn83b, Arn83c, Arn83d, Arn83f, Arn83g, Arn83h, Arn83i, Arn83e, Arn84, BS84, Ber66, Bie81b, Bie81c, Bie82, Bis22, Bos84, Bri09, Bri10, Bro67, Bru73, Bru74, Buc80b, BY16, Cao91, Car86, CGM87, Cog11, Cog14, Coo93, Cor08, Dar93, Dau71, Dea88, Del22b, Dem83, Dil92, Dos75, Edw80, Edw90, Ehb22, Ehb23, Eis91, EC19, FGP81, FP12, Fer14, Fis66, Fow80, Fow82, GG85, Gil91, Gio21, Gol70, Gru17, Gua15, Ham76, Haw71, Haw72, Haw74, Haw77, HB86, Hol09, Hom67, Hoy74, Im96, JP13, Joh77, Kie71, Kle61, Kno74, Knu76, Kox13, Lam13, Lê17, Ler72, Lin68, Lin70a]. theory [Lin70b, Lüt84b, Man89, McC66, Moe74, Mue70, Muk74, Nab08, Net20, Neu97, Pap17, Par85, Pér09a, Pér09b, PS10, Pet14, Pye79, Rab74, RP23, Ros71, SV24, Sal74, Sam90, Sch86, ST94, Ses91, Sha73, She71b, She74, She77a, She77b, She79, She91b, She94c, She95a, She95b, Shn17, Spa91a, Spe67, Str81, Tay82, Taz93, Tho92, Tuc73, Ull89, War91, Wil80a, Wil80b, Zan13, Zan14, Zeh83, dJ19a, dJ19b, dJ21, vR63, vdW68, Bie81a, Haw91, Hen94, Hoy76, Mal91b, Pal98, Pal01, Pal05, Sau07, She98]. There [Ace19]. Thereon [Wei00]. thermodynamics [Bie83b, Bie83a, Bie87, Bie92, Dai21, Hoy74, Hoy76, Smi77]. Thermodynamik [Bie83a, Bie87, Bie92, Bie83b, Hoy74, Hoy76]. third [Gul87]. thirteenth [Lin68, Not24, Ung74]. thirty [Tru96]. thirty-fifth [Tru96]. Thomas [Bus98, Loh66, Big19, Bus98, Dal86, Gou14, Kal77, Loh66, Loh79, Ste00a, Ste08, Zan13]. Thomson [Bie92, Bie92, Smi77, Top71, Wis81]. though [DF18]. thought [Jaë20, Mal10, Whi61]. thoughts [Rit15]. thousand [Fri19]. Threat [BG09, Hen97]. Three [Fri19, BG94, EQ21]. Thyrrheion [Hüt00]. Tides [Pal98]. time [Kin73, MZ24]. times [Ses91]. Tishboret [FG23]. Tissot [Pap17]. tool [Ace11b, Con18, Ehb22, Ehb23, Kje08]. Topological [Epp98]. Topologie [Arb81, Neu97]. Topologique [Arb79]. Topology [Arb79, Arb81, Epp98, Ull00, Wei80, JW20, Joh79, Joh81, Neu97, Rod15, Tay85, Wei79]. Torsion [Mar06]. tra [Bot77a]. trace [Høy08, Ray20]. Traces [Men07, Hog86]. tracing [Oss18]. tradition [CG15, Mat85, Men23]. Traditional [Fer96]. traditions [Cor13, Moz16b]. traité [Arn83e]. Traitement [Ses77]. traités [Lam13]. Trajectoires [Eis87]. trajectories [FT13, Gui16, Roq11, Eis87]. transfer [Haw88]. transform [Dea81, Dea92, Dea82, HJB85]. Transformation [Str92, Cog14, Gra80, Lor95]. Transformations [PB06, Cou10, FN23, Kay23, Shk72]. Transformed [Net99]. transforms [Dea85]. Transition [Ruf00]. Transitions [Jae98]. Translation [Dau08a, Dau08b, Gui16, Hen01, RGM19, Rei92]. transmission [FG23, Jon83, Mal10, Rei92, WH19]. transport [IR16]. transversal [Del22c]. transversals [Del22b]. trapezoid [Oss18]. Traum [Car86]. Traum- [Car86]. travail [Jao74]. Travaux [Dah80]. traversale [AB19]. treatise [Car14, DF18, Del20, Gol77, GJZ^{+}23, Van21, Fis99]. treatises [Lam13, Ste08]. Treatment [Dob98, Dob99a, Duk05b, Arn83d, GC23, Kau85, Ses77, She72, She73b, She93b]. Treatments [Bri09]. tree [BGTZ14]. Trepidation [Man98b, Not17]. triads [Pot74]. triangles [LS84]. triangulorum [BS77]. Trigonometric [Che15, Coo93, Dau71]. trois [Jaë10]. tropical [Swe80]. trous [Eis91]. True [Boo85, Not18]. Truly [Bie07]. Tschirnhaus [SS05]. Tubes [Yea11]. turbulence [Eck21]. Turkey [YT18]. turn [ZH12]. turning [McH71]. Tusi [Rās74]. twelfth [Eva76, Not24]. Twentieth [Wei00, Bel91, Hol17, Tay74]. Twentieth-Century [Wei00, Hol17]. twenty [Cho11]. twilight [Gol77]. twixt [Wes64]. Two [Car14, Men07, Pou04, Sal74, Ses91, DJ07b, Lüt94, Moz24a, Neu81, Sch84, Wes64, Bus98, EG78]. Two-Step [Men07]. Tycho [Car20, Mae98, Ros81, Str81]. type [Pet14, SdJ23]. Types [Arb81].

U.S. [Bul09]. Überlieferung [Bre94b]. Übersetzung [Hen01]. Übersicht [Neu81]. Übertragungsprinzip [Rei92]. UCLA [BS22]. ultrasonic [Kat08]. Umdeutung [DJ07a, DJ07b]. Umfeld [Rei03]. Umwälzungsprozeß [SS82]. Unbekannte [Kno74]. uncertainty [Pet11]. Underestimated [GH00]. understanding [Bas14, BC12, BB20, Nis11]. understood [Pou20]. Unendliche [Bus65b]. unflattenable [GP09]. Unification [vD04]. unified [CT24]. uniform [Bri10, HB86, Not17, Vie21]. unique [dJH20]. Uniqueness [Coo93]. unit [Gul87]. universal [Aok92, Aok96, Duc11, Str92, Wil70]. Universalis [Boc09]. universe [PR11]. universellen [Str92]. Universität [HY86]. University [HY86, Par88]. Unknown [Høy08, Kno74]. Unpublished [Del08, Gug77, RKH^{+}21, Sch82a]. unsolved [Kra97]. untapped [RR78]. Untergruppensatz [Joh84]. until [Bus19]. unum [Bus65b]. Unusual [Ste00b]. Unveröffentlichte [Sch82a]. unwelcome [Hen97]. upper [Sch73]. uranium [dAM97]. Urbino [Fra13]. Uruk [Ste00b]. Urysohn [Arb79]. usages [MS15a]. use [Ace11b, Fow91, Man92, RA20]. used [Gul87]. useful [Kin73]. Uses [Sid18b]. Using [Ses03].

v [Bie83a, Arn83f, DF20, Str92, vdW85b]. V-A-reciprocal [Str92]. V-A-Wechselwirkung [Str92]. V-VII [DF20]. V. [She91b]. vacuum [Kra12]. Vallée [BN93]. Vallejo [CPFA07]. Value [Bel98, Car21, Gul87, Sch71, Sme70]. values [Lau89, Moz16b]. Variational [Fra94]. Variations [Cog15, Fra85]. Vasistha [Abr80]. ve [FG23]. ve-ha-Tishboret [FG23]. Vector [Cap02, Rei89, Roc21]. Vegard [Dav90]. Vektorkalküls [Rei89]. Velocities [Mae06, Gra65]. Velocity [Cap02, Gua15, Sak82]. Ventimiglia [BN84]. Venus [Bri01, HA98, Rid78, dJ19b, vdW82]. verge [DJ07a, DJ07b]. Verhältnis [Hoy76, Sch84]. vermied [Art85]. Versagen [Bie85]. version [BS22, Hen01, Rec94]. Versuch [Bie83b]. Versuche [Bie87, Bie92]. versus [Bro16, Pet14, Roc21, Wil87]. very [Kin73]. VI [Arn83g, Sid18b]. via [Spa02]. vibration [Dos75]. vicaria [Mae90]. Victorian [Mit19a, Mit19b]. vie [Dug76]. vier [Neu73b]. Viète [BS77, Bie07, Boc09, Oak18, Rās74, Ste08]. view [Cor13, Jah87, Kno86, vdW76]. views [Arn83g, Ros69, LM98]. VII [Arn83h, DF20]. VIII [Arn83i]. Vinci [Fie97, Thr94]. vindication [Bre94a]. Virial [Bie82]. Virialtheorem [Bie82]. virtual [IR16]. vis [vB17]. visibility [Yaz18]. vision [Gio24, Gou18]. Vita [FGP81]. Vitali [Cog22, MT13]. vitriol [Sch74b]. Vitriolen [Sch74b]. viva} [vB17]. Vleck [DJ07a, DJ07b]. void [Gra65, McG66]. Voigt [CRT10]. Vollständigkeit [ST94]. Voltaic [Hon94]. Volterra [Ull00, Gue84]. volume [Sei88]. vor [Joh84, Sch74b]. voreuklidische [Art85, Art88]. Vorgeschichte [Ger82, Hof66]. Vorlesung [ST94, Ull89]. vs [Cor08].

W [SPJ01]. W. [Arb81]. Waerden [Mat85, vdW72b]. Waerden/Seidenberg [Mat85]. Wahrscheinlichkeitsrechnung [Sch74a]. wahrscheinlichkeitstheoretischer [Sch75]. Wallis [Dut82, Fow91, Ste01]. Walter [Kat08, Zan14]. Waqar [GC23]. waren [Bie87]. Wärmeäquivalent [Hoy78]. Wärmelehre [Bie81a, Bie81b, Bie81c, Bie82]. was [Car18, Car86]. Water [Dar03, Ped00, Eis86]. Water-Filled [Ped00]. Waterston [Hoy78]. Waterstons [Hoy78]. wave [Bac89, Blu22, Ger69, Sha73]. Waves [Dar03, BYS^{+}21a, BYS^{+}21b, Don91, Kub70, RKH^{+}21]. Weber [vdW72a, vdW72a]. Wechselwirkung [Str92]. Wechselwirkungen [Neu81]. Wedderburn [Par85]. wedge [vB17]. Weierstraß [Neu81, Dug73, Haw77, Neu81, Spa91b, Ull89, Vie21, Ull89]. Weierstrassian [Man75]. Weierstraßschen [Spa91b]. Weil [vdW71]. well [Hen01]. Welle [Bac89]. Wellenmechanik [Ger69]. were [Jaë20, Bie87]. West [Ung74]. Western [MB03]. Weyl [Ker86]. Whaples [DS12]. Whittaker [Cou10, Cou14, FN23]. Who [Wei78]. whose [Art85]. widespread [Sei76]. Wie [Bie87]. Wiechert [Sch82c, Sch84, Sch82c, Sch84, Sch85]. Wigandus [Bus98, Bus98]. Wilbraham [HH79]. Wilhelm [Haw82]. Wilhelmian [Pye82]. Will [Dia94]. William [McC66, AG09, Man92, Mat78, Smi77, Wis81]. Wilson [Ras80]. Wirken [Sch73]. Wissenschaft [Mit65, Sch82c]. Wissenschaften [Sch88]. Witelos [Loh68, Loh68]. witness [Bre94b]. witnesses [Vin05]. Wonders [Ste01]. Woodin [Rit15]. Work [She03, Bel10, Bul09, BN93, CPFA07, Del08, DF18, EQ21, Fen98, Fer07, Ham76, IR16, Jao74, Kra92, Lüt84a, Lüt84b, OS14, Ros15, Rou80, Sch82a, Sha78, She71a, She73c, She76, She78, She89b, She89a, She91a, She91b, She94a, She95b, Tay82, Tay85, Taz93, Twe92, Wil82, vdW88a, SS05]. works [Dah80, Dug76, ED09, Hog86, JW22, Mar93, Sch73]. workshop [AB19]. worrying [Yea11]. writing [Yiw20]. writings [CM15, GL19]. wurde [Car86]. Würfe [Nab08].

X [Ewa69, Arn84, For69]. X-rays [Ewa69, For69]. XI [Rās78]. XII [Rās78]. XIII [Ras83]. XIV [LD01, Ras83].

Yūsuf [Bre94b]. Ya [She91b]. Yakov [Fre74]. Yan [LA84]. Yang [Yon69]. Ye [LA84]. Year [Bel98, Abr84a, Bie81a, Bie81b, Bie81c, GS08, GS09, Her70, Not18, Swe80, vdW78b]. years [Cho11, Fri19, HB86, Joh84, Sch82a, Top71]. Yi [Gua10, LA84]. Ying [Gua10]. Young [Bik78]. youthful [GL19]. Yunus [Kin73].

Zîj [CG94]. Zacut [GC18]. Zahlbegriffs [Spa91b]. Zahlen [Spa91a]. zahlentheoretische [Hof60]. Zahlentheorie [Sch86]. Zaragosa [Rec94]. Zarqalluh [Moz24b]. Zeit [Sch82a]. zero [Kra12]. zero-point [Kra12]. zeta [Dut96b]. zhang [Lam94, Lam97]. Zij [Gol03, CG15, Moz22]. Zipf [RC23]. Ziqpu [RSW04]. zodiac [Bri10]. zum [Bec67, Bre94b, Sch82b, Pet63]. Zur [Pet61, Bie83a, Bus98, Hoy76, Kno74, Loh66, Neu73a, Sch75, Sch82c, Str92, Ull89, Vol87, Was71, Bur67, CFH^{+}70, Ger82, Ras71, Sch86, Scr61, Vol89]. Zusammenhang [Hoy74, Rei92]. Zusammenhänge [Hof60]. zwei [Neu81, Bus98]. zweiten [Bie81a, Bie81b, Bie82, Bie83b, Bie87, Hof65]. zwischen [Bre94b, Neu81, Neu97, Spa91b, vdW72b].

References

Ascher:1972:NRA

[AA72] Marcia Ascher and Robert Ascher. Numbers and relations from ancient Andean quipus. Archive for History of Exact Sciences, 8(4):288–320, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=8&issue=4&spage=288.

Achilles:2012:EPT

[AB12] Rüdiger Achilles and Andrea Bonfiglioli. The early proofs of the theorem of Campbell, Baker, Hausdorff, and Dynkin. Archive for History of Exact Sciences, 66(3):295–358, May 2012. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=66&issue=3&spage=295.

Anglade:2017:NIB

[AB17] Marie Anglade and Jean-Yves Briend. La notion d’involution dans le Brouillon Project de Girard Desargues. (French) [The notion of involution in the Brouillon Project of Girard Desargues]. Archive for History of Exact Sciences, 71(6):543–588, November 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-017-0196-5.

Anglade:2019:DTA

[AB19] Marie Anglade and Jean-Yves Briend. Le diamètre et la traversale: dans l’atelier de Girard Desargues. (French) [The diameter and the cross: in the workshop of Girard Desargues]. Archive for History of Exact Sciences, 73(4):385–426, July 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Anglade:2022:ENB

[AB22] Marie Anglade and Jean-Yves Briend. Nombrils, bruslans, autrement foyerz: la géométrie projective en action dans le brouillon project de Girard Desargues. Archive for History of Exact Sciences, 76(2):173–206, March 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://

link.springer.com/article/10.1007/s00407-021-00282-3.

Aboav:2008:EBD

[Abo08] David Aboav. Euclid’s book on divisions of figures: a conjecture as to its origin. Archive for History of Exact Sciences, 62(6):603–612, November 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=6&spage=603.

Abraham:1980:MSV

[Abr80] George Abraham. The motion of the sun in the Vasis.t.ha Siddhānta. Archive for History of Exact Sciences, 22(1–2):1– 3, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
22&issue=1&spage=1.

Abraham:1982:AFM

[Abr82] George Abraham. Algebraic formulae for the Moon, Saturn and Jupiter in the Pancasiddhāntikā. Archive for History of Exact Sciences, 26(4):287–297, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=26&issue=4&spage=287.

Abraham:1984:LYP

[Abr84a] George Abraham. Length of the year in pre-epicyclic planetary theory. Archive for History of Exact Sciences, 30(3–4):189–195, September 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
30&issue=3&spage=189.

Abraham:1984:MME

[Abr84b] George Abraham. The motion of Mars in Egyptian planetary tables. Archive for History of Exact Sciences, 30(1):1– 6, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
30&issue=1&spage=1.

Abraham:1985:TMS

[Abr85] George Abraham. Theories for the motion of the Sun in the Pancasiddhāntikā. Archive for History of Exact Sciences, 34(3):221–230, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=34&issue=3&spage=221.

Abraham:1986:MMR

[Abr86] George Abraham. The motion of the Moon in the Romaka Siddhānta. Archive for History of Exact Sciences, 35(4):325–328, December 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=4&spage=325.

Avellone:2002:FPG

[ABZ02] Maurizio Avellone, Aldo Brigaglia, and Carmela Zappulla. The foundations of projective geometry in Italy from De Paolis to Pieri. Archive for History of Exact Sciences, 56(5):363– 425, July 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=5&spage=363.

Acerbi:2000:PPC

[Ace00] F. Acerbi. Plato: Parmenides 149a7-c3. A proof by complete induction? Archive for History of Exact Sciences, 55(1):57– 76, August 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=1&spage=57.

Acerbi:2003:DMR

[Ace03a] F. Acerbi. Drowning by multiples remarks on the Fifth Book of Euclid’s Elements, with special emphasis on Prop.

8. Archive for History of Exact Sciences, 57(3):175–242, April 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=3&spage=175.

Acerbi:2003:SH

[Ace03b] F. Acerbi. On the Shoulders of Hipparchus. Archive for History of Exact Sciences, 57(6):465–502, September 2003. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=57&issue=6&spage=465.

Acerbi:2005:RPN

[Ace05] F. Acerbi. A reference to perfect numbers in Plato’s Theaetetus. Archive for History of Exact Sciences, 59(4):319–348, May 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=4&spage=319.

Acerbi:2008:EP

[Ace08] Fabio Acerbi. Euclid’s Pseudaria. Archive for History of Exact Sciences, 62(5):511–551, September 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=5&spage=511.

Acerbi:2009:MDA

[Ace09] Fabio Acerbi. The meaning of πλασματικóν in Diophantus’ Arithmetica. Archive for History of Exact Sciences, 63(1):5–31, January 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
63&issue=1&spage=5.

Acerbi:2011:GBM

[Ace11a] Fabio Acerbi. The geometry of burning mirrors in Greek antiquity. Analysis, heuristic, projections, lemmatic fragmentation. Archive for History of Exact Sciences, 65(5):471–497, September 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=5&spage=471.

Acerbi:2011:LSF

[Ace11b] Fabio Acerbi. The language of the “Givens”: its forms and its use as a deductive tool in Greek mathematics. Archive for History of Exact Sciences, 65(2):119–153, March 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=2&spage=119.

Acerbi:2019:TNC

[Ace19] F. Acerbi. There is no consequentia mirabilis in Greek mathematics. Archive for History of Exact Sciences, 73(3):217–242, May 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Adelmann:2009:LWB

[AG09] Clemens Adelmann and Eberhard H.-A. Gerbracht. Letters from William Burnside to Robert Fricke: automorphic functions, and the emergence of the Burnside Problem. Archive for History of Exact Sciences, 63(1):33–50, January 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=1&spage=33.

Aydin:2015:REA

[AH15] Nuh Aydin and Lakhdar Hammoudi. Root extraction by Al-Kashi and Stevin. Archive for History of Exact Sciences, 69 (3):291–310, May 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0150-3.

Aiton:1972:LMR

[Ait72] E. J. Aiton. Leibniz on motion in a resisting medium. Archive for History of Exact Sciences, 9(3):257–274, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=9&issue=3&spage=
257.

Aldrich:2009:BEM

[Ald09] John Aldrich. Burnside’s engagement with the “modern theory of statistics”. Archive for History of Exact Sciences, 63(1):51– 79, January 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
63&issue=1&spage=51.

Alvarez:1999:HSP

[Alv99] Carlos Alvarez. On the history of Souslin’s Problem. Archive for History of Exact Sciences, 54(3):181–242, August 1999. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=54&issue=3&spage=181.

Avigad:2014:CO

[AM14] Jeremy Avigad and Rebecca Morris. The concept of “character” in Dirichlet’s theorem on primes in an arithmetic progression. Archive for History of Exact Sciences, 68(3):265–326, May 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-013-0126-0.

Ambrosi:2012:PEG

[Amb12] Gerhard Michael Ambrosi. Pre-Euclidean geometry and Aeginetan coin design: some further remarks. Archive for History of Exact Sciences, 66(5):557–583, September 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=66&issue=5&spage=557.

Andersen:1985:CMI

[And85] Kirsti Andersen. Cavalieri’s method of indivisibles. Archive for History of Exact Sciences, 31(4):291–367, December 1985. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=31&issue=4&spage=291.

Andersson:1994:PDH

[And94] K. G. Andersson. Poincaré’s discovery of homoclinic points. Archive for History of Exact Sciences, 48(2):133–147, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=48&issue=2&spage=
133.

Anonymous:1981:EP

[Ano81] Anonymous. Errata for part I. Archive for History of Exact Sciences, 25(2–3):267, December 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=25&issue=2&spage=267. See [Joh79].

Aoki:1992:MTN

[Aok92] Shinko Aoki. The Moon-test in Newton’s Principia: Accuracy of inverse-square law of universal gravitation. Archive for History of Exact Sciences, 44(2):147–190, June 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=2&spage=147. See corrections

and additions [Aok96].

Aoki:1996:CAS

[Aok96] Shinko Aoki. Corrections and additions for “The Moontest in Newton’s Principia: Accuracy of inverse-square law of universal gravitation”. Archive for History of Exact Sciences, 49(4):393–396, December 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=49&issue=4&spage=393. See [Aok92].

Arboleda:1979:DET

[Arb79] L. C. Arboleda. Les Débuts de l’École Topologique Soviétique: Notes sur les lettres de Paul S. Alexandroff et Paul S. Urysohn à Maurice Fréchet. (French) [The beginnings of the Soviet Topology School: Notes on the letters of Paul S. Alexandroff et Paul S. Urysohn to Maurice Fréchet]. Archive for History of Exact Sciences, 20(1):73–89, March 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=20&issue=1&spage=73.

Arboleda:1981:RMF

[Arb81] L. C. Arboleda. Les recherches de M. Fréchet, P. Alexandrov, W. Sierpinski et K. Kuratowski sur la théorie des types de dimensions et les Débuts de la topologie Générale. (French) [The research of M. Fréchet, P. Alexandrov, W. Sierpinski and K. Kuratowski on the theory of types of dimensions and the beginnings of general topology]. Archive for History of Exact Sciences, 24(4):339–388, December 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=24&issue=4&spage=339.

Ariotti:1972:ACD

[Ari72] Piero E. Ariotti. Aspects of the conception and development of the pendulum in the 17^{th} century. Archive for History of Exact Sciences, 8(5):329–410, January 1972. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=8&issue=5&spage=329.

Arnold:1983:MPSa

[Arn83a] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). I. Physics in France after the revolution. Archive for History of Exact Sciences, 28(3):243–266, September 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=3&spage=243.

Arnold:1983:MPSb

[Arn83b] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). II. The Laplacian program. Archive for History of Exact Sciences, 28(3):267–287, September 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=3&spage=267.

Arnold:1983:MPSc

[Arn83c] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). III. Poisson: mathematician or physicist? Archive for History of Exact Sciences, 28(4):289–297, December 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=28&issue=4&spage=289.

Arnold:1983:MPSd

[Arn83d] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). IV. Disquiet with respect to Fourier’s treatment of heat. Archive for History of Exact Sciences, 28(4):299–320, December 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=28&issue=4&spage=299.

Arnold:1983:MPSi

[Arn83e] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). IX. Poisson’s closing synthesis: traité de physique mathématique. Archive for History of Exact Sciences, 29(1):73–94, March 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=29&issue=1&spage=73.

Arnold:1983:MPSe

[Arn83f] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). V. Fresnel and the circular screen. Archive for History of Exact Sciences, 28(4):321–342, December 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=4&spage=321.

Arnold:1983:MPSf

[Arn83g] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). VI. Elasticity: the crystallization of Poisson’s views on the nature of matter. Archive for History of Exact Sciences, 28(4):343–367, December 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=28&issue=4&spage=343.

Arnold:1983:MPSg

[Arn83h] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). VII. Mécanique physique. Archive for History of Exact Sciences, 29(1):37–51, March 1983. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=29&issue=1&spage=37.

Arnold:1983:MPSh

[Arn83i] D. H. Arnold. The mécanique physique of Siméon Dénis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). VIII. Applications of the mécanique physique. Archive for History of Exact Sciences, 29(1):53–72, March 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
29&issue=1&spage=53.

Arnold:1984:MPS

[Arn84] D. H. Arnold. The mécanique physique of Siméon Denis Poisson: The evolution and isolation in France of his approach to physical theory (1800–1840). X. Some perspective on Poisson’s contributions to the emergence of mathematical physics. Archive for History of Exact Sciences, 29(4):287–307, December 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
29&issue=4&spage=287.

Artmann:1985:VSD

[Art85] Benno Artmann. Über voreuklidische “Elemente”, deren Autor Proportionen vermied. (German) [On pre-Euclidean Elements, whose author avoided proportions]. Archive for History of Exact Sciences, 33(4):291–306, December 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=33&issue=4&spage=291.

Artmann:1988:VER

[Art88] Benno Artmann. Über voreuklidische ‘Elemente der Raumgeometrie’ aus der Schule des Eudoxos. (German) [On pre-Euclidean ‘elements of spatial geometry’ in the School of Eudoxos]. Archive for History of Exact Sciences, 39(2):121– 135, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
39&issue=2&spage=121.

Arthur:2013:LSI

[Art13] Richard T. W. Arthur. Leibniz’s syncategorematic infinitesimals. Archive for History of Exact Sciences, 67(5):553– 593, September 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-013-0119-z.

Ascher:1988:GCI

[Asc88] Marcia Ascher. Graphs in cultures (II): a study in ethnomathematics. Archive for History of Exact Sciences, 39(1):75–95, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
39&issue=1&spage=75.

Archibald:2014:IEB

[AT14] T. Archibald and R. Tazzioli. Integral equations between theory and practice: the cases of Italy and France to 1920. Archive for History of Exact Sciences, 68(5):547–597, September 2014. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0132-2.

Aubin:2017:BFM

[Aub17] David Aubin. Ballistics, fluid mechanics, and air resistance at Gâvre, 1829–1915. Archive for History of Exact Sciences, 71(6): 509–542, November 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-017-0195-6.

Ayoub:1980:PRC

[Ayo80] Raymond G. Ayoub. Paolo Ruffini’s contributions to the quintic. Archive for History of Exact Sciences, 23(3):253–277, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=3&spage=253.

Ayoub:1984:LFC

[Ayo84] Raymond Ayoub. The lemniscate and Fagnano’s contributions to elliptic integrals. Archive for History of Exact Sciences, 29(2):131–149, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=29&issue=2&spage=131.

Bachmakova:1963:MDA

[Bac63] I. G. Bachmakova. Les méthodes différentielles d’Archimède. Archive for History of Exact Sciences, 2(2):87–107, August 1963.

CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=2&spage=
87.

Bach:1989:FFA

[Bac89] Alexander Bach. Eine Fehlinterpretation mit Folgen: Albert Einstein und der Welle–Teilchen Dualismus. (German) [A misinterpretation with consequences: Albert Einstein and the wave– particle dualism]. Archive for History of Exact Sciences, 40(2): 173–206, June 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
40&issue=2&spage=173.

Bach:1990:BPD

[Bac90] Alexander Bach. Boltzmann’s probability distribution of 1877. Archive for History of Exact Sciences, 41(1):1–40, March 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=1&spage=1.

Bascelli:2014:GQU

[Bas14] Tiziana Bascelli. Galileo’s quanti: understanding infinitesimal magnitudes. Archive for History of Exact Sciences, 68(2): 121–136, March 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-013-0124-2.

Bos:1997:ES

[BB97] Henk J. M. Bos and Jed Z. Buchwald. Editorial statement. Archive for History of Exact Sciences, 51(1):1, March 1997. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=51&issue=1&spage=1.

Brack-Bernsen:2020:BAN

[BB20] Lis Brack-Bernsen. Babylonian astronomy: a new understanding of column Φ. Archive for History of Exact Sciences, 74(6): 605–640, November 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00407-020-00254-z.pdf.

Bernard:2012:NAF

[BC12] Alain Bernard and Jean Christianidis. A new analytical framework for the understanding of Diophantus’s Arithmetica I– III. Archive for History of Exact Sciences, 66(1):1–69, January 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
66&issue=1&spage=1.

Breard:2020:CBN

[BC20] Andrea Bréard and Constance A. Cook. Cracking bones and numbers: solving the enigma of numerical sequences on ancient Chinese artifacts. Archive for History of Exact Sciences, 74 (4):313–343, July 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-019-00245-9.

Beckmann:1967:NGB

[Bec67] Friedhelm Beckmann. Neue Gesichtspunkte zum 5. Buch Euklids. (German) [New criteria on the 5th Book of Euclid]. Archive for History of Exact Sciences, 4(1–2):1–144, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=4&issue=1&spage=
1.

Bechler:1973:NSM

[Bec73] Zev Bechler. Newton’s search for a mechanistic model of colour dispersion: a suggested interpretation. Archive for History of Exact Sciences, 11(1):1–37, March 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=11&issue=1&spage=1.

Beller:1985:PJI

[Bel85] Mara Beller. Pascual Jordan’s influence on the discovery of Heisenberg’s indeterminacy principle. Archive for History of Exact Sciences, 33(4):337–349, December 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=33&issue=4&spage=337.

Beller:1988:AJM

[Bel88] Eliyahu Beller. Ancient Jewish mathematical astronomy. Archive for History of Exact Sciences, 38(1):51–66, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=38&issue=1&spage=
51.

Bell:1991:LTC

[Bel91] James F. Bell. The late-twentieth century resolution of a mid-nineteenth century dilemma generated by the eighteenthcentury experiments of Ernst Chladni on the dynamics of rods. Archive for History of Exact Sciences, 43(3):251–273, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
43&issue=3&spage=251.

Beller:1998:NDA

[Bel98] Eliyahu Beller. A newly-discovered ancient value for the length of the year. Archive for History of Exact Sciences, 52(1):91–98, January 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
52&issue=1&spage=91.

Bell:2010:SEW

[Bel10] Jordan Bell. A summary of Euler’s work on the pentagonal number theorem. Archive for History of Exact Sciences, 64(3): 301–373, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Bernkopf:1966:DFS

[Ber66] Michael Bernkopf. The development of function spaces with particular reference to their origins in integral equation theory. Archive for History of Exact Sciences, 3(1):1–96, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=3&issue=1&spage=

Bernkopf:1968:HIM

[Ber68] Michael Bernkopf. A history of infinite matrices. Archive for History of Exact Sciences, 4(4):308–358, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=4&issue=4&spage=

308. A study of denumerably infinite linear systems as the first step in the history of operators defined on function spaces.

Berggren:1976:STA

[Ber76] J. L. Berggren. Spurious theorems in Archimedes’ equilibrium of planes: Book I. Archive for History of Exact Sciences, 16(2):87–103, June 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=2&spage=87.

Berggren:1991:PME

[Ber91] J. L. Berggren. Ptolemy’s maps of Earth and the heavens: a new interpretation. Archive for History of Exact Sciences, 43(2):133– 144, June 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
43&issue=2&spage=133.

Bernard:2003:SAP

[Ber03] Alain Bernard. Sophistic aspects of Pappus’s Collection. Archive for History of Exact Sciences, 57(2):93–150, March 2003. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=57&issue=2&spage=93.

Bartolozzi:1993:RSD

[BF93] Margherita Bartolozzi and Raffaella Franci. La regola dei segni dall’enunciato di R. Descartes (1637) alla dimostrazione di C. F. Gauss (1828). (Italian) [The rule of signs enunciated by R. Descartes (1637) as demonstrated by C. F. Gauss (1828)]. Archive for History of Exact Sciences, 45(4):335–374, December 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
45&issue=4&spage=335.

Barrow-Green:1994:OIP

[BG94] June Barrow-Green. Oscar II’s prize competition and the error in Poincaré’s memoir on the three body problem. Archive for History of Exact Sciences, 48(2):107–131, June 1994. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=48&issue=2&spage=107.

Bowen:1996:GCM

[BG96] Alan C. Bowen and Bernard R. Goldstein. Geminus and the concept of mean motion in Greco–Latin astronomy. Archive for History of Exact Sciences, 50(2):157–185, August 1996. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=50&issue=2&spage=157.

Buchwald:2008:AHE

[BG08] J. Z. Buchwald and J. J. Gray. The Archive for History of Exact Sciences expands its scope to include the history of modern biology. Archive for History of Exact Sciences, 62(6):601–602, November 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=6&spage=601.

Buchwald:2009:JUT

[BG09] J. Z. Buchwald and J. J. Gray. Journals under threat: a joint response from History of Science, Technology and Medicine Editors. Archive for History of Exact Sciences, 63(1):1–4, January 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
63&issue=1&spage=1.

Brazil:2014:HES

[BGTZ14] Marcus Brazil, Ronald L. Graham, Doreen A. Thomas, and Martin Zachariasen. On the history of the Euclidean Steiner tree problem. Archive for History of Exact Sciences, 68(3):327–354, May 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-013-0127-z.

Bajri:2015:RAS

[BHM15] Sanaa Bajri, John Hannah, and Clemency Montelle. Revisiting Al–Samaw’al’s table of binomial coefficients: Greek inspiration, diagrammatic reasoning and mathematical induction. Archive for History of Exact Sciences, 69(6):537–576, November 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-015-0156-x.

Bierhalter:1981:BMG

[Bie81a] Günter Bierhalter. Boltzmanns mechanische Grundlegung des zweiten Hauptsatzes der Wärmelehre aus dem Jahre 1866. (German) [Boltzmann’s mechanical foundation of the second main clause of the Theory of Heat from the year 1866]. Archive for History of Exact Sciences, 24(3):195–205, September 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
24&issue=3&spage=195.

Bierhalter:1981:CMG

[Bie81b] Günter Bierhalter. Clausius’ mechanische Grundlegung des zweiten Hauptsatzes der Wärmelehre aus dem Jahre 1871. (German) [Clausius’s mechanical foundation of the second main clause of the caloric theory from the year 1871]. Archive for History of Exact Sciences, 24(3):207–220, September 1981. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=24&issue=3&spage=207.

Bierhalter:1981:HHM

[Bie81c] Günter Bierhalter. Zu Hermann von Helmholtzens mechanischer Grundlegung der Wärmelehre aus dem Jahre 1884. (German) [On Hermann von Helmholtz’s of mechanical foundation of the caloric theory from the year 1884]. Archive for History of Exact Sciences, 25(1):71–84, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=25&issue=1&spage=71.

Bierhalter:1982:VSB

[Bie82] Günter Bierhalter. Das Virialtheorem in seiner Beziehung zu den mechanischen Grundlegungen des zweiten Hauptsatzes der Wärmelehre. (German) [The Virial Theorem in its relationship with mechanical foundations of the second main clause of caloric theory]. Archive for History of Exact Sciences, 27(3):199–211, September 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=3&spage=199.

Bierhalter:1983:VHM

[Bie83a] Günter Bierhalter. Die v. Helmholtzschen Monozykel–Analogien zur Thermodynamik und das Clausiussche Disgregationskonzept. (German) [The von Helmholtz monocycle — analogies to thermodynamics and the Clausius disgregation concept]. Archive for History of Exact Sciences, 29(1):95–100, March 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
29&issue=1&spage=95.

Bierhalter:1983:SVM

[Bie83b] Günter Bierhalter. Zu Szilys Versuch einer mechanischen Grundlegung des zweiten Hauptsatzes der Thermodynamik. (German) [On Szilys’ attempt at a mechanical foundation of the second main clause of thermodynamics]. Archive for History of Exact Sciences, 28(1):25–35, March 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=28&issue=1&spage=25.

Bierhalter:1985:MED

[Bie85] Günter Bierhalter. Die mechanischen Entropie- und Disgregationskonzepte aus dem 19. Jahrhundert: ihre Grundlagen, ihr Versagen und ihr Entstehungshintergrund. (German) [Mechanical entropy and disgregations concepts from the 19^{th} century: their basis, failure and developing background]. Archive for History of Exact Sciences, 32(1):17–41, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=1&spage=17.

Bierhalter:1987:WEW

[Bie87] Günter Bierhalter. Wie erfolgreich waren die im 19. Jahrhundert betriebenen Versuche einer mechanischen Grundlegung des zweiten Hauptsatzes der Thermodynamik?. (German) [Were 19^{th} century attempts at a mechanical foundation of the second main clause of thermodynamics successful?]. Archive for History of Exact Sciences, 37(1):77–99, March 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=37&issue=1&spage=77.

Bierhalter:1992:BBJ

[Bie92] Günter Bierhalter. Von L. Boltzmann bis J. J. Thomson: die Versuche einer mechanischen Grundlegung der Thermodynamik (1866–1890). (German) [From L. Boltzmann to J. J. Thomson: the search for a mechanical basis for thermodynamics (1866–1890)]. Archive for History of Exact Sciences, 44(1):25– 75, March 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
44&issue=1&spage=25.

Bien:2004:GBM

[Bie04] Reinhold Bien. Gauß and beyond: The making of Easter algorithms. Archive for History of Exact Sciences, 58(5):439– 452, July 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
58&issue=5&spage=439.

Bien:2007:VCC

[Bie07] Reinhold Bien. Viète’s controversy with Clavius over the truly Gregorian calendar. Archive for History of Exact Sciences, 61(1):39–66, January 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=1&spage=39.

Biggs:1983:MMC

[Big83] N. L. Biggs. De Morgan on map colouring and the separation axiom. Archive for History of Exact Sciences, 28(2):165– 170, June 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=2&spage=165.

Biggs:2019:THC

[Big19] Norman Biggs. Thomas Harriot on the coinage of England. Archive for History of Exact Sciences, 73(4):361–383, July 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-019-00228-w.pdf.

Bikerman:1978:CBL

[Bik78] J. J. Bikerman. Capillarity before Laplace: Clairaut, Segner, Monge, Young. Archive for History of Exact Sciences, 18(2): 103–122, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
18&issue=2&spage=103.

Bilodeau:1982:OED

[Bil82] G. G. Bilodeau. The origin and early development of nonanalytic infinitely differentiable functions. Archive for History of Exact Sciences, 27(2):115–135, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=2&spage=115.

Bistafa:2022:EFT

[Bis22] Sylvio R. Bistafa. Euler first theory of resonance. Archive for History of Exact Sciences, 76(3):207–221, May 2022. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-021-00280-5.

Britton:2000:NBP

[BJ00] John P. Britton and Alexander Jones. A new Babylonian planetary model in a Greek source. Archive for History of Exact Sciences, 54(4):349–373, January 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=4&spage=349.

Blaasjo:2016:DGA

[Blå16] Viktor Blåsjö. In defence of geometrical algebra. Archive for History of Exact Sciences, 70(3):325–359, May 2016. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
015-0169-5.

Blum:2022:ESB

[Blu22] Alexander S. Blum. Einstein’s second-biggest blunder: the mistake in the 1936 gravitational-wave manuscript of Albert Einstein and Nathan Rosen. Archive for History of Exact Sciences, 76(6):623–632, November 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00295-6.

Barone:1978:HAF

[BN78] Jack Barone and Albert Novikoff. A history of the axiomatic formulation of probability from Borel to Kolmogorov: Part I. Archive for History of Exact Sciences, 18(2):123– 190, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
18&issue=2&spage=123.

Brigaglia:1984:SGS

[BN84] Aldo Brigaglia and Pietro Nastasi. Le soluzioni di Girolamo Saccheri e Giovanni Ceva al ‘geometram quaero’ di Ruggero Ventimiglia: Geometria proiettiva italiana nel tardo seicento. (Italian) [The solutions of Girolamo Saccheri and Giovanni Ceva for ‘Geometram quaero’ by Ruggero Ventimiglia: Italian projective geometry in the late sixth century]. Archive for History of Exact Sciences, 30(1):7–44, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=30&issue=1&spage=7.

Bergia:1988:RCE

[BN88] S. Bergia and L. Navarro. Recurrences and continuity in Einstein’s research on radiation between 1905 and 1916. Archive for History of Exact Sciences, 38(1):79–99, March 1988. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=38&issue=1&spage=79.

Butzer:1993:AVP

[BN93] P. L. Butzer and R. J. Nessel. Aspects of de La Vallée Poussin’s work in approximation and its influence. Archive for History of Exact Sciences, 46(1):67–95, March 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=46&issue=1&spage=67.

Bockstaele:2009:BVD

[Boc09] Paul Bockstaele. Between Viète and Descartes: Adriaan van Roomen and the Mathesis Universalis. Archive for History of Exact Sciences, 63(4):433–470, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Bollinger:1972:GEH

[Bol72] Maja Bollinger. Geschichtliche Entwicklung des Homologiebegriffs. (German) [Historical development of the terms of homology]. Archive for History of Exact Sciences, 9(2):94–170, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
9&issue=2&spage=94.

Boos:1985:STG

[Boo85] William Boos. “The True” in Gottlob Frege’s “Über die Grundlagen der Geometrie” [“On the foundations of geometry”]. Archive for History of Exact Sciences, 34(1–2):141–192, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
34&issue=1&spage=141.

Borzacchini:2007:IMC

[Bor07] Luigi Borzacchini. Incommensurability, music and continuum: a cognitive approach. Archive for History of Exact Sciences, 61(3):273–302, May 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=3&spage=273.

Bos:1974:DHO

[Bos74] H. J. M. Bos. Differentials, higher-order differentials and the derivative in the Leibnizian calculus. Archive for History of Exact Sciences, 14(1):1–90, March 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=14&issue=1&spage=1.

Bos:1981:RCD

[Bos81] H. J. M. Bos. On the representation of curves in Descartes’ Géométrie. Archive for History of Exact Sciences, 24 (4):295–338, December 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=24&issue=4&spage=295.

Bos:1984:AMR

[Bos84] H. J. M. Bos. Arguments on motivation in the rise and decline of a mathematical theory; the “construction of equations”, 1637-ca.1750. Archive for History of Exact Sciences, 30(3–4):331–380, September 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=30&issue=3&spage=331.

Bottazzini:1977:FPM

[Bot77a] Umberto Bottazzini. Le funzioni a periodi multipli nella corrispondenza tra Hermite e Casorati. (Italian) [Multiple-periodic functions in the correspondence between Hermite and Casorati]. Archive for History of Exact Sciences, 18(1):39–88, March 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
18&issue=1&spage=39.

Bottazzini:1977:REB

[Bot77b] Umberto Bottazzini. Riemanns Einfluß auf E. Betti und F. Casorati. (German) [Riemann’s influence on E. Betti and F. Casorati]. Archive for History of Exact Sciences, 18(1):27– 37, March 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
18&issue=1&spage=27.

Boucard:2013:CFQ

[Bou13] Jenny Boucard. Cyclotomie et formes quadratiques dans l’oeuvre arithmétique d’Augustin-Louis Cauchy (1829–1840). Archive for History of Exact Sciences, 67(4):349–414, July 2013. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0115-3.

Britton:2011:PRD

[BPS11] John P. Britton, Christine Proust, and Steve Shnider. Plimpton 322: a review and a different perspective. Archive for History of Exact Sciences, 65(5):519–566, September 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=65&issue=5&spage=519.

Brakel:1985:PID

[Bra85] J. Brakel. The possible influence of the discovery of radio-active decay on the concept of physical probability. Archive for History of Exact Sciences, 31(4):369–385, December 1985. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=31&issue=4&spage=369.

Brackenridge:2003:NEQ

[Bra03] J. Bruce Brackenridge. Newton’s easy quadratures “omitted for the sake of brevity”. Archive for History of Exact Sciences, 57(4):313–336, May 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=57&issue=4&spage=313.

Braating:2007:NLG

[Brå07] Kajsa Bråting. A new look at E. G. Björling and the Cauchy sum theorem. Archive for History of Exact Sciences, 61(5):519–535, September 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=5&spage=519.

Breitenberger:1984:GGA

[Bre84] Ernst Breitenberger. Gauss’s geodesy and the Axiom of Parallels. Archive for History of Exact Sciences, 31(3):273–289, September 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
31&issue=3&spage=273.

Breger:1994:MAV

[Bre94a] Herbert Breger. The mysteries of adaequare: a vindication of Fermat. Archive for History of Exact Sciences, 46(3):193–219, February 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=3&spage=193.

Brentjes:1994:THA

[Bre94b] Sonja Brentjes. Textzeugen und Hypothesen zum arabischen Euklid in der Überlieferung von al-H .ağğāğ b. Yūsuf b. Mat.ar (zwischen 786 und 833). (German) [Text witness and hypotheses on the Arabic Euclid in the narration of al-H .ağğāğ b. Yūsuf b. Mat.ar (between 786 und 833)]. Archive for History of Exact Sciences, 47(1):53–92, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=1&spage=53.

Britton:2001:RST

[Bri01] John P. Britton. Remarks on a system: a text for Venus: ACT 1050. Archive for History of Exact Sciences, 55(6):525– 554, June 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=6&spage=525.

Britton:2007:SBL

[Bri07] John P. Britton. Studies in Babylonian lunar theory: Part I. Empirical elements for modeling lunar and solar anomalies. Archive for History of Exact Sciences, 61(2):83–145, March 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=2&spage=83.

Britton:2009:SBL

[Bri09] John P. Britton. Studies in Babylonian lunar theory. II. Treatments of lunar anomaly. Archive for History of Exact Sciences, 63(4):357–431, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Britton:2010:SBL

[Bri10] John P. Britton. Studies in Babylonian lunar theory: part III. The introduction of the uniform zodiac. Archive for History of Exact Sciences, 64(6):617–663, November 2010. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=64&issue=6&spage=617.

Briend:2021:MPD

[Bri21] Jean-Yves Briend. Mathématiques en perspective: Desargues, la Hire, le Poîvre. Archive for History of Exact Sciences, 75(6): 699–736, November 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00275-2.

Bromberg:1967:MDC

[Bro67] Joan Bromberg. Maxwell’s displacement current and his theory of light. Archive for History of Exact Sciences, 4(3):218–234, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
4&issue=3&spage=218.

Brownson:1981:EOC

[Bro81] C. D. Brownson. Euclid’s optics and its compatibility with linear perspective. Archive for History of Exact Sciences, 24(3):165–194, September 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=24&issue=3&spage=165.

Browne:2005:HMC

[Bro05] L. W. B. Browne. Halley’s method for calculating the Earth– Sun distance. Archive for History of Exact Sciences, 59(3):251– 266, March 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=3&spage=251.

Bromberg:2016:ELL

[Bro16] Joan Lisa Bromberg. Explaining the laser’s light: classical versus quantum electrodynamics in the 1960s. Archive for History of Exact Sciences, 70(3):243–266, May 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-015-0166-8.

Brush:1967:FSM

[Bru67] Stephen G. Brush. Foundations of statistical mechanics 1845– 1915. Archive for History of Exact Sciences, 4(3):145–183, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
4&issue=3&spage=145.

Brush:1968:HRP

[Bru68] Stephen G. Brush. A history of random processes. I. Brownian motion from Brown to Perrin. Archive for History of Exact Sciences, 5(1):1–36, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=1&spage=1.

Brush:1973:DKT

[Bru73] Stephen G. Brush. The development of the kinetic theory of gases. Archive for History of Exact Sciences, 11(1):38–96, March 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
11&issue=1&spage=38.

Brush:1974:DKT

[Bru74] Stephen G. Brush. The development of the kinetic theory of gases. Archive for History of Exact Sciences, 12(1):1–88, March 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
12&issue=1&spage=1.

Bruin:1981:ARE

[Bru81] Frans Bruin. Atmospheric refraction and extinction near the horizon. Archive for History of Exact Sciences, 25(1):1–17, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
25&issue=1&spage=1.

Bachmakova:1977:OTF

[BS77] I. G. Bachmakova and E. I. Slavutin. ≪Genesis triangulorum≫ de François Viète et ses recherches dans l’analyse indéterminée. (French) [≪Genesis triangulorum≫ of François Viète and his research in indeterminate analysis]. Archive for History of Exact Sciences, 16(4):289–306, December 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=4&spage=289.

Benis-Sinaceur:1984:DHA

[BS84] Hourya Benis-Sinaceur. De D. Hilbert à E. Artin: Les différents aspects du dix-septième problème et les filiations conceptuelles de la théorie des corps réels clos. (French) [From D. Hilbert to E. Artin: different aspects of the 17^{th} problem and the conceptual relations of the theory of closed real bodies]. Archive for History of Exact Sciences, 29(3):267–286, September 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
29&issue=3&spage=267.

Benis-Sinaceur:1985:TAS

[BS85] Hourya Benis-Sinaceur. La théorie d’artin et Schreier et l’analyse non-standard d’abraham Robinson. Archive for History of Exact Sciences, 34(3):257–264, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=34&issue=3&spage=257.

Berggren:2007:ASD

[BS07] J. L. Berggren and Nathan Sidoli. Aristarchus’s On the Sizes and Distances of the Sun and the Moon: Greek and Arabic texts. Archive for History of Exact Sciences, 61(3):213– 254, May 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=3&spage=213.

Belle:2022:GGC

[BS22] Riccardo Bellé and Beatrice Sisana. Galileo Galilei and the centers of gravity of solids: a reconstruction based on a newly discovered version of the conical frustum contained in manuscript UCLA 170/624. Archive for History of Exact Sciences, 76(5): 471–511, September 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00289-4.

Butzer:1988:OHC

[BSS88] P. L. Butzer, M. Schmidt, and E. L. Stark. Observations on the history of central B-splines. Archive for History of Exact Sciences, 39(2):137–156, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=2&spage=137.

Buchwald:1980:EID

[Buc80a] Jed Z. Buchwald. Experimental investigations of double refraction from Huygens to Malus. Archive for History of Exact Sciences, 21(4):311–373, December 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=4&spage=311.

Buchwald:1980:OTP

[Buc80b] Jed Z. Buchwald. Optics and the theory of the punctiform ether. Archive for History of Exact Sciences, 21(3):245–278, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
21&issue=3&spage=245.

Buchwald:2001:PDB

[Buc01] Jed Z. Buchwald. A potential disagreement between Helmholtz and Hertz. Archive for History of Exact Sciences, 55(4):365– 393, January 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=4&spage=365.

Buchwald:2006:DME

[Buc06] Jed Z. Buchwald. Discrepant measurements and experimental knowledge in the early modern era. Archive for History of Exact Sciences, 60(6):565–649, November 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=60&issue=6&spage=565.

Buchwald:2007:HMD

[Buc07] Jed Z. Buchwald. Huygens’ methods for determining optical parameters in birefringence. Archive for History of Exact Sciences, 61(1):67–81, January 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=1&spage=67.

Bullynck:2009:RGC

[Bul09] Maarten Bullynck. Reading Gauss in the computer age: on the U.S. reception of Gauss’s number-theoretical work (1938– 1989). Archive for History of Exact Sciences, 63(5):553–580, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Burckhardt:1967:GER

[Bur67] J. J. Burckhardt. Zur Geschichte der Entdeckung der 230 Raumgruppen. (German) [On the history of the discovery of the 230 space groups]. Archive for History of Exact Sciences, 4(3):235– 246, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
4&issue=3&spage=235.

Burckhardt:1971:BFS

[Bur71] J. J. Burckhardt. Der Briefwechsel von E. S. von Fedorow und A. Schoenflies, 1889–1908. (German) [Letters between E. S. von Fedorow and A. Schoenflies, 1889–1908]. Archive for History of Exact Sciences, 7(2):91–141, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=7&issue=2&spage=91.

Burckhardt:1972:BFF

[Bur72] J. J. Burckhardt. Der Briefwechsel von E. S. von Fedorow und F. Klein, 1893. (German) [Letters between E. S. von Fedorow und F. Klein, 1893]. Archive for History of Exact Sciences, 9(2):85–93, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=9&issue=2&spage=85.

Busard:1965:PGD

[Bus65a] H. L. L. Busard. The Practica Geometriae of Dominicus de Clavasio. Archive for History of Exact Sciences, 2(6):520–575, December 1965. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
2&issue=6&spage=520.

Busard:1965:URE

[Bus65b] H. L. L. Busard. Unendliche Reihen in A est unum calidum. (German) [Infinite series in A est unum calidum]. Archive for History of Exact Sciences, 2(5):387–397, August 1965. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=5&spage=
387.

Busard:1998:ZMT

[Bus98] H. L. L. Busard. Zwei mittelalterliche Texte zur theoretischen Mathematik: Die Arithmetica speculativa von Thomas Bradwardine und die Theorica numerorum von Wigandus Durnheimer. (German) [Two medieval texts of theoretical mathematics: The Arithmetica speculativa of Thomas Bradwardine and the Theorica numerorum of Wigandus Durnheimer]. Archive for History of Exact Sciences, 53(2):97– 124, June 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=2&spage=97.

Bussotti:2019:MCF

[Bus19] Paolo Bussotti. Michel Chasles’ foundational programme for geometry until the publication of his Aperçu historique. Archive for History of Exact Sciences, 73(3):261–308, May 2019. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Buchwald:2016:KTO

[BY16] Jed Z. Buchwald and Chen-Pang Yeang. Kirchhoff’s theory for optical diffraction, its predecessor and subsequent development: the resilience of an inconsistent theory. Archive for History of Exact Sciences, 70(5):463–511, September 2016. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
016-0176-1.

Buchwald:2021:CWH

[BYS^{+}21a] Jed Buchwald, Chen-Pang Yeang, Noah Stemeroff, Jenifer Barton, and Quinn Harrington. Correction to: What Heinrich Hertz discovered about electric waves in 1887-1888. Archive for History of Exact Sciences, 75(2):173, March 2021. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-020-00267-8.pdf. See [BYS^{+}21b].

Buchwald:2021:WHH

[BYS^{+}21b] Jed Buchwald, Chen-Pang Yeang, Noah Stemeroff, Jenifer Barton, and Quinn Harrington. What Heinrich Hertz discovered about electric waves in 1887–1888. Archive for History of Exact Sciences, 75(2):125–171, March 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). See correction [BYS^{+}21a].

Calinger:1976:ELP

[Cal76] Ronald Calinger. Euler’s “Letters to a princess of Germany” as an expression of his mature scientific outlook. Archive for History of Exact Sciences, 15(3):211–233, September 1976. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=15&issue=3&spage=211.

Cao:1991:RPA

[Cao91] Tian Yu Cao. The reggeization program 1962–1982: Attempts at reconciling quantum field theory with S-matrix theory. Archive for History of Exact Sciences, 41(3):239–283, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=3&spage=239.

Caparrini:2002:DVR

[Cap02] Sandro Caparrini. The discovery of the vector representation of moments and angular velocity. Archive for History of Exact Sciences, 56(2):151–181, January 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=56&issue=2&spage=151.

Carrier:1986:BEA

[Car86] Martin Carrier. Die begriffliche Entwicklung der Affinitätstheorie im 18. Jahrhundert Newtons Traum- und was daraus wurde. (German) [The conceptual development of affinity theory in 18^{th} century: Newton’s dream and what it became]. Archive for History of Exact Sciences, 36(4):327–389, December 1986. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=36&issue=4&spage=327.

Carman:2009:RNP

[Car09] Christián C. Carman. Rounding numbers: Ptolemy’s calculation of the Earth–Sun distance. Archive for History of Exact Sciences, 63(2):205–242, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Carman:2014:TPA

[Car14] Christián C. Carman. Two problems in Aristarchus’s treatise on the sizes and distances of the sun and moon. Archive for History of Exact Sciences, 68(1):35–65, January 2014. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0123-3.

Carman:2017:MCC

[Car17] Christián C. Carman. Martianus Capella’s calculation of the size of the Moon. Archive for History of Exact Sciences, 71 (2):193–210, March 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-016-0185-0.

Carman:2018:FCW

[Car18] Christián C. Carman. The first Copernican was Copernicus: the difference between Pre–Copernican and Copernican heliocentrism. Archive for History of Exact Sciences, 72(1):1–20, January 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Carman:2020:TBC

[Car20] Christián C. Carman. Tycho Brahe’s Calculi ad Corrigenda Elementa Orbitae Saturni and the technical aspects of his planetary model of Saturn. Archive for History of Exact Sciences, 74(6): 565–586, November 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Carman:2021:GIJ

[Car21] Christián C. Carman. The gravitational influence of Jupiter on the Ptolemaic value for the eccentricity of Saturn. Archive for History of Exact Sciences, 75(4):439–454, July 2021. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-020-00271-y.

Casey:1992:PR

[Cas92] James Casey. The principle of rigidification. Archive for History of Exact Sciences, 43(4):329–383, December 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=43&issue=4&spage=329.

Ciliberto:2013:PDP

[CD13] Ciro Ciliberto and Emma Sallent Del Colombo. Pasquale del Pezzo, Duke of Caianello, Neapolitan mathematician. Archive for History of Exact Sciences, 67(2):171–214, March 2013. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
012-0110-0.

Carman:2014:EAM

[CE14] Christián C. Carman and James Evans. On the epoch of the Antikythera mechanism and its eclipse predictor. Archive for History of Exact Sciences, 68(6):693–774, November 2014.

CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-014-0145-5;   http://www.scientificcomputing.
com/news/2014/12/worlds-oldest-computer-ancient-greek-
antikythera-mechanism-100-years-older-previously-believed.

Capobianco:2017:GAE

[CEF17] Giovanni Capobianco, Maria Rosaria Enea, and Giovanni Ferraro. Geometry and analysis in Euler’s integral calculus. Archive for History of Exact Sciences, 71(1):1–38, January 2017. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
016-0179-y.

Cenadelli:2010:SGS

[Cen10] Davide Cenadelli. Solving the Giant Stars problem: Theories of stellar evolution from the 1930s to the 1950s. Archive for History of Exact Sciences, 64(2):203–267, March 2010. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=64&issue=2&spage=203.

Contro:1970:KPC

[CFH^{+}70] Walter Salvator Contro, Johannes Viktor Feitzinger, Rolf Hartmann, Friedrich W. Ihloff, Hans-Georg Märtl, Friedemann Rex, Matthias Schramm, and Horst Zehe. Zur Kinematik der Planetenbewegung in Copernicus’ Commentariolus (German) [On the kinematics of planetary motion in Copernicus’ Commentariolus]. Archive for History of Exact Sciences, 6(5):360–371, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=5&spage=360.

Carazza:1980:NDK

[CG80] B. Carazza and G. P. Guidetti. La nascita dell’equazione di Klein–Gordon. (Italian) [The origin of the Klein–Gordon equation]. Archive for History of Exact Sciences, 22(4):373–383, December 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
22&issue=4&spage=373.

Carazza:1986:CEM

[CG86] B. Carazza and G. P. Guidetti. The Casimir electron model. Archive for History of Exact Sciences, 35(3):273–279, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=3&spage=273.

Chabas:1994:AAA

[CG94] José Chabás and Bernard R. Goldstein. Andalusian astronomy: al-Zîj al-Muqtabis of Ibn al-Kammâd. Archive for History of Exact Sciences, 48(1):1–41, March 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=48&issue=1&spage=1.

Chabas:2013:DTL

[CG13] José Chabás and Bernard R. Goldstein. Displaced tables in Latin: the Tables for the Seven Planets for 1340. Archive for History of Exact Sciences, 67(1):1–42, January 2013. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
012-0106-9.

Chabas:2015:IAK

[CG15] José Chabás and Bernard R. Goldstein. Ibn al-Kammad’s Muqtabis zij and the astronomical tradition of Indian origin in the Iberian Peninsula. Archive for History of Exact Sciences, 69(6):577–650, November 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-015-0158-8.

Chabas:2019:MMM

[CG19] José Chabás and Bernard R. Goldstein. The medieval Moon in a matrix: double argument tables for lunar motion. Archive for History of Exact Sciences, 73(4):335–359, July 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

CardonaSuarez:2020:LRK

[CG20] Carlos Alberto Cardona Suárez and Juliana Gutiérrez Valderrama. The law of refraction and Kepler’s heuristics. Archive for History of Exact Sciences, 74(1):45–75, January 2020. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-019-00236-w.pdf.

Christidis:1987:HRM

[CGM87] Theodore Christidis, Yorgos Goudaroulis, and Maria Mikou. The heuristic role of mathematics in the initial development of superconductivity theory. Archive for History of Exact Sciences, 37(2):183–191, June 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=37&issue=2&spage=183.

Chapman:1962:AHG

[Cha62] Sydney Chapman. Alexander von Humboldt and geomagnetic science. Archive for History of Exact Sciences, 2(1):41–51, August 1962. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
2&issue=1&spage=41.

Chabas:1991:ATJ

[Cha91] José Chabás. The astronomical tables of Jacob ben David Bonjorn. Archive for History of Exact Sciences, 42(4):279–314, December 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
42&issue=4&spage=279.

Chabas:2004:ACE

[Cha04] José Chabás. Astronomy for the Court in the early sixteenth century: Alfonso de Córdoba and his Tabule Atronomice Elisabeth Regine. Archive for History of Exact Sciences, 58(3):183– 217, March 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
58&issue=3&spage=183.

Chalmers:2015:RBM

[Cha15] Alan Chalmers. Robert Boyle’s mechanical account of hydrostatics and pneumatics: fluidity, the spring of the air and their relationship to the concept of pressure. Archive for History of Exact Sciences, 69(5):429–454, September 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-015-0159-7.

Chabas:2016:ATM

[Cha16] José Chabás. An analysis of the Tabulae magistrales by Giovanni Bianchini. Archive for History of Exact Sciences, 70(5): 543–552, September 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-016-0178-z.

Chen:1997:DPL

[Che97] Xiang Chen. The debate on the “polarity of light” during the optical revolution. Archive for History of Exact Sciences, 50(3–4):359–393, September 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=3&spage=359.

Chen:2015:TTE

[Che15] Jiang-Ping Jeff Chen. Trigonometric tables: explicating their construction principles in China. Archive for History of Exact Sciences, 69(5):491–536, September 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-015-0162-z.

Chorlay:2010:PSC

[Cho10] Renaud Chorlay. From problems to structures: the Cousin problems and the emergence of the sheaf concept. Archive for History of Exact Sciences, 64(1):1–73, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Chorlay:2011:SGF

[Cho11] Renaud Chorlay. ”local–global”: the first twenty years. Archive for History of Exact Sciences, 65(1):1–66, January 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=65&issue=1&spage=1.

Ciccoli:2015:FFR

[Cic15] N. Ciccoli. Fantappiè’s “final relativity” and deformations of Lie algebras. Archive for History of Exact Sciences, 69(3):311–326, May 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-015-0151-2.

Ciocci:2022:FCL

[Cio22] Argante Ciocci. Federico Commandino and the Latin edition of Pappus’ Collection. Archive for History of Exact Sciences, 76 (2):129–151, March 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00284-1.

Ciocci:2023:FCH

[Cio23a] Argante Ciocci. Federico Commandino and his Latin edition of Aristarchus’s On the Sizes and Distances of the Sun and the Moon. Archive for History of Exact Sciences, 77(1):1–23, January 2023. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL https://link.springer.com/
article/10.1007/s00407-022-00294-7.

Ciocci:2023:FCL

[Cio23b] Argante Ciocci. Federico Commandino and the Latin edition of Apollonius’s Conics (1566). Archive for History of Exact Sciences, 77(4):393–421, July 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-023-00307-z.

Chemla:2015:HDE

[CM15] Karine Chemla and Biao Ma. How do the earliest known mathematical writings highlight the state’s management of grains in early imperial China? Archive for History of Exact Sciences, 69(1):1–53, January 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-014-0139-3.

Cogliati:2011:GCK

[Cog11] Alberto Cogliati. On the genesis of the Cartan–Kähler theory. Archive for History of Exact Sciences, 65(4):397–435, July 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=4&spage=397.

Cogliati:2014:JTT

[Cog14] Alberto Cogliati. On Jacobi’s transformation theory of elliptic functions. Archive for History of Exact Sciences, 68(4):529–545, July 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-013-0131-3.

Cogliati:2015:VTC

[Cog15] Alberto Cogliati. Variations on a theme: Clifford’s parallelism in elliptic space. Archive for History of Exact Sciences, 69 (4):363–390, July 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0154-z.

Cogliati:2022:VGA

[Cog22] Alberto Cogliati. Vitali’s generalized absolute differential calculus. Archive for History of Exact Sciences, 76(1):15–43, January 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-021-00273-4.

Cohen:1998:NDM

[Coh98] I. Bernard Cohen. Newton’s determination of the masses and densities of the Sun, Jupiter, Saturn, and the Earth. Archive for History of Exact Sciences, 53(1):83–95, May 1998. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=53&issue=1&spage=83.

Collison:1977:OCB

[Col77] Mary Joan Collison. The origins of the cubic and biquadratic reciprocity laws. Archive for History of Exact Sciences, 17(1): 63–69, May 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
17&issue=1&spage=63.

Contro:1976:PH

[Con76] Walter S. Contro. Von Pasch zu Hilbert. (German) [From Pasch to Hilbert]. Archive for History of Exact Sciences, 15(3):283–295, September 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=15&issue=3&spage=283.

Confalonieri:2015:CIC

[Con15] Sara Confalonieri. The casus irreducibilis in Cardano’s Ars Magna and De Regula Aliza. Archive for History of Exact Sciences, 69(3):257–289, May 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-015-0149-9.

Confalonieri:2018:FAC

[Con18] Sara Confalonieri. A further analysis of Cardano’s main tool in the De Regula Aliza: on the origins of the splittings. Archive for History of Exact Sciences, 72(3):303–352, May 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Cooke:1993:UTS

[Coo93] Roger Cooke. Uniqueness of trigonometric series and descriptive set theory, 1870–1985. Archive for History of Exact Sciences, 45(4):281–334, December 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=45&issue=4&spage=281.

Corry:1997:DHA

[Cor97a] Leo Corry. David Hilbert and the axiomatization of physics (1894–1905). Archive for History of Exact Sciences, 51(2):83– 198, June 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=2&spage=83.

Corry:1997:HMP

[Cor97b] Leo Corry. Hermann Minkowski and the postulate of relativity. Archive for History of Exact Sciences, 51(4):273–314, December 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=4&spage=273.

Corry:1999:DHB

[Cor99] Leo Corry. David Hilbert between mechanical and electromagnetic reductionism (1910–1915). Archive for History of Exact Sciences, 53(6):489–527, January 1999. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=53&issue=6&spage=489.

Corry:2008:NCV

[Cor08] Leo Corry. Number crunching vs. number theory: computers and FLT, from Kummer to SWAC (1850–1960), and beyond. Archive for History of Exact Sciences, 62(4):393–455, July 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=4&spage=393.

Corry:2013:GAM

[Cor13] Leo Corry. Geometry and arithmetic in the medieval traditions of Euclid’s Elements: a view from Book II. Archive for History of Exact Sciences, 67(6):637–705, November 2013. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0121-5.

Coutinho:2004:QRN

[Cou04] S. C. Coutinho. Quotient rings of noncommutative rings in the first half of the 20th century. Archive for History of Exact Sciences, 58(3):255–281, March 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=3&spage=255.

Coutinho:2010:LCP

[Cou10] S. C. Coutinho. A lost chapter in the pre-history of algebraic analysis: Whittaker on contact transformations. Archive for History of Exact Sciences, 64(6):665–706, November 2010. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=64&issue=6&spage=665.

Coutinho:2014:WAD

[Cou14] S. C. Coutinho. Whittaker’s analytical dynamics: a biography. Archive for History of Exact Sciences, 68(3):355–407, May 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-013-0133-1.

Castelao:2007:NSE

[CPFA07] José-Miguel Pacheco Castelao, F. Javier Pérez-Fernández, and Carlos-Oswaldo Suárez Alemán. Numerical solving of equations in the work of José Mariano Vallejo. Archive for History of Exact Sciences, 61(5):537–552, September 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=5&spage=537.

Capecchi:2007:PCC

[CR07] Danilo Capecchi and Giuseppe C. Ruta. Piola’s contribution to continuum mechanics. Archive for History of Exact Sciences, 61(4):303–342, July 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=4&spage=303.

Carman:2019:PPM

[CR19] Christián C. Carman and Gonzalo L. Recio. Ptolemaic planetary models and Kepler’s laws. Archive for History of Exact Sciences, 73(1):39–124, January 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Craik:2013:CCR

[Cra13] Alex D. D. Craik. “continuity and change”: representing mass conservation in fluid mechanics. Archive for History of Exact Sciences, 67(1):43–80, January 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-012-0108-7.

Capecchi:2010:CVM

[CRT10] Danilo Capecchi, Giuseppe Ruta, and Patrizia Trovalusci. From classical to Voigt’s molecular models in elasticity. Archive for History of Exact Sciences, 64(5):525–559, 2010. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Cardin:2024:LCS

[CT24] Franco Cardin and Rossana Tazzioli. Levi-Civita simplifies Einstein. The Ricci rotation coefficients and unified field theories. Archive for History of Exact Sciences, 78(1):87–126, January 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00322-0.

Dahan:1980:TCS

[Dah80] Amy Dahan. Les travaux de Cauchy sur les substitutions. Étude de son approche du concept de groupe. (French) [The works of Cauchy on substitutions: Study of his approach to the group concept]. Archive for History of Exact Sciences, 23(4):279–319, December 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=23&issue=4&spage=279.

Dais:2021:IGD

[Dai21] Photis Dais. Impact of Gibbs’ and Duhem’s approaches to thermodynamics on the development of chemical thermodynamics. Archive for History of Exact Sciences, 75(2):175–248, March 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Dale:1982:BLE

[Dal82] A. I. Dale. Bayes or Laplace? An examination of the origin and early applications of Bayes’ Theorem. Archive for History of Exact Sciences, 27(1):23–47, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=1&spage=23.

Dale:1986:NDR

[Dal86] A. I. Dale. A newly-discovered result of Thomas Bayes. Archive for History of Exact Sciences, 35(2):101–113, June 1986. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=35&issue=2&spage=101.

Dalmedico:2001:HEM

[Dal01] Amy Dahan Dalmédico. History and epistemology of models: Meteorology (1946–1963) as a case study. Archive for History of Exact Sciences, 55(5):395–422, April 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=5&spage=395.

Martins:1997:BCU

[dAM97] Roberto de Andrade Martins. Becquerel and the choice of uranium compounds. Archive for History of Exact Sciences, 51(1): 67–81, March 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=1&spage=67.

Darrigol:1993:ERG

[Dar93] Olivier Darrigol. The electrodynamic revolution in Germany as documented by early German expositions of “Maxwell’s theory”. Archive for History of Exact Sciences, 45(3):189–280, September 1993. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL http://www.jstor.org/stable/
41133949;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=45&issue=3&spage=
189.

Darrigol:2002:BHE

[Dar02] Olivier Darrigol. Between hydrodynamics and elasticity theory: The first five births of the Navier–Stokes equation. Archive for History of Exact Sciences, 56(2):95–150, January 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=2&spage=95.

Darrigol:2003:SHE

[Dar03] Olivier Darrigol. The spirited horse, the engineer, and the mathematician: Water waves in nineteenth-century hydrodynamics. Archive for History of Exact Sciences, 58(1):21–95, November 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
58&issue=1&spage=21.

Darrigol:2007:AOH

[Dar07] Olivier Darrigol. The acoustic origins of harmonic analysis. Archive for History of Exact Sciences, 61(4):343–424, July 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=4&spage=343.

Darrigol:2020:DNS

[Dar20] Olivier Darrigol. Deducing Newton’s second law from relativity principles: a forgotten history. Archive for History of Exact Sciences, 74(1):1–43, January 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Dauben:1971:TBG

[Dau71] Joseph W. Dauben. The trigonometric background to Georg Cantor’s theory of sets. Archive for History of Exact Sciences, 7(3):181–216, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=7&issue=3&spage=181.

Dauben:2008:ESS

[Dau08a] Joseph W. Dauben. Erratum: Suan Shu Shu: a book on numbers and computations: English translation with commentary. Archive for History of Exact Sciences, 62(3):347, May 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=3&spage=347. See [Dau08b].

Dauben:2008:SSS

[Dau08b] Joseph W. Dauben. Suan Shu Shu: a book on numbers and computations: English translation with commentary. Archive for History of Exact Sciences, 62(2):91–178, March 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

62&issue=2&spage=91. Translated from the Chinese and with commentary by Joseph W. Dauben. See erratum [Dau08a].

Davies:1986:CRB

[Dav86] Mansel Davies. Charles Rugeley Bury and his contributions to physical chemistry. Archive for History of Exact Sciences, 36(1): 75–90, March 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
36&issue=1&spage=75.

Davies:1990:CRB

[Dav90] Mansel Davies. C. R. Bury, L. Vegard, and the electronic interpretation of the periodic table. A note. Archive for History of Exact Sciences, 41(2):185–187, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=41&issue=2&spage=185.

Davis:2003:MAL

[Dav03] A. E. L. Davis. The mathematics of the area law: Kepler’s successful proof in Epitome Astronomiae Copernicanae (1621). Archive for History of Exact Sciences, 57(5):355– 393, July 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=5&spage=355.

deClark:2016:DMP

[dC16] Sybil G. de Clark. The dimensions of the magnetic pole: a controversy at the heart of early dimensional analysis. Archive for History of Exact Sciences, 70(3):293–324, May 2016. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
015-0168-6.

DeYoung:2012:FAR

[De 12] Gregg De Young. Further adventures of the Rome 1594 Arabic redaction of Euclid’s Elements. Archive for History of Exact Sciences, 66(3):265–294, May 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=3&spage=265.

DeRisi:2016:DEA

[De 16] Vincenzo De Risi. The development of Euclidean axiomatics. Archive for History of Exact Sciences, 70(6):591–676, November 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00407-015-0173-9.pdf.

Deakin:1981:DLT

[Dea81] Michael A. B. Deakin. The development of the Laplace transform, 1737–1937 I. Euler to Spitzer, 1737–1880. Archive for History of Exact Sciences, 25(4):343–390, December 1981. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=25&issue=4&spage=343.

Deakin:1982:DLT

[Dea82] Michael A. B. Deakin. The development of the Laplace Transform, 1737–1937 II. Poincaré to Doetsch, 1880–1937. Archive for History of Exact Sciences, 26(4):351–381, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
26&issue=4&spage=351.

Deakin:1985:EII

[Dea85] Michael A. B. Deakin. Euler’s invention of integral transforms. Archive for History of Exact Sciences, 33(4):307–319, December 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
33&issue=4&spage=307.

Deakin:1988:NCA

[Dea88] Michael A. B. Deakin. Nineteenth century anticipations of modern theory of dynamical systems. Archive for History of Exact Sciences, 39(2):183–194, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=2&spage=183.

Deakin:1992:ALT

[Dea92] Michael A. B. Deakin. The ascendancy of the Laplace transform and how it came about. Archive for History of Exact Sciences, 44(3):265–286, September 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=3&spage=265.

Detlefsen:1976:CRN

[DEHY76] Michael Detlefsen, Douglas K. Erlandson, J. Clark Heston, and Charles M. Young. Computation with Roman numerals. Archive for History of Exact Sciences, 15(2):141–148, June 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
15&issue=2&spage=141.

Delsedime:1970:LNA

[Del70] Piero Delsedime. L’infini numérique dans l’Arénaire d’Archimède. (French) [Numerical infinity in the Arénaire of Archimedes]. Archive for History of Exact Sciences, 6(5):345–359, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=5&spage=345.

DelCentina:2008:UMS

[Del08] Andrea Del Centina. Unpublished manuscripts of Sophie Germain and a revaluation of her work on Fermat’s Last Theorem. Archive for History of Exact Sciences, 62(4):349–392, July 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=4&spage=349.

Delcourt:2011:AGH

[Del11] Jean Delcourt. Analyse et géométrie, histoire des courbes gauches De Clairaut à Darboux. (French) [Analysis and geometry: history of ruled curves De Clairaut to Darboux]. Archive for History of Exact Sciences, 65(3):229–293, May 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=3&spage=229.

DelCentina:2016:KSC

[Del16a] Andrea Del Centina. On Kepler’s system of conics in Astronomiae pars optica. Archive for History of Exact Sciences, 70(6): 567–589, November 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-016-0175-2.

Centina:2016:PPL

[Del16b] Andrea Del Centina. Poncelet’s porism: a long story of renewed discoveries, I. Archive for History of Exact Sciences, 70(1):1–122, January 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-015-0163-y.

DelCentina:2016:PPL

[Del16c] Andrea Del Centina. Poncelet’s porism: a long story of renewed discoveries, II. Archive for History of Exact Sciences, 70 (2):123–173, March 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0164-x.

DelCentina:2020:PMH

[Del20] Andrea Del Centina. Pascal’s mystic hexagram, and a conjectural restoration of his lost treatise on conic sections. Archive for History of Exact Sciences, 74(5):469–521, September 2020. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

DelCentina:2022:BPJ

[Del22a] Andrea Del Centina. Brianchon and Poncelet’s joint memoir, the nine-point circle, and beyond. Archive for History of Exact Sciences, 76(4):363–390, July 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00286-7.

DelCentina:2022:CTT

[Del22b] Andrea Del Centina. Carnot’s theory of transversals and its applications by Servois and Brianchon: the awakening of synthetic geometry in France. Archive for History of Exact Sciences, 76 (1):45–128, January 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00276-1.

DelCentina:2022:DCI

[Del22c] Andrea Del Centina. Desargues’s concepts of involution and transversal, their origin, and possible sources of inspiration. Archive for History of Exact Sciences, 76(6):573–622, November 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-022-00296-5.

Demidov:1982:SPD

[Dem82] S. S. Demidov. The study of partial differential equations of the first order in the 18^{th} and 19^{th} centuries. Archive for History of Exact Sciences, 26(4):325–350, December 1982. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=26&issue=4&spage=325.

Demidov:1983:HTL

[Dem83] S. S. Demidov. On the history of the theory of linear differential equations. Archive for History of Exact Sciences, 28(4):369–387, December 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=4&spage=369.

Centina:2012:CBS

[DF12] Andrea Del Centina and Alessandra Fiocca. The correspondence between Sophie Germain and Carl Friedrich Gauss. Archive for History of Exact Sciences, 66(6):585–700, November 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
66&issue=6&spage=585.

DelCentina:2018:MTN

[DF18] Andrea Del Centina and Alessandra Fiocca. “A masterly though neglected work”, Boscovich’s treatise on conic sections. Archive for History of Exact Sciences, 72(4):453–495, July 2018. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Centina:2020:BEB

[DF20] Andrea Del Centina and Alessandra Fiocca. Borelli’s edition of books V-VII of Apollonius’s Conics, and Lemma 12 in Newton’s Principia. Archive for History of Exact Sciences, 74 (3):255–279, May 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-019-00244-w.

Dragovic:2020:PPF

[DG20] Vladimir Dragović and Irina Goryuchkina. Polygons of Petrović and Fine, algebraic ODEs, and contemporary mathematics. Archive for History of Exact Sciences, 74(6):523–564, November 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Dhombres:1986:QAH

[Dho86] J. Dhombres. Quelques aspects de l’histoire des équations fonctionnelles liés à l’évolution du concept de fonction. (French) [Some aspects of the history of functional equations connected to the concept of a function]. Archive for History of Exact Sciences, 36(2):91–181, June 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=36&issue=2&spage=91.

Dias:1994:WSS

[Dia94] Penha Maria Cardoso Dias. “Will someone say exactly what the H-theorem proves?” A study of Burbury’s Condition A and Maxwell’s Proposition II. Archive for History of Exact Sciences, 46(4):341–366, December 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=46&issue=4&spage=341.

Dias:1999:ESB

[Dia99] Penha Maria Cardoso Dias. Euler’s “harmony” between the principles of “rest” and “least action”: The conceptual making of analytical mechanics. Archive for History of Exact Sciences, 54(1):67–86, April 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=1&spage=67.

Dill:1992:KTR

[Dil92] Ellis Harold Dill. Kirchhoff’s theory of rods. Archive for History of Exact Sciences, 44(1):1–23, March 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=1&spage=1.

Duncan:2007:VURa

[DJ07a] Anthony Duncan and Michel Janssen. On the verge of Umdeutung [reinterpretation] in Minnesota: Van Vleck and the Correspondence Principle. Part one. Archive for History of Exact Sciences, 61(6):553–624, November 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=6&spage=553.

Duncan:2007:VURb

[DJ07b] Anthony Duncan and Michel Janssen. On the verge of Umdeutung [reinterpretation] in Minnesota: Van Vleck and the Correspondence Principle. Part two. Archive for History of Exact Sciences, 61(6):625–671, November 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=61&issue=6&spage=625.

deJong:2019:SBPa

[dJ19a] Teije de Jong. A study of Babylonian planetary theory I. The outer planets. Archive for History of Exact Sciences, 73(1):1–37, January 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s00407-018-0216-0.pdf.

deJong:2019:SBPb

[dJ19b] Teije de Jong. A study of Babylonian planetary theory II. The planet Venus. Archive for History of Exact Sciences, 73 (4):309–333, July 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00407-019-00224-0.pdf.

deJong:2021:SBP

[dJ21] Teije de Jong. A study of Babylonian planetary theory III. The planet Mercury. Archive for History of Exact Sciences, 75(5): 491–522, September 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-020-00269-6.

deJong:2020:BOU

[dJH20] Teije de Jong and Hermann Hunger. Babylonian observations of a unique planetary configuration. Archive for History of Exact Sciences, 74(6):587–603, November 2020. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-020-00252-1.pdf.

deOliveira:1988:ACC

[dO88] A. J. Franco de Oliveira. Anastácio da Cunha and the concept of convergent series. Archive for History of Exact Sciences, 39(1):1–12, March 1988. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=1&spage=1.

Dobson:1998:NPR

[Dob98] Geoffrey. J. Dobson. Newton’s problems with rigid body dynamics in the light of his treatment of the precession of the equinoxes. Archive for History of Exact Sciences, 53(2):125– 145, June 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=2&spage=125.

Dobson:1999:ACI

[Dob99a] Geoffrey J. Dobson. Against Chandrasekhar’s interpretation of Newton’s treatment of the precession of the equinoxes. Archive for History of Exact Sciences, 53(6):577–597, January 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=6&spage=577.

Dobson:1999:NER

[Dob99b] Geoffrey J. Dobson. Newton’s errors with the rotational motion of fluids. Archive for History of Exact Sciences, 54(3):243– 254, August 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=3&spage=243.

Dobson:2001:LPB

[Dob01] Geoffrey J. Dobson. On Lemmas 1 and 2 to Proposition 39 of Book 3 of Newton’s Principia. Archive for History of Exact Sciences, 55(4):345–363, January 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=4&spage=345.

Domingues:2023:GAA

[Dom23] João Caramalho Domingues. Geometry and analysis in Anastácio da Cunha’s calculus. Archive for History of Exact Sciences, 77(6):579–600, November 2023. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://

link.springer.com/article/10.1007/s00407-023-00313-1.

Doncel:1991:PHC

[Don91] Manuel G. Doncel. On the process of Hertz’s conversion to Hertzian waves. Archive for History of Exact Sciences, 43(1): 1–27, March 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
43&issue=1&spage=1.

Dong:2021:HMR

[Don21] Hao Dong. Hobbes’s model of refraction and derivation of the sine law. Archive for History of Exact Sciences, 75(3):323–348, May 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Dostrovsky:1975:EVT

[Dos75] Sigalia Dostrovsky. Early vibration theory: Physics and music in the seventeenth century. Archive for History of Exact Sciences, 14(3):169–218, September 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=14&issue=3&spage=169.

Dias:1995:CIC

[DPC95] Penha Maria Cardoso Dias, Simone Pinheiro Pinto, and Deisemar Hollanda Cassiano. The conceptual import of Carnot’s theorem to the discovery of the entropy. Archive for History of Exact Sciences, 49(2):135–161, June 1995. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=49&issue=2&spage=135.

Dumbaugh:2012:CEA

[DS12] Della Dumbaugh and Joachim Schwermer. The collaboration of Emil Artin and George Whaples: Artin’s mathematical circle extends to America. Archive for History of Exact Sciences, 66(5):465–484, September 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=5&spage=465.

Ducheyne:2011:TUG

[Duc11] Steffen Ducheyne. Testing universal gravitation in the laboratory, or the significance of research on the mean density of the earth and big G, 1798–1898: changing pursuits and long-term methodological-experimental continuity. Archive for History of Exact Sciences, 65(2):181–227, March 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=2&spage=181.

Dugac:1973:EAK

[Dug73] Pierre Dugac. Eléments d’analyse de Karl Weierstrass. (French) [Elements of analysis of Karl Weierstrass]. Archive for History of Exact Sciences, 10(1–2):41–176, January 1973. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=10&issue=1&spage=41.

Dugac:1976:NDV

[Dug76] Pierre Dugac. Notes et documents sur la vie et l’oeuvre de René Baire. (French) [Notes and documents on the life and works of René Baire]. Archive for History of Exact Sciences, 15(4):297–383, December 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=15&issue=4&spage=297.

Duke:2002:ABA

[Duk02a] Dennis W. Duke. Associations between the ancient star catalogues. Archive for History of Exact Sciences, 56(5):435– 450, July 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=5&spage=435.

Duke:2002:HCS

[Duk02b] Dennis W. Duke. Hipparchus’ coordinate system. Archive for History of Exact Sciences, 56(5):427–433, July 2002. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=56&issue=5&spage=427.

Duke:2005:EIM

[Duk05a] Dennis Duke. The Equant in India: The mathematical basis of ancient Indian planetary models. Archive for History of Exact Sciences, 59(6):563–576, October 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=6&spage=563.

Duke:2005:PTO

[Duk05b] Dennis Duke. Ptolemy’s treatment of the outer planets. Archive for History of Exact Sciences, 59(2):169–187, January 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=2&spage=169.

Duke:2007:SLA

[Duk07] Dennis W. Duke. The second lunar anomaly in ancient Indian astronomy. Archive for History of Exact Sciences, 61(2):147– 157, March 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=2&spage=147.

Duke:2008:MML

[Duk08] Dennis W. Duke. Mean motions and longitudes in Indian astronomy. Archive for History of Exact Sciences, 62(5):489–509, September 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=5&spage=489.

Duke:2009:MMP

[Duk09] Dennis Duke. Mean motions in Ptolemy’s planetary hypotheses. Archive for History of Exact Sciences, 63(6):635–654, November 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
63&issue=6&spage=635.

Duke:2010:GAB

[Duk10] Dennis Duke. Greek angles from Babylonian numbers. Archive for History of Exact Sciences, 64(3):375–394, May 2010. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=64&issue=3&spage=375.

Dutka:1981:IBF

[Dut81] Jacques Dutka. The incomplete beta function — a historical profile. Archive for History of Exact Sciences, 24(1):11– 29, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
24&issue=1&spage=11.

Dutka:1982:WPB

[Dut82] Jacques Dutka. Wallis’s product, Brouncker’s continued fraction, and Leibniz’s series. Archive for History of Exact Sciences, 26(2):115–126, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=26&issue=2&spage=115.

Dutka:1984:EHH

[Dut84] Jacques Dutka. The early history of the hypergeometric function. Archive for History of Exact Sciences, 31(1):15–34, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
31&issue=1&spage=15.

Dutka:1985:PRF

[Dut85] Jacques Dutka. On the problem of random flights. Archive for History of Exact Sciences, 32(3–4):351–375, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=3&spage=351.

Dutka:1986:SRT

[Dut86] Jacques Dutka. On square roots and their representations. Archive for History of Exact Sciences, 36(1):21–39, March 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
36&issue=1&spage=21.

Dutka:1988:SPP

[Dut88] Jacques Dutka. On the St. Petersburg paradox. Archive for History of Exact Sciences, 39(1):13–39, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=39&issue=1&spage=13.

Dutka:1990:RAM

[Dut90] Jacques Dutka. Robert Adrain and the method of least squares. Archive for History of Exact Sciences, 41(2):171– 184, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=2&spage=171.

Dutka:1991:EHF

[Dut91] Jacques Dutka. The early history of the factorial function. Archive for History of Exact Sciences, 43(3):225–249, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
43&issue=3&spage=225.

Dutka:1993:SME

[Dut93] Jacques Dutka. ”Eratosthenes’ measurement of the Earth reconsidered”. Archive for History of Exact Sciences, 46(1):55– 66, March 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=1&spage=55.

Dutka:1995:EHB

[Dut95] Jacques Dutka. On the early history of Bessel functions. Archive for History of Exact Sciences, 49(2):105–134, June 1995. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=49&issue=2&spage=105.

Dutka:1996:GPD

[Dut96a] Jacques Dutka. On Gauss’ priority in the discovery of the method of least squares. Archive for History of Exact Sciences, 49(4):355–370, December 1996. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=49&issue=4&spage=355.

Dutka:1996:SSD

[Dut96b] Jacques Dutka. On the summation of some divergent series of Euler and the zeta functions. Archive for History of Exact Sciences, 50(2):187–200, August 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=2&spage=187.

Dutka:1997:NKE

[Dut97] Jacques Dutka. A note on ‘Kepler’s equation’. Archive for History of Exact Sciences, 51(1):59–65, March 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=51&issue=1&spage=59.

Earman:2001:LCR

[Ear01] John Earman. Lambda: The constant that refuses to die. Archive for History of Exact Sciences, 55(3):189–220, January 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=3&spage=189.

Eastwood:1970:MDL

[Eas70] Bruce S. Eastwood. Metaphysical derivations of a law of refraction: Damianos and Grosseteste. Archive for History of Exact Sciences, 6(3):224–236, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=3&spage=224.

Evans:2019:BST

[EC19] James Evans and Christián C. Carman. Babylonian solar theory on the Antikythera mechanism. Archive for History of Exact Sciences, 73(6):619–659, November 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Eckert:2002:EFS

[Eck02] Michael Eckert. Euler and the Fountains of Sanssouci. Archive for History of Exact Sciences, 56(6):451–468, November 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=6&spage=451.

Eckert:2021:PFG

[Eck21] Michael Eckert. Pipe flow: a gateway to turbulence. Archive for History of Exact Sciences, 75(3):249–282, May 2021. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-020-00263-y.pdf.

Eckert:2024:EPH

[Eck24] Michael Eckert. The efflux problem: how hydraulics became divorced from hydrodynamics. Archive for History of Exact Sciences, 78(2):127–152, March 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-023-00320-2.

Esteve:2009:EBI

[ED09] Ma Rosa Massa Esteve and Amadeu Delshams. Euler’s beta integral in Pietro Mengoli’s works. Archive for History of Exact Sciences, 63(3):325–356, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Edwards:1975:BKP

[Edw75] Harold M. Edwards. The background of Kummer’s proof of Fermat’s last theorem for regular primes. Archive for History of Exact Sciences, 14(3):219–236, September 1975. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=14&issue=3&spage=219. See

postscript [Edw77].

Edwards:1977:PBK

[Edw77] Harold M. Edwards. Postscript to “The background of Kummer’s proof … ”. Archive for History of Exact Sciences, 17(4): 381–394, July 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
17&issue=4&spage=381. See [Edw75].

Edwards:1980:GIT

[Edw80] Harold M. Edwards. The genesis of ideal theory. Archive for History of Exact Sciences, 23(4):321–378, December 1980. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=23&issue=4&spage=321.

Edwards:1990:NGT

[Edw90] Harold M. Edwards. A note on Galois theory. Archive for History of Exact Sciences, 41(2):163–169, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=41&issue=2&spage=163.

Earman:1978:EHT

[EG78] John Earman and Clark Glymour. Einstein and Hilbert: Two months in the history of General Relativity. Archive for History of Exact Sciences, 19(3):291–308, September 1978. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=19&issue=3&spage=291.

Ehberger:2022:LDT

[Ehb22] Markus Ehberger. “The language of Dirac’s theory of radiation”: the inception and initial reception of a tool for the quantum field theorist. Archive for History of Exact Sciences, 76(6): 531–571, November 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00293-8. See correction

[Ehb23].

Ehberger:2023:CLD

[Ehb23] Markus Ehberger. Correction to: “The language of Dirac’s theory of radiation”: the inception and initial reception of a tool for the quantum field theorist. Archive for History of Exact Sciences, 77(1):121–122, January 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00300-y. See

[Ehb22].

Ehrlich:2006:RNA

[Ehr06] Philip Ehrlich. The rise of non-Archimedean mathematics and the roots of a misconception I: the emergence of non-Archimedean systems of magnitudes. Archive for History of Exact Sciences, 60(1):1–121, January 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=60&issue=1&spage=1.

Eisenstaedt:1982:HSS

[Eis82] J. Eisenstaedt. Histoire et singularités de la solution de Schwarzschild (1915–1923). (French) [History and singularities of Schwarzschild’s solution (1915–1923)]. Archive for History of Exact Sciences, 27(2):157–198, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=2&spage=157.

Eisenstaedt:1986:RGE

[Eis86] Jean Eisenstaedt. La relativité générale à l’étiage: 1925– 1955. (French) [General Relativity at low-water level: 1925– 1955]. Archive for History of Exact Sciences, 35(2):115–185, June 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=2&spage=115.

Eisenstaedt:1987:TIS

[Eis87] J. Eisenstaedt. Trajectoires et impasses de la solution de Schwarzschild. (French) [Trajectories and blocks in Schwarzschild’s solution]. Archive for History of Exact Sciences, 37(4):275–357, December 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
37&issue=4&spage=275.

Eisenstaedt:1991:IGP

[Eis91] J. Eisenstaedt. De l’influence de la gravitation sur la propagation de la lumière en théorie newtonienne. L’archéologie des trous noirs. (French) [On the influence of gravitation on the propagation of light in Newtonian theory. The archaelogy of black holes]. Archive for History of Exact Sciences, 42(4):315–386, December 1991. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=42&issue=4&spage=315.

Eisenstaedt:1996:LBN

[Eis96] Jean Eisenstaedt. L’optique balistique newtonienne à l’épreuve des satellites de Jupiter. (French) [Newtonian ballistic optics tested by Jupiter’s satellites]. Archive for History of Exact Sciences, 50(2):117–156, August 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=2&spage=117.

Edwards:1982:DSB

[ENP82] Harold Edwards, Olaf Neumann, and Walter Purkert. Dedekinds “Bunte Bemerkungen” zu Kroneckers “Grundzüge”. (German) [Dedekind’s “multicolored remark” on Kronecker’s “Outlines”]. Archive for History of Exact Sciences, 27(1):49–85, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=1&spage=49.

Epple:1998:TMS

[Epp98] Moritz Epple. Topology, matter, and space, I: Topological notions in 19th-century natural philosophy. Archive for History of Exact Sciences, 52(4):297–392, February 1998. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=52&issue=4&spage=297.

Esquisabel:2021:FPI

[EQ21] Oscar M. Esquisabel and Federico Raffo Quintana. Fiction, possibility and impossibility: three kinds of mathematical fictions in Leibniz’s work. Archive for History of Exact Sciences, 75(6): 613–647, November 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00277-0.

Evans:1976:SEG

[Eva76] Gillian R. Evans. The “sub-euclidean” geometry of the earlier middle ages, up to the mid-twelfth century. Archive for History of Exact Sciences, 16(2):105–118, June 1976. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=2&spage=105.

Ewald:1969:MMC

[Ewa69] P. P. Ewald. The myth of myths; comments on P. Forman’s paper on “The discovery of the diffraction of X-rays in crystals”. Archive for History of Exact Sciences, 6(1):72–81, January 1969. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=6&issue=1&spage=

72. See [For69].

Feigenbaum:1985:BTM

[Fei85] L. Feigenbaum. Brook Taylor and the method of increments. Archive for History of Exact Sciences, 34(1–2):1–140, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
34&issue=1&spage=1.

Fenster:1998:LED

[Fen98] Della Dumbaugh Fenster. Leonard Eugene Dickson and his work in the Arithmetics of Algebras. Archive for History of Exact Sciences, 52(2):119–159, February 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=52&issue=2&spage=119.

Ferreiros:1996:TLE

[Fer96] José Ferreirós. Traditional logic and the early history of sets, 1854–1908. Archive for History of Exact Sciences, 50(1):5– 71, March 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
50&issue=1&spage=5.

Ferraro:1999:FMD

[Fer99] Giovanni Ferraro. The first modern definition of the sum of a divergent series: An aspect of the rise of 20th century mathematics. Archive for History of Exact Sciences, 54(2):101– 135, June 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=2&spage=101.

Ferraro:2007:FAG

[Fer07] Giovanni Ferraro. The foundational aspects of Gauss’s work on the hypergeometric, factorial and digamma functions. Archive for History of Exact Sciences, 61(5):457–518, September 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=5&spage=457.

Ferreiro:2014:MIA

[Fer14] Larrie D. Ferreiro. The mutual influence of aircraft aerodynamics and ship hydrodynamics in theory and experiment. Archive for History of Exact Sciences, 68(2):241–263, March 2014. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0129-x.

Friedman:2023:FTT

[FG23] Michael Friedman and David Garber. On fluidity of the textual transmission in Abraham bar Hiyya’s H .ibbur ha-Meshih.ah ve-ha-Tishboret. Archive for History of Exact Sciences, 77(2): 123–174, March 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00297-4.

Fenaroli:1981:GSS

[FGP81] Giuseppina Fenaroli, Ubaldo Garibaldi, and M. Antonietta Penco. Giochi, scommesse sulla vita, tabelle di mortalità: Nascita del calcolo probabilistico, statistica e teoria delle popolazioni. (Italian) [Giochi, bets on life, mortality tables: the birth of the calculus of probability, statistics, and population theory]. Archive for History of Exact Sciences, 25(4):329–341, December 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
25&issue=4&spage=329.

Fincke:2021:BSA

[FHR21] Jeanette C. Fincke, Wayne Horowitz, and Eshbal Ratzon. BM 76829: a small astronomical fragment with important implications for the Late Babylonian Astronomy and the Astronomical Book of Enoch. Archive for History of Exact Sciences, 75(3): 349–368, May 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Field:1984:LAJ

[Fie84] J. V. Field. A Lutheran astrologer: Johannes Kepler. Archive for History of Exact Sciences, 31(3):189–272, September 1984. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=31&issue=3&spage=189.

Field:1997:RAP

[Fie97] J. V. Field. Rediscovering the Archimedean Polyhedra: Piero della Francesca, Luca Pacioli, Leonardo da Vinci, Albrecht Dürer, Daniele Barbaro, and Johannes Kepler. Archive for History of Exact Sciences, 50(3–4):241–289, September 1997. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=50&issue=3&spage=241.

Fisher:1966:DMT

[Fis66] Charles S. Fisher. The death of a mathematical theory: a study in the sociology of knowledge. Archive for History of Exact Sciences, 3(2):137–159, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=3&issue=2&spage=137.

Fisher:1981:IIQ

[Fis81] Gordon Fisher. The infinite and infinitesimal quantities of du Bois–Reymond and their reception. Archive for History of Exact Sciences, 24(2):101–163, June 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=24&issue=2&spage=101.

Fisch:1999:MPT

[Fis99] Menachem Fisch. The making of Peacock’s Treatise on Algebra: a case of creative indecision. Archive for History of Exact Sciences, 54(2):137–179, June 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=2&spage=137.

Fraser:2023:CTJ

[FN23] Craig Fraser and Michiyo Nakane. Canonical transformations from Jacobi to Whittaker. Archive for History of Exact Sciences, 77(3):241–343, May 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00303-9.

Forman:1969:DDX

[For69] Paul Forman. The discovery of the diffraction of X-rays by crystals; a critique of the myths. Archive for History of Exact Sciences, 6(1):38–71, January 1969. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=6&issue=1&spage=

38. See comments [Ewa69].

Fowler:1980:BIE

[Fow80] D. H. Fowler. Book II of Euclid’s Elements and a pre-Eudoxan theory of ratio. Archive for History of Exact Sciences, 22(1–2): 5–36, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
22&issue=1&spage=5.

Fowler:1981:ARE

[Fow81] D. H. Fowler. Anthyphairetic ratio and Eudoxan proportion. Archive for History of Exact Sciences, 24(2):69–72, June 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
24&issue=2&spage=69.

Fowler:1982:BIE

[Fow82] D. H. Fowler. Book II of Euclid’s Elements and a pre-Eudoxan theory of ratio. part 2: Sides and diameters. Archive for History of Exact Sciences, 26(3):193–209, September 1982. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=26&issue=3&spage=193.

Fowler:1991:ATU

[Fow91] D. H. Fowler. An approximation technique, and its use by Wallis and Taylor. Archive for History of Exact Sciences, 41(3):189–233, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=41&issue=3&spage=189.

Ferraro:2012:LTA

[FP12] Giovanni Ferraro and Marco Panza. Lagrange’s theory of analytical functions and his ideal of purity of method. Archive for History of Exact Sciences, 66(2):95–197, March 2012. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=66&issue=2&spage=95.

Fraser:1983:JLE

[Fra83] Craig Fraser. J. L. Lagrange’s early contributions to the principles and methods of mechanics. Archive for History of Exact Sciences, 28(3):197–241, September 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=28&issue=3&spage=197.

Fraser:1985:JLC

[Fra85] Craig Fraser. J. L. Lagrange’s changing approach to the foundations of the calculus of variations. Archive for History of Exact Sciences, 32(2):151–191, June 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=32&issue=2&spage=151.

Fraser:1989:CAA

[Fra89] Craig G. Fraser. The calculus as algebraic analysis: Some observations on mathematical analysis in the 18^{th} century. Archive for History of Exact Sciences, 39(4):317–335, December 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
39&issue=4&spage=317.

Fraser:1994:OEV

[Fra94] Craig G. Fraser. The origins of Euler’s variational calculus. Archive for History of Exact Sciences, 47(2):103–141, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
47&issue=2&spage=103.

Franchella:1997:ODK

[Fra97] Miriam Franchella. On the origins of Dénes König’s infinity lemma. Archive for History of Exact Sciences, 51(1):3– 27, March 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=1&spage=3.

Franklin:1998:SPE

[Fra98] Allan Franklin. Selectivity and the production of experimental results “Any fool can take data. It’s taking good data that counts.” E. Commins (private communication). Archive for History of Exact Sciences, 53(5):399–485, November 1998. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=53&issue=5&spage=399.

Franklin:2004:DMA

[Fra04] Allan Franklin. Doing much about nothing. Archive for History of Exact Sciences, 58(4):323–379, May 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=4&spage=323.

Frank:2013:MTC

[Fra13] Martin Frank. Mathematics, technics, and courtly life in Late Renaissance Urbino. Archive for History of Exact Sciences, 67 (3):305–330, May 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-013-0113-5.

Freudenthal:1971:DCP

[Fre71] H. Freudenthal. Did Cauchy plagiarize Bolzano? Archive for History of Exact Sciences, 7(5):375–392, January 1971.

CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=7&issue=5&spage=
375.

Frenkel:1974:YJI

[Fre74] Victor J. Frenkel. Yakov (James) Il’ich Frenkel (1894–1952): Materials for his scientific biography. Archive for History of Exact Sciences, 13(1):1–26, March 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=13&issue=1&spage=1.

Freudenthal:1977:WAW

[Fre77] Hans Freudenthal. What is algebra and what has it been in history? Archive for History of Exact Sciences, 16(3):189–200, September 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
16&issue=3&spage=189.

Freudenthal:1991:NGI

[Fre91] Hans Freudenthal. Nichteuklidische Geometrie im Altertum?. (German) [Non-Euclidean geometry in antiquity]. Archive for History of Exact Sciences, 43(3):189–197, September 1991. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=43&issue=3&spage=189.

Friberg:2014:GDP

[Fri14] Jöran Friberg. Geometric division problems, quadratic equations, and recursive geometric algorithms in Mesopotamian mathematics. Archive for History of Exact Sciences, 68(1):1–34, January 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-013-0122-4.

Friberg:2019:TTY

[Fri19] Jöran Friberg. Three thousand years of sexagesimal numbers in Mesopotamian mathematical texts. Archive for History of Exact Sciences, 73(2):183–216, March 2019. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-019-00221-3.pdf.

Friedman:2020:HNC

[Fri20] Michael Friedman. How to notate a crossing of strings? On Modesto Dedò’s notation of braids. Archive for History of Exact Sciences, 74(4):281–312, July 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-019-00238-8.

Froman:1994:HBT

[Frö94] Per Olof Fröman. Historical background of the Tachyon concept. Archive for History of Exact Sciences, 48(3–4):373–380, September 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
48&issue=3&spage=373.

Fenster:2005:DCA

[FS05] Della D. Fenster and Joachim Schwermer. A delicate collaboration: Adrian Albert and Helmut Hasse and the Principal Theorem in Division Algebras in the early 1930’s. Archive for History of Exact Sciences, 59(4):349–379, May 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=4&spage=349.

Fenster:2007:BCF

[FS07] Della D. Fenster and Joachim Schwermer. Beyond class field theory: Helmut Hasse’s arithmetic in the theory of algebras in early 1931. Archive for History of Exact Sciences, 61(5):425–456, September 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=5&spage=425.

Figari:2013:ECT

[FT13] Rodolfo Figari and Alessandro Teta. Emergence of classical trajectories in quantum systems: the cloud chamber problem in the analysis of Mott (1929). Archive for History of Exact Sciences, 67(2):215–234, March 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-012-0111-z.

Galuzzi:1980:PDT

[Gal80] Massimo Galuzzi. Il problema delle tangenti nella “Géométrie” di Descartes. (Italian) [The problem of tangents in Descartes’s Geometry]. Archive for History of Exact Sciences, 22(1–2):37– 51, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
22&issue=1&spage=37.

Galuzzi:2001:GNA

[Gal01] Massimo Galuzzi. Galois’ note on the approximative solution of numerical equations (1830). Archive for History of Exact Sciences, 56(1):29–37, November 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=56&issue=1&spage=29.

Galeriu:2017:ECH

[Gal17] Calin Galeriu. Electric charge in hyperbolic motion: the early history. Archive for History of Exact Sciences, 71(4):363–378, July 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Gaarding:1989:HMD

[Går89a] Lars Gårding. History of the mathematics of double refraction. Archive for History of Exact Sciences, 40(4):355–385, December 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
40&issue=4&spage=355.

Gario:1989:RSS

[Gar89b] Paola Gario. Resolution of singularities of surfaces by P. Del Pezzo. A mathematical controversy with C. Segre. Archive for History of Exact Sciences, 40(3):247–274, September 1989. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=40&issue=3&spage=247.

Gario:1991:SGS

[Gar91] Paola Gario. Singolarità e geometria sopra una superficie nella corrispondenza di C. Segre a G. Castelnuovo. (Italian) [Singularity and geometry over a surface in the correspondence of C. Segre to G. Castelnuovo]. Archive for History of Exact Sciences, 43(2):145–188, June 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=43&issue=2&spage=145.

Goldstein:1991:IDO

[GB91] Bernard R. Goldstein and Alan C. Bowen. The introduction of dated observations and precise measurement in Greek astronomy. Archive for History of Exact Sciences, 43(2):93– 132, June 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
43&issue=2&spage=93.

Goldstein:2004:PBC

[GC04] Bernard R. Goldstein and José Chabás. Ptolemy, Bianchini, and Copernicus: Tables for planetary latitudes. Archive for History of Exact Sciences, 58(5):453–473, July 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=5&spage=453.

Goldstein:2017:AAT

[GC17] Bernard R. Goldstein and José Chabás. Analysis of the astronomical tables for 1340 compiled by Immanuel ben Jacob Bonfils. Archive for History of Exact Sciences, 71(1):71–108, January 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-016-0181-4.

Goldstein:2018:NEA

[GC18] Bernard R. Goldstein and José Chabás. New evidence on Abraham Zacut’s astronomical tables. Archive for History of Exact Sciences, 72(1):21–62, January 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Goldstein:2023:JIW

[GC23] Bernard R. Goldstein and José Chabás. Joseph Ibn Waqar and the treatment of retrograde motion in the Middle Ages. Archive for History of Exact Sciences, 77(2):175–199, March 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-022-00301-x.

Goldstein:2024:TRS

[GC24] Bernard R. Goldstein and José Chabás. Tables for the radii of the Sun, the Moon, and the shadow from John of Gmunden to Longomontanus. Archive for History of Exact Sciences, 78 (1):67–86, January 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-023-00318-w.

Gray:2023:JLP

[GCWL23] Jeremy Gray, Joshua L. Cherry, Eric-Jan Wagenmakers, and Alexander Ly. The Jeffreys–Lindley paradox: an exchange. Archive for History of Exact Sciences, 77(4):443–449, July 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00310-4. See [WL23].

Gearhart:1985:EEE

[Gea85] C. A. Gearhart. Epicycles, eccentrics, and ellipses: The predictive capabilities of Copernican planetary models. Archive for History of Exact Sciences, 32(3–4):207–222, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=3&spage=207.

Gearhart:2010:SSS

[Gea10] Clayton A. Gearhart. “astonishing successes” and “bitter disappointment”: The specific heat of hydrogen in quantum theory. Archive for History of Exact Sciences, 64(2):113–202, March 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
64&issue=2&spage=113.

Gerber:1969:GW

[Ger69] Johannes Gerber. Geschichte der Wellenmechanik. (German) [History of wave mechanics]. Archive for History of Exact Sciences, 5(5):349–416, January 1969. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=5&spage=349.

Gericke:1982:VEK

[Ger82] H. Gericke. Zur Vorgeschichte und Entwicklung des Krümmungsbegriffs. (German) [On the prehistory and development of the curvature term]. Archive for History of Exact Sciences, 27(1):1– 21, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=1&spage=1.

Grattan-Guinness:1970:BCS

[GG70] I. Grattan-Guinness. Bolzano, Cauchy and the “new analysis” of the early nineteenth century. Archive for History of Exact Sciences, 6(5):372–400, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=5&spage=372.

Gavroglu:1985:HLT

[GG85] Kostas Gavroglu and George Goudaroulis. From the history of low temperature physics: Prejudicial attitudes that hindered the initial development of superconductivity theory. Archive for History of Exact Sciences, 32(3–4):377–383, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=3&spage=377.

Gomez-Garcia:2021:SBD

[GGHPMR21] Francisco Gómez-García, Pedro J. Herrero-Piñeyro, and Antonio Mellado-Romero. The six books of Diophantus’ Arithmetic increased and reduced to specious: the lost manuscript of Jacques Ozanam (1640–1718). Archive for History of Exact Sciences, 75(5):557–611, September 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://

link.springer.com/article/10.1007/s00407-021-00274-3.

Gluchoff:2000:MUP

[GH00] Alan Gluchoff and Frederick Hartmann. On a “much Underestimated” Paper of Alexander. Archive for History of Exact Sciences, 55(1):1–41, August 2000. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=1&spage=1.

Gillings:1974:RRM

[Gil74] R. J. Gillings. The Recto of the Rhind mathematical papyrus: how did the ancient Egyptian scribe prepare it? Archive for History of Exact Sciences, 12(4):291–298, August 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
12&issue=4&spage=291.

Gillings:1975:WRB

[Gil75] R. J. Gillings. What is the relation between the EMLR and the RMP recto? Archive for History of Exact Sciences, 14(3):159–167, September 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=14&issue=3&spage=159.

Gilain:1991:HDT

[Gil91] Christian Gilain. Sur l’histoire du théorème fondamental de l’algèbre: théorie des équations et calcul intégral. (French) [On the Fundamental Theorem of Algebra: theory of the equations of integral calculus]. Archive for History of Exact Sciences, 42(2):91–136, June 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=42&issue=2&spage=91.

Giovannini:2021:DHF

[Gio21] Eduardo N. Giovannini. David Hilbert and the foundations of the theory of plane area. Archive for History of Exact Sciences, 75(6):649–698, November 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-021-00278-z.

Giovanelli:2024:PPP

[Gio24] Marco Giovanelli. The practice of principles: Planck’s vision of a relativistic general dynamics. Archive for History of Exact Sciences, 78(3):305–360, May 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-024-00326-4.

Gispert:1983:FAF

[Gis83] Hélène Gispert. Sur les fondements de l’analyse en France. (French) [On the basis of analysis in France]. Archive for History of Exact Sciences, 28(1):37–106, March 1983. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=28&issue=1&spage=37.

Gysembergh:2023:PTM

[GJZ^{+}23] Victor Gysembergh, Alexander Jones, Emanuel Zingg, Pascal Cotte, and Salvatore Apicella. Ptolemy’s treatise on the meteoroscope recovered. Archive for History of Exact Sciences, 77 (2):221–240, March 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00302-w.

Ginoux:2012:BOA

[GL12] Jean-Marc Ginoux and René Lozi. Blondel et les oscillations auto-entretenues. (French) [Blondel and self-sustained oscillations]. Archive for History of Exact Sciences, 66(5):485–530, September 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
66&issue=5&spage=485.

Gauthier:2019:YWL

[GL19] Sébastien Gauthier and François Lê. On the youthful writings of Louis J. Mordell on the Diophantine equation y^{2} − k = x^{3}. Archive for History of Exact Sciences, 73(5):427–468, September 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Grasshoff:2017:PPO

[GM17] Gerd Graßhoff and Florian Mittenhuber. Of paths and places: the origin of Ptolemy’s Geography. Archive for History of Exact Sciences, 71(6):483–508, November 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://

link.springer.com/article/10.1007/s00407-017-0194-7.

Goetsch:1968:AB

[Goe68] H. Goetsch. Die Algebra der Babylonier. (German) [The algebra of the Babylonians]. Archive for History of Exact Sciences, 5(2):79–153, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=2&spage=79.

Goldberg:1970:ATE

[Gol70] Stanley Goldberg. The Abraham theory of the electron: The symbiosis of experiment and theory. Archive for History of Exact Sciences, 7(1):7–25, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=7&issue=1&spage=7.

Goldstein:1977:IMT

[Gol77] Bernard R. Goldstein. Ibn Mu^{c}ādh’s treatise on twilight and the height of the atmosphere. Archive for History of Exact Sciences, 17(2):97–118, July 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=2&spage=97.

Goldstein:2003:AZH

[Gol03] Bernard R. Goldstein. An anonymous zij in Hebrew for 1400 A.D.: a preliminary report. Archive for History of Exact Sciences, 57(2):151–171, March 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=57&issue=2&spage=151.

Goulding:2014:THO

[Gou14] Robert Goulding. Thomas Harriot’s optics, between experiment and imagination: the case of Mr Bulkeley’s glass. Archive for History of Exact Sciences, 68(2):137–178, March 2014. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0125-1.

Goulding:2018:BVI

[Gou18] Robert Goulding. Binocular vision and image location before Kepler. Archive for History of Exact Sciences, 72(5):497–546, September 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Gandon:2009:PDL

[GP09] Sébastien Gandon and Yvette Perrin. Le problème de la définition de l’aire d’une surface gauche: Peano et Lebesgue. (French) [The problem of the definition of an unflattenable ruled surface: Peano and Lebesgue]. Archive for History of Exact Sciences, 63(6):665–704, November 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=6&spage=665.

Grant:1961:NOC

[Gra61] Edward Grant. Nicole Oresme and the commensurability or incommensurability of the celestial motions. Archive for History of Exact Sciences, 1(4):420–458, July 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=4&spage=420.

Grant:1965:BGE

[Gra65] Edward Grant. Bradwardine and Galileo: Equality of velocities in the void. Archive for History of Exact Sciences, 2(4):344– 364, April 1965. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
2&issue=4&spage=344.

Gray:1980:ORP

[Gra80] Jeremy J. Gray. Olinde Rodrigues’ paper of 1840 on transformation groups. Archive for History of Exact Sciences, 21(4):375–385, December 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=4&spage=375.

Gray:1984:SRR

[Gra84] J. D. Gray. The shaping of the Riesz representation theorem: a chapter in the history of analysis. Archive for History of Exact Sciences, 31(2):127–187, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=31&issue=2&spage=127.

Graham:2015:PA

[Gra15] Daniel W. Graham. On Philolaus’ astronomy. Archive for History of Exact Sciences, 69(2):217–230, March 2015. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
014-0147-3.

Greenberg:1996:INP

[Gre96] John L. Greenberg. Isaac Newton and the problem of the Earth’s shape. Archive for History of Exact Sciences, 49(4):371–391, December 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
49&issue=4&spage=371.

Grossman:1998:NQA

[Gro98] Nathaniel Grossman. Newton’s quiescence of the apsides and radially-symmetric attractions to a center. Archive for History of Exact Sciences, 52(2):109–117, February 1998. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=52&issue=2&spage=109.

Grupe:2017:SPT

[Gru17] Dirk Grupe. Stephen of Pisa’s theory of the oscillating deferents of the inner planets (1h. 12th C.). Archive for History of Exact Sciences, 71(5):379–407, September 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://

link.springer.com/article/10.1007/s00407-017-0192-9.

Gaarding:1994:NHA

[GS94] Lars Gårding and Christian Skau. Niels Henrik Abel and solvable equations. Archive for History of Exact Sciences, 48(1):81– 103, March 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
48&issue=1&spage=81.

Gray:2008:SBG

[GS08] J. M. K. Gray and J. M. Steele. Studies on Babylonian goal-year astronomy I: a comparison between planetary data in goal-year texts, almanacs and normal star almanacs. Archive for History of Exact Sciences, 62(5):553–600, September 2008. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=62&issue=5&spage=553.

Gray:2009:SBG

[GS09] J. M. K. Gray and J. M. Steele. Studies on Babylonian goal-year astronomy II: the Babylonian calendar and goalyear methods of prediction. Archive for History of Exact Sciences, 63(6):611–633, November 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=6&spage=611.

Guan:2010:NIS

[Gua10] Yuzhen Guan. A new interpretation of Shen Kuo’s Ying Biao Yi. Archive for History of Exact Sciences, 64(6):707–719, November 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
64&issue=6&spage=707.

Guan:2015:ETJ

[Gua15] Yuzhen Guan. Eclipse theory in the Jing chu li: Part I. The adoption of lunar velocity. Archive for History of Exact Sciences, 69(1):103–123, January 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-014-0146-4.

Guerraggio:1984:MVP

[Gue84] Angelo Guerraggio. Le memorie di Volterra e Peano sul movimento dei poli. (Italian) [The memoirs of Volterra and Peano on the movement of the poles]. Archive for History of Exact Sciences, 31(2):97–126, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=31&issue=2&spage=97.

Guggenheimer:1977:JCT

[Gug77] H. Guggenheimer. The Jordan curve theorem and an unpublished manuscript by Max Dehn. Archive for History of Exact Sciences, 17(2):193–200, July 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=2&spage=193.

Guicciardini:2016:LTR

[Gui16] Niccolò Guicciardini. Lost in translation? Reading Newton on inverse-cube trajectories. Archive for History of Exact Sciences, 70(2):205–241, March 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0170-z.

Gulbekian:1987:OVS

[Gul87] Edward Gulbekian. The origin and value of the stadion unit used by Eratosthenes in the third century B.C. Archive for History of Exact Sciences, 37(4):359–363, December 1987. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=37&issue=4&spage=359.

Gimeno:2021:OCL

[GXB21] Gonzalo Gimeno, Mercedes Xipell, and Marià Baig. Operator calculus: the lost formulation of quantum mechanics. Archive for History of Exact Sciences, 75(3):283–322, May 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Garber:1979:EP

[GZ79] Daniel Garber and Sandy Zabell. On the emergence of probability. Archive for History of Exact Sciences, 21(1):33–53, March 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
21&issue=1&spage=33.

Gesztelyi:2012:HEO

[GZKB^{+}12] Rudolf Gesztelyi, Judit Zsuga, Adam Kemeny-Beke, Balazs Varga, Bela Juhasz, et al. The Hill equation and the origin of quantitative pharmacology. Archive for History of Exact Sciences, 66(4):427–438, July 2012. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=4&spage=427.

Hamilton:1998:BVT

[HA98] Norman T. Hamilton and Asger Aaboe. A Babylonian Venus text computed according to System A: ACT No. 1050. Archive for History of Exact Sciences, 53(3–4):215–221, November 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=3&spage=215.

Hahn:2002:PSA

[Hah02] Alexander J. Hahn. The pendulum swings again: a mathematical reassessment of Galileo’s experiments with inclined planes. Archive for History of Exact Sciences, 56(4):339– 361, June 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=4&spage=339.

Hall:1960:CD

[Hal60] A. Rupert Hall. Cartesian dynamics. Archive for History of Exact Sciences, 1(2):172–178, December 1960. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=2&spage=172.

Hald:1988:MSP

[Hal88] A. Hald. On de Moivre’s solutions of the problem of duration of play 1708–1718. Archive for History of Exact Sciences, 38(2):109–134, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=38&issue=2&spage=109.

Hald:1990:EBP

[Hal90] A. Hald. Evaluations of the beta probability integral by Bayes and Price. Archive for History of Exact Sciences, 41(2):139– 156, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=2&spage=139.

Hamburg:1976:TEC

[Ham76] Robin Rider Hamburg. The theory of equations in the 18th century: The work of Joseph Lagrange. Archive for History of Exact Sciences, 16(1):17–36, March 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=1&spage=17.

Hanle:1977:CAE

[Han77] Paul A. Hanle. The coming of age of Erwin Schrödinger: His quantum statistics of ideal gases. Archive for History of Exact Sciences, 17(2):165–192, July 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=2&spage=165.

Hannah:2011:CRP

[Han11] John Hannah. Conventions for recreational problems in Fibonacci’s Liber Abbaci. Archive for History of Exact Sciences, 65(2):155–180, March 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=????
&volume=65&issue=2&spage=155.

Harari:2003:CER

[Har03] Orna Harari. The concept of existence and the role of constructions in Euclid’s Elements. Archive for History of Exact Sciences, 57(1):1–23, January 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=57&issue=1&spage=1.

Hartshorne:2008:PHS

[Har08] Robin Hartshorne. Publication history of von Staudt’s Geometrie der Lage. Archive for History of Exact Sciences, 62(3):297– 299, May 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=3&spage=297.

Hawkins:1971:OTG

[Haw71] Thomas Hawkins. The origins of the theory of group characters. Archive for History of Exact Sciences, 7(2):142–170, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
7&issue=2&spage=142.

Hawkins:1972:HNL

[Haw72] Thomas Hawkins. Hypercomplex numbers, Lie groups, and the creation of group representation theory. Archive for History of Exact Sciences, 8(4):243–287, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=8&issue=4&spage=243.

Hawkins:1974:NLF

[Haw74] Thomas Hawkins. New light on Frobenius’ creation of the theory of group characters. Archive for History of Exact Sciences, 12(3):217–243, September 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=12&issue=3&spage=217.

Hawkins:1977:WTM

[Haw77] Thomas Hawkins. Weierstrass and the theory of matrices. Archive for History of Exact Sciences, 17(2):119–163, July 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
17&issue=2&spage=119.

Hawkins:1982:WKS

[Haw82] Thomas Hawkins. Wilhelm Killing and the structure of Lie algebras. Archive for History of Exact Sciences, 26(2):127– 192, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
26&issue=2&spage=127.

Hawkins:1988:HPT

[Haw88] Thomas Hawkins. Hesses’s principle of transfer and the representation of Lie algebras. Archive for History of Exact Sciences, 39(1):41–73, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=1&spage=41.

Hawkins:1991:JBL

[Haw91] Thomas Hawkins. Jacobi and the birth of Lie’s Theory of Groups. Archive for History of Exact Sciences, 42(3):187–278, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
42&issue=3&spage=187.

Hawkins:2005:FCP

[Haw05] Thomas Hawkins. Frobenius, Cartan, and the Problem of Pfaff. Archive for History of Exact Sciences, 59(4):381–436, May 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=4&spage=381.

Hawkins:2008:CFO

[Haw08a] Thomas Hawkins. Continued fractions and the origins of the Perron–Frobenius theorem. Archive for History of Exact Sciences, 62(6):655–717, November 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=6&spage=655.

Hawkins:2008:FSA

[Haw08b] Thomas Hawkins. Frobenius and the symbolical algebra of matrices. Archive for History of Exact Sciences, 62(1):23–57, January 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=1&spage=23.

Hlawka:1986:ETG

[HB86] Edmund Hlawka and Christa Binder. Über die Entwicklung der Theorie der gleichverteilung in den Jahren 1909 bis 1916. (German) [On the development of the theory of uniform distributions in the years 1909 to 1916]. Archive for History of Exact Sciences, 36(3):197–249, September 1986. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=36&issue=3&spage=197.

Hertz:1995:HHL

[HD95] H. G. Hertz and Manuel G. Doncel. Heinrich Hertz’s laboratory notes of 1887. Archive for History of Exact Sciences, 49(3):197–270, September 1995. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=49&issue=3&spage=197.

Heilbron:1968:SPR

[Hei68] John L. Heilbron. The scattering of α and β particles and Rutherford’s atom. Archive for History of Exact Sciences, 4(4):247–307, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=4&issue=4&spage=247.

Heimann:1970:MMC

[Hei70] P. M. Heimann. Maxwell and the modes of consistent representation. Archive for History of Exact Sciences, 6(3):171–213, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=3&spage=171.

Heller:2022:FKP

[Hel22] Henning Heller. Felix Klein’s projective representations of the groups S_{6} and A_{7}. Archive for History of Exact Sciences, 76(5): 431–470, September 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00290-x.

Hentschel:1994:EFF

[Hen94] Klaus Hentschel. Erwin Finlay Freundlich and testing Einstein’s Theory of Relativity. Archive for History of Exact Sciences, 47(2):143–201, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=2&spage=143.

Hentschel:1997:UDP

[Hen97] Klaus Hentschel. An unwelcome discovery: The pole effect in the electric arc, a threat to early 20th century precision spectrometry. Archive for History of Exact Sciences, 51(3):199–271, September 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=3&spage=199.

Hentschel:2001:BFS

[Hen01] Klaus Hentschel. Das Brechungsgesetz in der Fassung von Snellius Rekonstruktion seines Entdeckungspfades und eine Übersetzung seines lateinischen Manuskriptes sowie ergänzender Dokumente. (German) [The Law of Refraction in the version of Snellius’ reconstruction of its discovery path and a translation of its Latin manuscript as well as supplementing documents]. Archive for History of Exact Sciences, 55(4):297–344, January 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=4&spage=297.

Hermelink:1963:TJ

[Her63] Heinrich Hermelink. Tabulae Jahen. (German). Archive for History of Exact Sciences, 2(2):108–112, August 1963. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=2&spage=
108.

Herrmann:1970:AIJ

[Her70] Dieter B. Herrmann. Das Astronomentreffen im Jahre 1798 auf dem Seeberg bei Gotha. (German) [The astronomers’ meeting in the year 1798 in Seeberg bei Gotha]. Archive for History of Exact Sciences, 6(4):326–344, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=4&spage=326.

Hessenbruch:2000:RER

[Hes00] Arne Hessenbruch. Rutherford’s 1901 experiment on radiation energy and his creation of a stable detector. Archive for History of Exact Sciences, 54(5):403–420, February 2000. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=54&issue=5&spage=403.

Hon:2005:HEM

[HG05a] Giora Hon and Bernard R. Goldstein. How Einstein made asymmetry disappear: Symmetry and Relativity in 1905. Archive for History of Exact Sciences, 59(5):437–544, July 2005. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=59&issue=5&spage=437.

Hon:2005:LRD

[HG05b] Giora Hon and Bernard R. Goldstein. Legendre’s revolution (1794): The definition of symmetry in solid geometry. Archive for History of Exact Sciences, 59(2):107–155, January 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=2&spage=107.

Hewitt:1979:GWP

[HH79] Edwin Hewitt and Robert E. Hewitt. The Gibbs–Wilbraham phenomenon: an episode in Fourier analysis. Archive for History of Exact Sciences, 21(2):129–160, June 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=2&spage=129.

Heideman:1985:GHF

[HJB85] Michael T. Heideman, Don H. Johnson, and C. Sidney Burrus. Gauss and the history of the Fast Fourier Transform. Archive for History of Exact Sciences, 34(3):265–277, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
34&issue=3&spage=265.

Hofmann:1960:ZMF

[Hof60] J. E. Hofmann. Über zahlentheoretische Methoden Fermats und Eulers, ihre Zusammenhänge und ihre Bedeutung. (German) [On the number-theoretical method of Fermat and Euler, its relationships and its meaning]. Archive for History of Exact Sciences, 1(2):122–159, December 1960. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=2&spage=122.

Hofmann:1965:FIA

[Hof65] Jos. E. Hofmann. Aus der Frühzeit der Infinitesimalmethoden: Auseinandersetzung um die algebraische Quadratur algebraischer Kurven in der zweiten Hälfte des 17. Jahrhunderts. (German) [From the early period of infinitesimal methods: Arguments about the algebraic quadrature of algebraic curves in the second half of the 17^{th} century]. Archive for History of Exact Sciences, 2(4):271–343, April 1965. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=4&spage=271.

Hofmann:1966:KCH

[Hof66] Jos. E. Hofmann. Über die Kreismessung von Chr. Huygens, ihre Vorgeschichte, ihren Inhalt, ihre Bedeutung und ihr Nachwirken. (German) [On the circle measurement of Chr. Huygens, their prehistory, their contents, their meaning and their having a lasting effect]. Archive for History of Exact Sciences, 3(2):102–136, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=3&issue=2&spage=102.

Hogendijk:1984:GAC

[Hog84] Jan P. Hogendijk. Greek and Arabic constructions of the regular heptagon. Archive for History of Exact Sciences, 30 (3–4):197–330, September 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=30&issue=3&spage=197.

Hogendijk:1986:ATL

[Hog86] Jan P. Hogendijk. Arabic traces of lost works of Apollonius. Archive for History of Exact Sciences, 35(3):187–253, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=3&spage=187.

Hollings:2009:EDA

[Hol09] Christopher Hollings. The early development of the algebraic theory of semigroups. Archive for History of Exact Sciences, 63 (5):497–536, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Hollings:2014:ESG

[Hol14] Christopher Hollings. Embedding semigroups in groups: not as simple as it might seem. Archive for History of Exact Sciences, 68(5):641–692, September 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-014-0138-4.

Hollings:2017:NCP

[Hol17] Christopher D. Hollings. ‘Nobody could possibly misunderstand what a group is’: a study in early twentieth-century group axiomatics. Archive for History of Exact Sciences, 71(5):409–481, September 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://

link.springer.com/article/10.1007/s00407-017-0193-8;
 https://link.springer.com/content/pdf/10.1007/s00407-
017-0193-8.pdf.

Home:1967:FHT

[Hom67] Roderick W. Home. Francis Hauksbee’s theory of electricity. Archive for History of Exact Sciences, 4(3):203–217, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
4&issue=3&spage=203.

Hong:1994:CVC

[Hon94] Sungook Hong. Controversy over Voltaic Contact phenomena, 1862–1900. Archive for History of Exact Sciences, 47 (3):233–289, September 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=3&spage=233.

Hoyer:1973:RSE

[Hoy73] Ulrich Hoyer. Über die Rolle der Stabilitätsbetrachtungen in der Entwicklung der Bohrschen Atomtheorie. (German) [On the role of stability conditions in the development of Bohr’s atomic theory]. Archive for History of Exact Sciences, 10(3–5):177–206, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=10&issue=3&spage=177.

Hoyer:1974:ZCT

[Hoy74] Ulrich Hoyer. Über den Zusammenhang der Carnotschen Theorie mit der Thermodynamik. (German) [On the connection of Carnot theory with thermodynamics]. Archive for History of Exact Sciences, 13(4):359–375, December 1974. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=13&issue=4&spage=359.

Hoyer:1976:VCT

[Hoy76] Ulrich Hoyer. Das Verhältnis der Carnotschen Theorie zur klassischen Thermodynamik. (German) [The relationship of Carnot’s Theory to classical thermodynamics]. Archive for History of Exact Sciences, 15(2):149–197, June 1976. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=15&issue=2&spage=149.

Hoyer:1978:WMW

[Hoy78] Ulrich Hoyer. Über Waterstons mechanisches Wärmeäquivalent. (German) [On Waterston’s mechanical equivalent of heat]. Archive for History of Exact Sciences, 19(4):371–381, December 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=4&spage=371.

Hoyer:1980:BP

[Hoy80] Ulrich Hoyer. Von Boltzmann zu Planck. (German) [From Boltzmann to Planck]. Archive for History of Exact Sciences, 23(1): 47–86, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=1&spage=47.

Hoyrup:1988:JNC

[Høy88] Jens Høyrup. Jordanus de Nemore, 13^{th} century mathematical innovator: an essay on intellectual context, achievement, and failure. Archive for History of Exact Sciences, 38(4):307–363, September 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
38&issue=4&spage=307.

Hoyrup:2008:UHT

[Høy08] Jens Høyrup. The “Unknown Heritage”: trace of a forgotten locus of mathematical sophistication. Archive for History of Exact Sciences, 62(6):613–654, November 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=6&spage=613.

Huttig:2000:CST

[Hüt00] Manfred Hüttig. The conical sundial from Thyrrheion — reconstruction and error analysis of a displaced antique sundial. Archive for History of Exact Sciences, 55(2):163–176, December 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=2&spage=163.

Haubold:1986:JIT

[HY86] Hans Joachim Haubold and Eiichi Yasui. Jun Ishiwaras Text über Albert Einsteins Gastvortrag an der Universität zu Kyoto am 14. Dezember 1922. (German) [Jun Ishiwara’s text of Albert Einstein’s guest lecture at Kyoto University on 14 December 1922]. Archive for History of Exact Sciences, 36(3):271–279, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=36&issue=3&spage=271. Translated from

the Japanese by Hans Joachim Haubold and Eiichi Yasui.

Iversen:2019:BPI

[IJ19a] Paul Iversen and Alexander Jones. The Back Plate Inscription and eclipse scheme of the Antikythera Mechanism revisited. Archive for History of Exact Sciences, 73(5):469–511, September 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00407-019-00229-9.pdf. See correction [IJ19b].

Iversen:2019:CBP

[IJ19b] Paul Iversen and Alexander Jones. Correction to: The Back Plate Inscription and eclipse scheme of the Antikythera Mechanism revisited. Archive for History of Exact Sciences, 73(5): 513–516, September 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00407-019-00234-y.pdf. See

[IJ19a].

Illy:1989:ETL

[Ill89] József Illy. Einstein teaches Lorentz, Lorentz teaches Einstein: their collaboration in General Relativity, 1913–1920. Archive for History of Exact Sciences, 39(3):247–289, September 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
39&issue=3&spage=247.

Im:1996:ECF

[Im96] Gyeong Soon Im. Experimental constraints on formal quantum mechanics: The emergence of Born’s quantum theory of collision processes in Göttingen, 1924–1927. Archive for History of Exact Sciences, 50(1):73–101, March 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=1&spage=73.

Iohara:2021:MJA

[IM21] Kenji Iohara and Philippe Malbos. Maurice Janet’s algorithms on systems of linear partial differential equations. Archive for History of Exact Sciences, 75(1):43–81, January 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Iurato:2016:RVW

[IR16] Giuseppe Iurato and Giuseppe Ruta. On the role of virtual work in Levi-Civita’s parallel transport. Archive for History of Exact Sciences, 70(5):553–565, September 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-016-0177-0.

Jaeger:1998:ECP

[Jae98] Gregg Jaeger. The Ehrenfest classification of phase transitions: Introduction and evolution. Archive for History of Exact Sciences, 53(1):51–81, May 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=53&issue=1&spage=51.

Jaeck:2010:ESA

[Jaë10] Frédéric Jaëck. Éléments structurels en analyse fonctionnelle: trois notes de Fréchet sur les opérations linéaires. Archive for History of Exact Sciences, 64(4):461–483, 2010. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Jaeck:2020:WWG

[Jaë20] Frédéric Jaëck. What were the genuine Banach spaces in 1922? Reflection on axiomatisation and progression of the mathematical thought. Archive for History of Exact Sciences, 74(2): 109–129, March 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-019-00239-7.

Jahnke:1987:MPA

[Jah87] Hans Niels Jahnke. Motive und Probleme der Arithmetisierung der Mathematik in der ersten Hälfte des 19. Jahrhunderts — Cauchys Analysis in der Sicht des Mathematikers Martin Ohm. (German) [Motives and problems of the arithmetisation of mathematics in the first half of the 19^{th} century — Cauchy’s analysis in the view of the mathematician Martin Ohm]. Archive for History of Exact Sciences, 37(2):101– 182, June 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
37&issue=2&spage=101.

Jami:1988:HCD

[Jam88] Catherine Jami. Une histoire chinoise du “nombre π”. (French) [A Chinese history of the “number π”]. Archive for History of Exact Sciences, 38(1):39–50, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=38&issue=1&spage=39.

Jaouiche:1974:LDQ

[Jao74] Khalil Jaouiche. Le livre du qarast.ūn de Tābit ibn Qurra: Étude sur l’origine de la notion de travail et du calcul du moment statique d’une barre homogène. (French) [The book of qarast.ūn of Tābit ibn Qurra: Study of the origin of the notion of work and of the calculation of the static moment of a homogeneous bar]. Archive for History of Exact Sciences, 13(4):325–347, December 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

13&issue=4&spage=325. Étude sur l’origine de la notion de travail et du calcul du moment statique d’une barre homogène.

Jensen:1972:LTA

[Jen72] Claus Jensen. The lunar theories of al-baghdadi. Archive for History of Exact Sciences, 8(4):321–328, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=8&issue=4&spage=
321.

Johnson:1977:PDT

[Joh77] Dale M. Johnson. Prelude to dimension theory: The geometrical investigations of Bernard Bolzano. Archive for History of Exact Sciences, 17(3):261–295, September 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=3&spage=261.

Johnson:1979:PID

[Joh79] Dale M. Johnson. The problem of the invariance of dimension in the growth of modern topology, part I. Archive for History of Exact Sciences, 20(2):97–188, June 1979. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=20&issue=2&spage=97. See errata [Ano81].

Johnson:1981:PID

[Joh81] Dale M. Johnson. The problem of the invariance of dimension in the growth of modern topology, part II. Archive for History of Exact Sciences, 25(2–3):85–266, December 1981. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=25&issue=2&spage=85.

Johnsen:1984:UPH

[Joh84] Karsten Johnsen. Der “Untergruppensatz” von P. Hoyer — 25 Jahre vor O. Schreier. (German) [The subgroup theorem of P. Hoher — 25 years before O. Schreier]. Archive for History of Exact Sciences, 29(4):377–383, December 1984. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=29&issue=4&spage=377.

Jones:1983:DTD

[Jon83] Alexander Jones. The development and transmission of 248day schemes for lunar motion in ancient astronomy. Archive for History of Exact Sciences, 29(1):1–36, March 1983. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=29&issue=1&spage=1.

Jones:1999:LSO

[Jon99] Alexander Jones. A likely source of an observation report in Ptolemy’s Almagest. Archive for History of Exact Sciences, 54(3):255–258, August 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=3&spage=255.

Jones:2004:SBO

[Jon04] Alexander Jones. A study of Babylonian observations of planets near normal stars. Archive for History of Exact Sciences, 58(6):475–536, September 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=6&spage=475.

Jones:2010:OJP

[Jon10] Alexander Jones. Obituary: John P. Britton (1939–2010). Archive for History of Exact Sciences, 64(6):613–615, November 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
64&issue=6&spage=613.

Ji:2013:HDT

[JP13] Lizhen Ji and Athanase Papadopoulos. Historical development of Teichmüller theory. Archive for History of Exact Sciences, 67 (2):119–147, March 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-012-0104-y.

Jongmans:1994:PNP

[JS94] François Jongmans and Eugene Seneta. A probabilistic “New Principle” of the 19^{th} century. Archive for History of Exact Sciences, 47(1):93–102, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=1&spage=93.

Ji:2020:PSM

[JW20] Lizhen Ji and Chang Wang. Poincaré’s stated motivations for topology. Archive for History of Exact Sciences, 74(4):381– 400, July 2020. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL https://link.springer.com/
article/10.1007/s00407-020-00247-y.

Ji:2022:PWL

[JW22] Lizhen Ji and Chang Wang. Poincaré’s works leading to the Poincaré Conjecture. Archive for History of Exact Sciences, 76 (3):223–260, May 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00283-2.

Kainzinger:2011:MSS

[Kai11] Albert Kainzinger. The mathematics in the structures of Stonehenge. Archive for History of Exact Sciences, 65(1):67–97, January 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=1&spage=67.

Kalmar:1977:THR

[Kal77] Martin Kalmar. Thomas Hariot’s De Reflexione Corporum Rotundorum: an early solution to the problem of impact. Archive for History of Exact Sciences, 16(3):201–230, September 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
16&issue=3&spage=201.

Katz:1985:DFC

[Kat85] Victor J. Katz. Differential forms — Cartan to de Rham. Archive for History of Exact Sciences, 33(4):321–336, December 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
33&issue=4&spage=321.

Katzir:2003:DPE

[Kat03] Shaul Katzir. The discovery of the piezoelectric effect. Archive for History of Exact Sciences, 57(1):61–91, January 2003. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=57&issue=1&spage=61.

Katzir:2008:UFS

[Kat08] Shaul Katzir. From ultrasonic to frequency standards: Walter Cady’s discovery of the sharp resonance of crystals. Archive for History of Exact Sciences, 62(5):469–487, September 2008. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=62&issue=5&spage=469.

Kaunzner:1985:FLB

[Kau85] Wolfgang Kaunzner. Über eine frühe lateinische Bearbeitung der Algebra al-Khwārizmīs in MS Lyell 52 der Bodleian Library Oxford. (German) [On early Latin treatment of algebra: Al-Khwārizmī in MS Lyell 52 the Bodleian Library Oxford]. Archive for History of Exact Sciences, 32(1):1–16, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=1&spage=1.

Kay:2023:FKS

[Kay23] L. D. Kay. Felix Klein, Sophus Lie, contact transformations, and connexes. Archive for History of Exact Sciences, 77(4):373– 391, July 2023. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL https://link.springer.com/
article/10.1007/s00407-023-00305-1.

Kragh:2000:CBM

[KC00] Helge Kragh and Bruno Carazza. Classical behavior of macroscopic bodies from quantum principles: Early discussions. Archive for History of Exact Sciences, 55(1):43–56, August 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=1&spage=43.

Kerszberg:1986:PWI

[Ker86] Pierre Kerszberg. Le principe de Weyl et l’invention d’une cosmologie non-statique. (French) [Weyl’s Principle and the invention of a nonstatic cosmology]. Archive for History of Exact Sciences, 35(1):1–89, March 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=35&issue=1&spage=1.

Keyser:1992:NLH

[Key92] Paul Keyser. A new look at Heron’s “steam Engine”. Archive for History of Exact Sciences, 44(2):107–124, June 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=44&issue=2&spage=107.

Keyser:2022:ASD

[Key22] Paul T. Keyser. The Archimedean ‘sambuke’ of Damis in Biton. Archive for History of Exact Sciences, 76(2):153–172, March 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-021-00281-4.

Kiernan:1971:DGT

[Kie71] B. Melvin Kiernan. The development of Galois theory from Lagrange to Artin. Archive for History of Exact Sciences, 8(1–2):40–154, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=8&issue=1&spage=40.

King:1973:IYV

[Kin73] David A. King. Ibn Yunus’ very useful tables for reckoning time by the sun. Archive for History of Exact Sciences, 10(3–5):342–394, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=10&issue=3&spage=342.

Kjeldsen:2001:JNC

[Kje01] Tinne Hoff Kjeldsen. John von Neumann’s conception of the Minimax Theorem: a journey through different mathematical contexts. Archive for History of Exact Sciences, 56(1):39–68, November 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=1&spage=39.

Kjeldsen:2002:DMG

[Kje02] Tinne Hoff Kjeldsen. Different motivations and goals in the historical development of the theory of systems of linear inequalities. Archive for History of Exact Sciences, 56(6):469–538, November 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=6&spage=469.

Kjeldsen:2008:MTG

[Kje08] Tinne Hoff Kjeldsen. From measuring tool to geometrical object: Minkowski’s development of the concept of convex bodies. Archive for History of Exact Sciences, 62(1):59–89, January 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=1&spage=59.

King:1984:CKR

[KL84] M. Christine King and Keith J. Laidler. Chemical kinetics and the radiation hypothesis. Archive for History of Exact Sciences, 30(1):45–86, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=30&issue=1&spage=45.

Klein:1961:MPB

[Kle61] Martin J. Klein. Max Planck and the beginnings of the quantum theory. Archive for History of Exact Sciences, 1(5):459–479, October 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
1&issue=5&spage=459.

Knobloch:1974:USL

[Kno74] Eberhard Knobloch. Unbekannte Studien von Leibniz zur Eliminations- und Explikationstheorie. (German) [Unknown studies from Leibniz on the elimination and explanation theory]. Archive for History of Exact Sciences, 12(2):142–173, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
12&issue=2&spage=142.

Knorr:1976:AMC

[Kno76] Wilbur R. Knorr. Archimedes and the measurement of the circle: a new interpretation. Archive for History of Exact Sciences, 15(2):115–140, June 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=15&issue=2&spage=115.

Knorr:1978:AEP

[Kno78] Wilbur R. Knorr. Archimedes and the Elements: Proposal for a revised chronological ordering of the Archimedean corpus. Archive for History of Exact Sciences, 19(3):211–290, September 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=3&spage=211.

Knorr:1981:AIS

[Kno81] Wilbur Knorr. Aristotle and incommensurability: Some further reflections. Archive for History of Exact Sciences, 24(1): 1–9, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

24&issue=1&spage=1. See correction [Kno82].

Knorr:1982:CMA

[Kno82] Wilbur Knorr. A correction to my article “Aristotle and Incommensurability”. Archive for History of Exact Sciences, 27(4):391–392, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=4&spage=391. See [Kno81].

Knorr:1986:ADC

[Kno86] Wilbur R. Knorr. Archimedes’ dimension of the circle: a view of the genesis of the extant text. Archive for History of Exact Sciences, 35(4):281–324, December 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=35&issue=4&spage=281.

Knorr:1992:WCD

[Kno92] Wilbur R. Knorr. When circles don’t look like circles: an optical theorem in Euclid and Pappus. Archive for History of Exact Sciences, 44(4):287–329, December 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=4&spage=287.

Knobloch:1999:GLD

[Kno99] Eberhard Knobloch. Galileo and Leibniz: Different approaches to infinity. Archive for History of Exact Sciences, 54(2):87– 99, June 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=2&spage=87.

Knudsen:1976:FEP

[Knu76] Ole Knudsen. The Faraday effect and physical theory, 1845– 1873. Archive for History of Exact Sciences, 15(3):235–281, September 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
15&issue=3&spage=235.

Koppelman:1971:COR

[Kop71] Elaine Koppelman. The calculus of operations and the rise of abstract algebra. Archive for History of Exact Sciences, 8(3):155– 242, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
8&issue=3&spage=155.

Kox:1988:HAL

[Kox88] A. J. Kox. Hendrik Antoon Lorentz, the ether, and the General Theory of Relativity. Archive for History of Exact Sciences, 38(1):67–78, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=38&issue=1&spage=67.

Kox:2013:HAL

[Kox13] A. J. Kox. Hendrik Antoon Lorentz’s struggle with quantum theory. Archive for History of Exact Sciences, 67(2):149–170, March 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://link.springer.com/article/10.
1007/s00407-012-0107-8;   http://link.springer.com/content/
pdf/10.1007/s00407-012-0107-8.pdf.

Kragh:1981:GDR

[Kra81] Helge Kragh. The genesis of Dirac’s Relativistic Theory of Electrons. Archive for History of Exact Sciences, 24(1):31– 67, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
24&issue=1&spage=31.

Kragh:1992:RCW

[Kra92] Helge Kragh. Relativistic collisions: The work of Christian Møller in the early 1930s. Archive for History of Exact Sciences, 43(4):299–328, December 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=43&issue=4&spage=299.

Kragh:1997:ORS

[Kra97] Helge Kragh. The origin of radioactivity: From solvable problem to unsolved non-problem. Archive for History of Exact Sciences, 50(3–4):331–358, September 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=3&spage=331.

Kragh:2003:MNP

[Kra03] Helge Kragh. Magic number: a partial history of the finestructure constant. Archive for History of Exact Sciences, 57(5): 395–431, July 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=5&spage=395.

Kragh:2010:AMF

[Kra10] Helge Kragh. An anthropic myth: Fred Hoyle’s carbon-12 resonance level. Archive for History of Exact Sciences, 64(6):721–751, November 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=64&issue=6&spage=721.

Kragh:2012:PDE

[Kra12] Helge Kragh. Preludes to dark energy: zero-point energy and vacuum speculations. Archive for History of Exact Sciences, 66(3):199–240, May 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=3&spage=199.

Kragh:2023:THE

[Kra23] Helge Kragh. A terminological history of early elementary particle physics. Archive for History of Exact Sciences, 77(1): 73–120, January 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00299-2.

Krieger:1968:SBA

[Kri68] H. Krieger. Über Stirlings Bestimmung von ∑_{n=0}^{∞}(−1)^{n}/(2n+1) und ∑_{n=0}^{∞} 1/n^{2}. (German) [On Stirling’s determination of ∑_{n=0}^{∞}(−1)^{n}/(2n + 1) and ∑_{n=0}^{∞} 1/n^{2}]. Archive for History of Exact Sciences, 5(1):37–46, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=1&spage=37.

Koch:1990:NIK

[KSFW90] Ulla Koch, Joachim Schaper, Susanne Fischer, and Michael Wegelin. Eine neue Interpretation der Kudurru–Symbole. Archive for History of Exact Sciences, 41(2):93–114, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=41&issue=2&spage=
93.

Kubli:1970:LBE

[Kub70] Fritz Kubli. Louis de Broglie und die Entdeckung der Materiewellen. (German) [Louis de Broglie and the discovery of matter waves]. Archive for History of Exact Sciences, 7(1):26– 68, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
7&issue=1&spage=26.

Kushner:1989:CSS

[Kus89] David Kushner. The controversy surrounding the secular acceleration of the Moon’s mean motion. Archive for History of Exact Sciences, 39(4):291–316, December 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=4&spage=291.

Kyriazis:2021:PAP

[Kyr21] G. A. Kyriazis. On Peirce’s 1878 article ‘The probability of induction’: a conceptualistic appraisal. Archive for History of Exact Sciences, 75(1):1–20, January 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Lam:1984:LYH

[LA84] Lay-Yong Lam and Tian-Se Ang. Li Ye and his Yi Gu Yan Duan (old mathematics in expanded sections). Archive for History of Exact Sciences, 29(3):237–266, September 1984. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=29&issue=3&spage=237.

Lacki:2000:EAQ

[Lac00] Jan Lacki. The early axiomatizations of quantum mechanics: Jordan, von Neumann and the continuation of Hilbert’s program. Archive for History of Exact Sciences, 54(4):279–318, January 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=4&spage=279.

Laidler:1985:CKO

[Lai85] Keith J. Laidler. Chemical kinetics and the origins of physical chemistry. Archive for History of Exact Sciences, 32(1):43– 75, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=1&spage=43.

Laidler:1986:DTC

[Lai86] Keith J. Laidler. The development of theories of catalysis. Archive for History of Exact Sciences, 35(4):345–374, December 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=4&spage=345.

Laidler:1988:CKO

[Lai88] Keith J. Laidler. Chemical kinetics and the Oxford College Laboratories. Archive for History of Exact Sciences, 38(3):197–283, September 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
38&issue=3&spage=197.

Lam:1979:CSC

[Lam79] Lay Yong Lam. Chu Shih-chieh’s Suan-hsüeh ch’i-meng (introduction to mathematical studies). Archive for History of Exact Sciences, 21(1):1–31, March 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=1&spage=1.

Lam:1986:COO

[Lam86] Lay Yong Lam. The conceptual origins of our numeral system and the symbolic form of algebra. Archive for History of Exact Sciences, 36(3):183–195, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=36&issue=3&spage=183.

Lam:1987:LES

[Lam87] Lay Yong Lam. Linkages: Exploring the similarities between the Chinese rod numeral system and our numeral system. Archive for History of Exact Sciences, 37(4):365–392, December 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
37&issue=4&spage=365.

Lam:1988:CGR

[Lam88] Lay Yong Lam. A Chinese genesis: Rewriting the history of our numeral system. Archive for History of Exact Sciences, 38(2): 101–108, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

38&issue=2&spage=101. See correction [Lam89].

Lam:1989:CAC

[Lam89] Lay Yong Lam. Correction for the article: “A Chinese genesis: rewriting the history of our numeral system”. Archive for History of Exact Sciences, 39(4):379, 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). See [Lam88].

Lam:1994:JZS

[Lam94] Lay Yong Lam. Jiu zhang suanshu (nine chapters on the mathematical art): an overview. Archive for History of Exact Sciences, 47(1):1–51, June 1994. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=1&spage=1.

Lam:1997:ZQS

[Lam97] Lay Yong Lam. Zhang Qiujian Suanjing (the mathematical classic of Zhang Qiujian): an overview. Archive for History of Exact Sciences, 50(3–4):201–240, September 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=3&spage=201.

Lamande:2013:QCT

[Lam13] Pierre Lamandé. Quelques conceptions de la théorie des proportions dans des traités de la seconde moitié du dix septième siècle. (French) [Some conceptions of the theory of proportions in the treatises of the second half of the 17th Century]. Archive for History of Exact Sciences, 67(6):595–636, November 2013. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0120-6.

Laugwitz:1965:BBG

[Lau65] Detlef Laugwitz. Bemerkungen zu Bolzanos Größenlehre. (German) [Remarks on Bolzano’s size teachings [????]]. Archive for History of Exact Sciences, 2(5):398–409, August 1965. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=5&spage=
398.

Laugwitz:1989:DVI

[Lau89] Detlef Laugwitz. Definite values of infinite sums: Aspects of the foundations of infinitesimal analysis around 1820. Archive for History of Exact Sciences, 39(3):195–245, September 1989. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=39&issue=3&spage=195.

Landry-Deron:2001:MEC

[LD01] Isabelle Landry-Deron. Les Mathématiciens envoyés en Chine par Louis XIV en 1685. (French) [The mathematicians sent to China by Louis XIV in 1685]. Archive for History of Exact Sciences, 55(5):423–463, April 2001. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=5&spage=423.

Le:2017:ACG

[Lê17] François Lê. Alfred Clebsch’s “geometrical clothing” of the theory of the quintic equation. Archive for History of Exact Sciences, 71(1):39–70, January 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-016-0180-5.

Lervig:1972:SCT

[Ler72] Philip Lervig. On the structure of Carnot’s theory of heat. Archive for History of Exact Sciences, 9(3):222–239, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
9&issue=3&spage=222.

Lindberg:1968:TPI

[Lin68] David C. Lindberg. The theory of Pinhole images from antiquity to the thirteenth century. Archive for History of Exact Sciences, 5(2):154–176, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=2&spage=154.

Lindberg:1970:RRB

[Lin70a] David C. Lindberg. A reconsideration of Roger Bacon’s theory of pinhole images. Archive for History of Exact Sciences, 6(3):214– 223, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=3&spage=214.

Lindberg:1970:TPI

[Lin70b] David C. Lindberg. The theory of Pinhole images in the fourteenth century. Archive for History of Exact Sciences, 6(4):299– 325, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=4&spage=299.

Little:2022:ECS

[Lit22] John B. Little. The eclectic content and sources of Clavius’s Geometria Practica. Archive for History of Exact Sciences, 76 (4):391–424, July 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00288-5.

Liu:2005:RRS

[Liu05] Ciyuan Liu. The regular records of solar eclipse in Ancient China and a computer readable table. Archive for History of Exact Sciences, 59(2):157–168, January 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=2&spage=157.

Li:1998:AEM

[LM98] Wenlin Li and Jean-Claude Martzloff. Aperçu sur les échanges mathématiques entre la Chine et la France (1880–1949). (French) [Views of the mathematical exchanges between China and France (1880–1949)]. Archive for History of Exact Sciences, 53(3–4):181–200, November 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=53&issue=3&spage=181.

Lohne:1961:NSS

[Loh61] Johs. Lohne. Newton’s “proof” of the sine law and his mathematical principles of colors. Archive for History of Exact Sciences, 1(4):389–405, July 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=4&spage=389.

Lohne:1966:DRT

[Loh66] J. A. Lohne. Dokumente zur Revalidierung von Thomas Harriot als Algebraiker. (German) [Documents on the revalidation of Thomas Harriot as an algebraicist]. Archive for History of Exact Sciences, 3(3):185–205, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=3&issue=3&spage=185.

Lohne:1968:EEW

[Loh68] J. A. Lohne. Der Eigenartige Einfluß Witelos auf die Entwicklung der Dioptrik. (German) [The strange influence of Witelos on the development of the Dioptrik]. Archive for History of Exact Sciences, 4(5):414–426, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=4&issue=5&spage=414.

Lohne:1979:ETH

[Loh79] J. A. Lohne. Essays on Thomas Harriot. Archive for History of Exact Sciences, 20(3–4):189–312, September 1979. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=20&issue=3&spage=189. I.

Billiard balls and laws of collision, II. Ballistic parabolas, III. A survey of Harriot’s scientific writings.

Lorch:1995:PMT

[Lor95] R. P. Lorch. Ptolemy and Maslama on the transformation of circles into circles in stereographic projection. Archive for History of Exact Sciences, 49(3):271–284, September 1995. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=49&issue=3&spage=271.

Lorenat:2016:SAG

[Lor16] Jemma Lorenat. Synthetic and analytic geometries in the publications of Jakob Steiner and Julius Plücker (1827–1829). Archive for History of Exact Sciences, 70(4):413–462, July 2016. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
015-0174-8.

Loveland:2001:BCS

[Lov01] Jeff Loveland. Buffon, the certainty of sunrise, and the probabilistic reductio ad absurdum. Archive for History of Exact Sciences, 55(5):465–477, April 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=5&spage=465.

Lazaroff-Puck:2015:GLD

[LP15] Cameron Lazaroff-Puck. Gearing up for Lagrangian dynamics. Archive for History of Exact Sciences, 69(5):455–490, September 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-015-0157-9.

Lam:1984:RAT

[LS84] Lay Yong Lam and Kang Sheng Shen. Right-angled triangles in ancient China. Archive for History of Exact Sciences, 30(2):87– 112, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
30&issue=2&spage=87.

Lam:1986:MPS

[LS86] Lay-Yong Lam and Kangsheng Shen. Mathematical problems on surveying in ancient China. Archive for History of Exact Sciences, 36(1):1–20, March 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=36&issue=1&spage=1.

Lutzen:1979:HOC

[Lüt79] Jesper Lützen. Heaviside’s operational calculus and the attempts to rigorise it. Archive for History of Exact Sciences, 21(2):161– 200, June 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
21&issue=2&spage=161.

Lutzen:1984:JLW

[Lüt84a] Jesper Lützen. Joseph Liouville’s work on the figures of equilibrium of a rotating mass of fluid. Archive for History of Exact Sciences, 30(2):113–166, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=30&issue=2&spage=113.

Lutzen:1984:SLW

[Lüt84b] Jesper Lützen. Sturm and Liouville’s work on ordinary linear differential equations. The emergence of Sturm–Liouville theory. Archive for History of Exact Sciences, 29(4):309–376, December 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
29&issue=4&spage=309.

Lutzen:1994:NCT

[Lüt94] Jesper Lützen. Note by the communicator of two papers, by J. Barrow–Green and by K. G. Andersson, on Poincaré. Archive for History of Exact Sciences, 48(2):105–106, June 1994. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=48&issue=2&spage=105.

Lutzen:1995:IBM

[Lüt95] Jesper Lützen. Interactions between mechanics and differential geometry in the 19^{th} century. Archive for History of Exact Sciences, 49(1):1–72, March 1995. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=49&issue=1&spage=1.

Macuglia:2023:SEC

[Mac23] Daniele Macuglia. SHAKE and the exact constraint satisfaction of the dynamics of semi-rigid molecules in Cartesian coordinates, 1973–1977. Archive for History of Exact Sciences, 77 (4):345–371, July 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-023-00306-0.

Maeyama:1988:KMM

[Mae88] Y. Maeyama. The Keplerian and mean motions. A geometrical study. Archive for History of Exact Sciences, 38(4):365–383, September 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
38&issue=4&spage=365.

Maeyama:1990:KHV

[Mae90] Y. Maeyama. Kepler’s hypothesis vicaria. Archive for History of Exact Sciences, 41(1):53–92, March 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=41&issue=1&spage=53.

Maeyama:1998:DSO

[Mae98] Y. Maeyama. Determination of the Sun’s orbit: Hipparchus, Ptolemy, al-Battani, Copernicus, Tycho Brahe. Archive for History of Exact Sciences, 53(1):1–49, May 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=53&issue=1&spage=1.

Maeyama:2006:ELV

[Mae06] Y. Maeyama. On the extreme lunar velocities. Archive for History of Exact Sciences, 60(3):269–283, May 2006. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=60&issue=3&spage=269.

Magnus:2022:GLS

[Mag22] Jan R. Magnus. Gauss on least-squares and maximum-likelihood estimation. Archive for History of Exact Sciences, 76(4):425– 430, July 2022. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL https://link.springer.com/
article/10.1007/s00407-022-00291-w.

Mahoney:1968:ALG

[Mah68] Michael S. Mahoney. Another look at Greek geometrical analysis. Archive for History of Exact Sciences, 5(3–4):318–348, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
5&issue=3&spage=318.

Maieru:1982:QPE

[Mai82] Luigi Maierù. Il quinto postulato Euclideo da C. Clavio [1589] a G. Saccheri [1733]. Archive for History of Exact Sciences, 27(4):297–334, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=4&spage=297.

Maieru:1989:NPP

[Mai89] L. Maierú. Le “nonnullae propositiones de parabola” di Marino Ghetaldi. (Italian) [The nonnull propositions of the parabola of Marino Ghetaldi]. Archive for History of Exact Sciences, 40(3):207–245, September 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=40&issue=3&spage=207.

Maieru:1990:SCC

[Mai90] Luigi Maierù. ”… in Christophorum Clavium de Contactu Linearum Apologia”: Considerazioni attorno alla polemica fra Peletier e Clavio circa l’angolo di contatto (1579–1589). (Italian) [“… in Christophorum Clavium de Contactu Linearum Apologia”: Considerations around the controversy between Peletier and Clavius about the contact angle (1579– 1589)]. Archive for History of Exact Sciences, 41(2):115– 137, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=2&spage=115.

Maieru:1994:SDM

[Mai94a] Luigi Maierù. La “Filosofia delle Matematiche” di P. Galluppi. (Italian) [The ‘Philosophy of Mathematics’ of P. Galluppi]. Archive for History of Exact Sciences, 47(3):203–231, September 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
47&issue=3&spage=203.

Maieru:1994:RFM

[Mai94b] Luigi Maierù. Rapporto fra “metodo” e “contenuti matematici” in P. Gesualdo Melacrinò da Reggio Calabria. (Italian) [Relationship between ‘method’ and ‘contained mathematics’ in Fater Gesualdo Melacrinò of Reggio Calabria]. Archive for History of Exact Sciences, 48(1):43–79, March 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=48&issue=1&spage=43.

Malley:1991:HCC

[Mal91a] Marjorie Malley. A heated controversy on cold light. Archive for History of Exact Sciences, 42(2):173–186, June 1991. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=42&issue=2&spage=173.

Maltese:1991:RRT

[Mal91b] Giulio Maltese. The rejection of the Ricci tensor in Einstein’s first tensorial Theory of Gravitation. Archive for History of Exact Sciences, 41(4):363–381, December 1991. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=41&issue=4&spage=363.

Malet:1993:JGT

[Mal93] Antoni Malet. James Gregorie on tangents and the “Taylor” rule for series expansions. Archive for History of Exact Sciences, 46(2):97–137, June 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=46&issue=2&spage=97.

Maltese:2000:RML

[Mal00] Giulio Maltese. On the relativity of motion in Leonhard Euler’s science. Archive for History of Exact Sciences, 54(4):319–348, January 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=4&spage=319.

Malpangotto:2010:GCG

[Mal10] Michela Malpangotto. Graphical choices and geometrical thought in the transmission of Theodosius’ spherics from antiquity to the Renaissance. Archive for History of Exact Sciences, 64(1):75–112, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Malpangotto:2016:OMC

[Mal16] Michela Malpangotto. The original motivation for Copernicus’s research: Albert of Brudzewo’s Commentariolum super Theoricas novas Georgii Purbachii. Archive for History of Exact Sciences, 70(4):361–411, July 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-015-0171-y.

Manning:1975:EWA

[Man75] Kenneth R. Manning. The emergence of the Weierstrassian approach to complex analysis. Archive for History of Exact Sciences, 14(4):297–383, December 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=14&issue=4&spage=297.

Mancha:1989:EBT

[Man89] J. L. Mancha. Egidius of Baisiu’s theory of pinhole images. Archive for History of Exact Sciences, 40(1):1–35, March 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
40&issue=1&spage=1.

Mancha:1992:AUP

[Man92] J. L. Mancha. Astronomical use of pinhole images in William of Saint–Cloud’s Almanach Planetarum (1292). Archive for History of Exact Sciences, 43(4):275–298, December 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=43&issue=4&spage=275.

Mancha:1998:HRA

[Man98a] J. L. Mancha. Heuristic reasoning: Approximation procedures in Levi ben Gerson’s astronomy. Archive for History of Exact Sciences, 52(1):13–50, January 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=52&issue=1&spage=13.

Mancha:1998:IAK

[Man98b] J. L. Mancha. On Ibn Al-Kammad’s table for trepidation. Archive for History of Exact Sciences, 52(1):1–11, January 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
52&issue=1&spage=1.

Mancha:2004:ABR

[Man04] J. L. Mancha. Al-Bit.ūjī’s theory of the motions of the fixed stars. Archive for History of Exact Sciences, 58(2):143–182, January 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
58&issue=2&spage=143.

Maracchia:1980:AI

[Mar80] Silvio Maracchia. Aristotele e l’incommensurabilità. (Italian) [Aristotle and incommensurability]. Archive for History of Exact Sciences, 21(3):201–228, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=3&spage=201.

Martin:1993:ELI

[Mar93] M. Carmen Hernández Martín. Esbozo de una logica interna en la obra arquimediana. (Spanish) [Outline of internal logic in the works of Archimedes]. Archive for History of Exact Sciences, 46(2):139–151, June 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=46&issue=2&spage=139.

Martinez:2006:RCT

[Mar06] Alberto A. Martínez. Replication of Coulomb’s torsion balance experiment. Archive for History of Exact Sciences, 60 (6):517–563, November 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=60&issue=6&spage=517.

Marx:2012:ICP

[Mar12] Christian Marx. Investigations of the coordinates in Ptolemy’s Geographike Hyphegesis Book 8. Archive for History of Exact Sciences, 66(5):531–555, September 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=5&spage=531.

Marx:2021:MPS

[Mar21] Christian Marx. On the making of Ptolemy’s star catalog. Archive for History of Exact Sciences, 75(1):21–42, January 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Masia:2015:DHA

[Mas15] Ramon Masià. On dating Hero of Alexandria. Archive for History of Exact Sciences, 69(3):231–255, May 2015. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
014-0148-2.

Masia:2016:NRA

[Mas16] Ramon Masià. A new reading of Archytas’ doubling of the cube and its implications. Archive for History of Exact Sciences, 70 (2):175–204, March 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0165-9.

Mathews:1978:WRH

[Mat78] Jerold Mathews. William Rowan Hamilton’s paper of 1837 on the arithmetization of analysis. Archive for History of Exact Sciences, 19(2):177–200, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=19&issue=2&spage=177.

Mathews:1985:NOT

[Mat85] Jerold Mathews. A neolithic oral tradition for the van der Waerden/Seidenberg origin of mathematics. Archive for History of Exact Sciences, 34(3):193–220, September 1985. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=34&issue=3&spage=193.

McCarthy:2003:AKS

[MB03] Daniel P. McCarthy and John G. Byrne. Al-Khwārizmī’s sine tables and a Western table with the Hindu norm of R = 150. Archive for History of Exact Sciences, 57(3):243– 266, April 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=3&spage=243.

McCormmach:1966:ATJ

[McC66] Russell McCormmach. The atomic theory of John William Nicholson. Archive for History of Exact Sciences, 3(2):160–184, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
3&issue=2&spage=160.

McCarthy:1996:LPT

[McC96] Daniel McCarthy. The lunar and Paschal tables of De ratione paschali attributed to Anatolius of Laodicea. Archive for History of Exact Sciences, 49(4):285–320, December 1996. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=49&issue=4&spage=285.

McGuire:1966:BVN

[McG66] J. E. McGuire. Body and void and Newton’s De Mundi systemate: Some new sources. Archive for History of Exact Sciences, 3(3):206–248, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=3&issue=3&spage=206.

McHugh:1971:HSO

[McH71] James A. M. McHugh. An historical survey of ordinary linear differential equations with a large parameter and turning points. Archive for History of Exact Sciences, 7(4):277–324, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
7&issue=4&spage=277.

McLarty:2011:ENF

[McL11] Colin McLarty. Emmy Noether’s first great mathematics and the culmination of first-phase logicism, formalism, and intuitionism. Archive for History of Exact Sciences, 65(1):99–117, January 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=1&spage=99.

Meli:2006:ICF

[Mel06] Domenico Bertoloni Meli. Inherent and centrifugal forces in Newton. Archive for History of Exact Sciences, 60(3):319– 335, May 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
60&issue=3&spage=319.

Mendell:2007:TTT

[Men07] Henry Mendell. Two traces of two-step Eudoxan proportion theory in Aristotle: a tale of definitions in Aristotle, with a moral. Archive for History of Exact Sciences, 61(1):3–37, January 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=1&spage=3.

Mendell:2023:HTC

[Men23] Henry Mendell. Hero and the tradition of the circle segment. Archive for History of Exact Sciences, 77(5):451–499, September 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00308-y.

Merker:2002:AAC

[Mer02] Anne Merker. Aristote et l’arc-en-ciel: Enjeux philosophiques et étude scientifique. (French) [Aristotle and the rainbow: Philosophical issues and scientific studies]. Archive for History of Exact Sciences, 56(3):183–238, May 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=56&issue=3&spage=183.

Mercier:2007:SSM

[Mer07] Raymond Mercier. The standard scheme of the Moon and its mean quantities. Archive for History of Exact Sciences, 61(3): 255–272, May 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=3&spage=255.

Meschkowski:1965:BGC

[Mes65] Herbert Meschkowski. Aus den Briefbüchern Georg Cantors. (German) [From the diaries of George Cantor]. Archive for History of Exact Sciences, 2(6):503–519, December 1965. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=6&spage=
503.

Middleton:1969:PDB

[Mid69] W. E. Knowles Middleton. Paolo del Buono on the elasticity of air. Archive for History of Exact Sciences, 6(1):1–28, January 1969. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=1&spage=1.

Middleton:1973:GAB

[Mid73] W. E. Knowles Middleton. Giovanni Alfonso Borelli and the invention of the Heliostat. Archive for History of Exact Sciences, 10(3–5):329–341, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=10&issue=3&spage=329.

Miller:1973:SHP

[Mil73] Arthur I. Miller. A study of Henri Poincaré’s “Sur la Dynamique de l’Électron”. Archive for History of Exact Sciences, 10(3–5):207–328, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=10&issue=3&spage=207.

Mills:1983:CBC

[Mil83] Stella Mills. The controversy between Colin MacLaurin and George Campbell over complex roots, 1728–1729. Archive for History of Exact Sciences, 28(2):149–164, June 1983. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=28&issue=2&spage=149.

Mills:1985:IDL

[Mil85] Stella Mills. The independent derivations by Leonhard Euler and Colin MacLaurin of the Euler–MacLaurin summation formula. Archive for History of Exact Sciences, 33(1–3):1– 13, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
33&issue=1&spage=1.

Miller:2017:PRP

[Mil17] David Marshall Miller. The Parallelogram Rule from pseudo-Aristotle to Newton. Archive for History of Exact Sciences, 71 (2):157–191, March 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-016-0183-2.

Mittelstrass:1965:EMW

[Mit65] Jürgen Mittelstrass. Die Entdeckung der Möglichkeit von Wissenschaft. (German) [The development of probability in science]. Archive for History of Exact Sciences, 2(5):410–435, August 1965. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
2&issue=5&spage=410.

Mitchell:2019:CEC

[Mit19a] Daniel Jon Mitchell. Correction to: ”The etherealization of common sense?” Arithmetical and algebraic modes of intelligibility in late Victorian mathematics of measurement. Archive for History of Exact Sciences, 73(2):181, March 2019. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-018-0220-4.pdf. See [Mit19b].

Mitchell:2019:ECS

[Mit19b] Daniel Jon Mitchell. “The etherealization of common sense?” Arithmetical and algebraic modes of intelligibility in late Victorian mathematics of measurement. Archive for History of Exact Sciences, 73(2):125–180, March 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). See correction [Mit19a].

Moesgaard:1974:TIQ

[Moe74] Kristian Peder Moesgaard. Thabit Ibn Qurra between Ptolemy and Copernicus: an analysis of Thabit’s solar theory. Archive for History of Exact Sciences, 12(3):199–216, September 1974. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=12&issue=3&spage=199.

Molland:1978:EBG

[Mol78] A. G. Molland. An examination of Bradwardine’s geometry. Archive for History of Exact Sciences, 19(2):113–175, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=2&spage=113.

Monna:1972:CFC

[Mon72] A. F. Monna. The concept of function in the 19^{th} and 20^{th} centuries, in particular with regard to the discussions between Baire, Borel and Lebesgue. Archive for History of Exact Sciences, 9(1):57–84, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=9&issue=1&spage=57.

Mozaffari:2014:MYA

[Moz14] S. Mohammad Mozaffari. Muh.yi al-Din al-Maghribi’s lunar measurements at the Maragha observatory. Archive for History of Exact Sciences, 68(1):67–120, January 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-013-0130-4.

Mozaffari:2016:FSM

[Moz16a] S. Mohammad Mozaffari. A forgotten solar model. Archive for History of Exact Sciences, 70(3):267–291, May 2016. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
015-0167-7.

Mozaffari:2016:PLM

[Moz16b] S. Mohammad Mozaffari. Planetary latitudes in medieval Islamic astronomy: an analysis of the non-Ptolemaic latitude parameter values in the Maragha and Samarqand astronomical traditions. Archive for History of Exact Sciences, 70(5):513– 541, September 2016. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0172-x.

Mozaffari:2018:AMS

[Moz18a] S. Mohammad Mozaffari. An analysis of medieval solar theories. Archive for History of Exact Sciences, 72(2):191–243, March 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Mozaffari:2018:AOM

[Moz18b] S. Mohammad Mozaffari. Astronomical observations at the Maragha observatory in the 1260s–1270s. Archive for History of Exact Sciences, 72(6):591–641, November 2018. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Mozaffari:2019:IAF

[Moz19] S. Mohammad Mozaffari. Ibn al-Fahhad and the Great Conjunction of 1166 AD. Archive for History of Exact Sciences, 73(5): 517–549, September 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Mozaffari:2022:MCS

[Moz22] S. Mohammad Mozaffari. A mechanical concentric solar model in Khazini’s Mu‘tabar zij. Archive for History of Exact Sciences, 76(5):513–529, September 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00292-9.

Mozaffari:2024:HST

[Moz24a] S. Mohammad Mozaffari. Hipparchus’ selenelion and two pairs of lunar eclipses revisited. Archive for History of Exact Sciences, 78(4):361–373, July 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-024-00330-8.

Mozaffari:2024:IAZ

[Moz24b] S. Mohammad Mozaffari. Ibn al-Zarqalluh’s discovery of the annual equation of the Moon. Archive for History of Exact Sciences, 78(3):271–304, May 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-023-00323-z.

Ma:2022:PDH

[MP22] Minghui Ma and Ahti-Veikko Pietarinen. Peirce’s dragon-head logic (R 501, 1901). Archive for History of Exact Sciences, 76 (3):261–317, May 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00285-0.

Mawhin:2010:SDE

[MR10] Jean Mawhin and André Ronveaux. Schrödinger and Dirac equations for the hydrogen atom, and Laguerre polynomials. Archive for History of Exact Sciences, 64(4):429–460, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Moyon:2015:LFO

[MS15a] Marc Moyon and Maryvonne Spiesser. L’arithmétique des fractions dans l’oeuvre de Fibonacci: fondements & usages. Archive for History of Exact Sciences, 69(4):391–427, July 2015. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
015-0155-y.

Mozaffari:2015:SLO

[MS15b] S. Mohammad Mozaffari and John M. Steele. Solar and lunar observations at Istanbul in the 1570s. Archive for History of Exact Sciences, 69(4):343–362, July 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-015-0153-0.

McKinzie:1997:HLE

[MT97] Mark McKinzie and Curtis Tuckey. Hidden lemmas in Euler’s summation of the reciprocals of the squares. Archive for History of Exact Sciences, 51(1):29–57, March 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=51&issue=1&spage=29.

Maurey:2013:TRV

[MT13] Bernard Maurey and Jean-Pierre Tacchi. Sur le théorème de recouvrement de Vitali. (French) [On Vitali’s covering theorem]. Archive for History of Exact Sciences, 67(1):81–118, January 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://link.springer.com/article/10.
1007/s00407-012-0109-6.

Mueller:1970:HET

[Mue70] Ian Mueller. Homogeneity in Eudoxus’s theory of proportion. Archive for History of Exact Sciences, 7(1):1–6, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=7&issue=1&spage=
1.

Mukherji:1974:HMT

[Muk74] Viśvapriya Mukherji. A history of the meson theory of nuclear forces from 1935 to 1952. Archive for History of Exact Sciences, 13(1):27–102, March 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=13&issue=1&spage=27.

Miolo:2024:LCE

[MZ24] Laure Miolo and Stefan Zieme. Lewis Caerleon and the equation of time: tabular astronomical practices in late fifteenthcentury England. Archive for History of Exact Sciences, 78 (2):183–243, March 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-023-00324-y.

Nabonnand:2008:TWS

[Nab08] Philippe Nabonnand. La théorie des Würfe [casts] de von Staudt — une irruption de lálgèbre dans la géométrie pure. (French) [The casts of von Staudt theory — an eruption of algebra in pure geometry]. Archive for History of Exact Sciences, 62(3):201–242, May 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=3&spage=201.

Nauenberg:1994:NEC

[Nau94] Michael Nauenberg. Newton’s early computational method for dynamics. Archive for History of Exact Sciences, 46(3):221–252, February 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=3&spage=221.

Nauenberg:2010:EAC

[Nau10] M. Nauenberg. The early application of the calculus to the inverse square force problem. Archive for History of Exact Sciences, 64(3):269–300, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Nauenberg:2011:PBP

[Nau11] Michael Nauenberg. Proposition 10, Book 2, in the Principia, revisited. Archive for History of Exact Sciences, 65(5):567–587, September 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=5&spage=567.

Nauenberg:2014:OMF

[Nau14] Michael Nauenberg. Orbital motion and force in Newton’s Principia; the equivalence of the descriptions in Propositions 1 and 6. Archive for History of Exact Sciences, 68(2):179–205, March 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://link.springer.com/article/10.
1007/s00407-014-0136-6.

Nauenberg:2015:SLS

[Nau15] Michael Nauenberg. Solution to the long-standing puzzle of Huygens’ “anomalous suspension”. Archive for History of Exact Sciences, 69(3):327–341, May 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-015-0152-1.

Navarro:1998:GEF

[Nav98] Luis Navarro. Gibbs, Einstein and the foundations of statistical mechanics. Archive for History of Exact Sciences, 53(2):147– 180, June 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=2&spage=147.

Nadal:1984:CHS

[NB84] R. Nadal and J.-P. Brunet. Le “Commentaire” d’Hipparque I. La sphère mobile. Archive for History of Exact Sciences, 29(3):201–236, September 1984. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=29&issue=3&spage=201.

Nadal:1989:CHI

[NB89] R. Nadal and J.-P. Brunet. Le “Commentaire” d’Hipparque II. Position de 78 étoiles. (French) [Hipparchus’ Commentary II. Position of 78 stars]. Archive for History of Exact Sciences, 40(4):305–354, December 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=40&issue=4&spage=305.

Nemenyi:1962:MCI

[Nem62] P. F. Neményi. The main concepts and ideas of fluid dynamics in their historical development. Archive for History of Exact Sciences, 2(1):52–86, August 1962. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=1&spage=52.

Netz:1999:ATC

[Net99] Reviel Netz. Archimedes transformed: The case of a result stating a maximum for a cubic equation. Archive for History of Exact Sciences, 54(1):1–47, April 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=1&spage=1.

Netuka:2020:LCC

[Net20] Ivan Netuka. Lebesgue’s criticism of Carl Neumann’s method in potential theory. Archive for History of Exact Sciences, 74 (1):77–108, January 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Neuenschwander:1973:BFG

[Neu73a] E. Neuenschwander. Beiträge zur Frühgeschichte der griechischen Geometrie I. (German) [Contributions to the early history of Greek geometry I]. Archive for History of Exact Sciences, 11(2–3):127–133, June 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=11&issue=2&spage=127.

Neuenschwander:1973:EVB

[Neu73b] E. Neuenschwander. Die ersten vier Bücher der Elemente Euklids. (German) [The first four books of Euclid’s Elements]. Archive for History of Exact Sciences, 9(4–5):325–380, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

9&issue=4&spage=325. Untersuchungen über den mathematischen Aufbau, die Zitierweise und die Entstehungsgeschichte.

Neuenschwander:1974:SBE

[Neu74] E. Neuenschwander. Die stereometrischen Bücher der Elemente Euklids. (German) [The stereometric books of Euclid’s Elements]. Archive for History of Exact Sciences, 14(2):91– 125, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

14&issue=2&spage=91. Untersuchungen über den mathematischen Aufbau und die Entstehungsgeschichte.

Neuenschwander:1978:NCP

[Neu78] E. Neuenschwander. Der Nachlaß von Casorati (1835–1890) in Pavia. (German) [The deduction of Casorati (1835–1890) in Pavia]. Archive for History of Exact Sciences, 19(1):1–89, March 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=1&spage=1.

Neuenschwander:1981:WZF

[Neu81] E. Neuenschwander. Über die Wechselwirkungen zwischen der französischen Schule, Riemann und Weierstraß. Eine Übersicht mit zwei Quellenstudien. (German) [On the reciprocal effects between the French school, Riemann and Weierstrass. An overview with two source studies]. Archive for History of Exact Sciences, 24(3):221–255, September 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=24&issue=3&spage=221.

Neumann:1997:EGT

[Neu97] Olaf Neumann. Die Entwicklung der Galois–Theorie zwischen Arithmetik und Topologie (1850 bis 1960). (German)

[The development of Galois theory between arithmetic and topology (1850 to 1960)]. Archive for History of Exact Sciences, 50(3–4):291–329, September 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=50&issue=3&spage=291.

Neumann:2006:CPG

[Neu06] Peter M. Neumann. The concept of primitivity in group theory and the second memoir of Galois. Archive for History of Exact Sciences, 60(4):379–429, July 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=60&issue=4&spage=379.

Niss:2005:HLI

[Nis05] Martin Niss. History of the Lenz–Ising Model 1920–1950: From ferromagnetic to cooperative phenomena. Archive for History of Exact Sciences, 59(3):267–318, March 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=3&spage=267.

Niss:2009:HLI

[Nis09] Martin Niss. History of the Lenz–Ising model 1950–1965: from irrelevance to relevance. Archive for History of Exact Sciences, 63(3):243–287, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Niss:2011:HLI

[Nis11] Martin Niss. History of the Lenz–Ising model 1965–1971: the role of a simple model in understanding critical phenomena. Archive for History of Exact Sciences, 65(6):625–658, November 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=6&spage=625.

Norton:1992:ENE

[Nor92] John D. Norton. Einstein, Nordström and the early demise of scalar, Lorentz-covariant theories of gravitation. Archive for History of Exact Sciences, 45(1):17–94, March 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=45&issue=1&spage=17.

Norton:2004:EIG

[Nor04] John D. Norton. Einstein’s investigations of Galilean covariant electrodynamics prior to 1905. Archive for History of Exact Sciences, 59(1):45–105, November 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=1&spage=45.

Nothaft:2017:CTM

[Not17] C. Philipp E. Nothaft. Criticism of trepidation models and advocacy of uniform precession in medieval Latin astronomy. Archive for History of Exact Sciences, 71(3):211–244, May 2017. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Nothaft:2018:JHT

[Not18] C. Philipp E. Nothaft. John Holbroke, the Tables of Cambridge, and the “true length of the year”: a forgotten episode in fifteenth-century astronomy. Archive for History of Exact Sciences, 72(1):63–88, January 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Nothaft:2023:MAG

[Not23] C. Philipp E. Nothaft. Measurements of altitude and geographic latitude in Latin astronomy, 1100–1300. Archive for History of Exact Sciences, 77(6):537–577, November 2023. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-023-00312-2.

Nothaft:2024:GLL

[Not24] C. Philipp E. Nothaft. Geographic longitude in Latin Europe during the twelfth and thirteenth centuries. Archive for History of Exact Sciences, 78(1):29–65, January 2024. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-023-00316-y.

Navarro:2004:PEN

[NP04] Luis Navarro and Enric Pérez. Paul Ehrenfest on the necessity of quanta (1911): Discontinuity, quantization, corpuscularity, and adiabatic invariance. Archive for History of Exact Sciences, 58(2):97–141, January 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=2&spage=97.

Navarro:2006:PEG

[NP06] Luis Navarro and Enric Pérez. Paul Ehrenfest: The genesis of the Adiabatic Hypothesis, 1911–1914. Archive for History of Exact Sciences, 60(2):209–267, March 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=60&issue=2&spage=209.

Nadal:2004:EAC

[NTP04a] Robert Nadal, Abdelkaddous Taha, and Pierre Pinel. Erratum: Le contenu astronomique des sphériques de Ménélaos: (French) [Erratum: The astronomical content of the Spherics of Menelaos]. Archive for History of Exact Sciences, 58(5): 437, July 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
58&issue=5&spage=437. See [NTP04b].

Nadal:2004:CAS

[NTP04b] Robert Nadal, Abdelkaddous Taha, and Pierre Pinel. Le contenu astronomique des sphériques de Ménélaos. (French) [The astronomical content of the spherics of menelaos]. Archive for History of Exact Sciences, 58(5):381–436, July 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=5&spage=381.  See erratum

[NTP04a].

Oaks:2009:PEA

[Oak09] Jeffrey A. Oaks. Polynomials and equations in Arabic algebra. Archive for History of Exact Sciences, 63(2):169–203, 2009. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Oaks:2018:FVR

[Oak18] Jeffrey A. Oaks. François Viète’s revolution in algebra. Archive for History of Exact Sciences, 72(3):245–302, May 2018. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Oswald:2014:CCF

[OS14] Nicola M. R. Oswald and Jörn J. Steuding. Complex continued fractions: early work of the brothers Adolf and Julius Hurwitz. Archive for History of Exact Sciences, 68(4):499–528, July 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-014-0135-7.

Osada:2013:INQ

[Osa13] Naoki Osada. Isaac Newton’s ‘Of Quadrature by Ordinates’. Archive for History of Exact Sciences, 67(4):457–476, July 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-013-0117-1.

Ossendrijver:2018:BTT

[Oss18] Mathieu Ossendrijver. Bisecting the trapezoid: tracing the origins of a Babylonian computation of Jupiter’s motion. Archive for History of Exact Sciences, 72(2):145–189, March 2018. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Oswald:2017:RBM

[Osw17] Nicola M. R. Oswald. On a relation between modular functions and Dirichlet series: found in the estate of Adolf Hurwitz. Archive for History of Exact Sciences, 71(4):345–361, July 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Paiow:1971:MS

[Pai71] Michael Paiow. Die Mathematische Staatsstelle. I. (German) [The mathematical public office. I]. Archive for History of Exact Sciences, 8(1–2):1–8, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=8&issue=1&spage=1.

Paiow:1974:MSI

[Pai74] Michael Paiow. Die Mathematische Staatsstelle. II. (German) [The mathematical public office. II]. Archive for History of Exact Sciences, 12(2):174–185, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=12&issue=2&spage=174.

Paiow:1977:MSI

[Pai77] Michael Paiow. Die Mathematische Staatsstelle. III. (German) [The mathematical public office. III]. Archive for History of Exact Sciences, 18(1):1–26, March 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=18&issue=1&spage=1.

Paiow:1982:MT

[Pai82] M. E. Paiow. Die Mathematische Theaetetsstelle. (German) [The mathematical Theaetet position]. Archive for History of Exact Sciences, 27(1):87–99, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=1&spage=87. With notes by M.

L. Gasparov.

Palladino:1985:ICO

[Pal85] Franco Palladino. Individuo e classe: l’origine e il significato di ∈ nel simbolismo di Peano. (Italian) [Individual and class: the origin and meaning of ∈ in Peano’s notation]. Archive for History of Exact Sciences, 34(3):241–256, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
34&issue=3&spage=241.

Palmieri:1998:REG

[Pal98] Paolo Palmieri. Re-examining Galileo’s Theory of Tides. Archive for History of Exact Sciences, 53(3–4):223–375, November 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=3&spage=223.

Palmieri:2001:OEC

[Pal01] Paolo Palmieri. The obscurity of the equimultiples Clavius’ and Galileo’s foundational studies of Euclid’s Theory of Proportions. Archive for History of Exact Sciences, 55(6):555– 597, June 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=6&spage=555.

Palmieri:2005:CDG

[Pal05] Paolo Palmieri. The cognitive development of Galileo’s Theory of Buoyancy. Archive for History of Exact Sciences, 59(2):189– 222, January 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=2&spage=189.

Palmieri:2006:NLG

[Pal06] P. Palmieri. A new look at Galileo’s search for mathematical proofs. Archive for History of Exact Sciences, 60(3):285– 317, May 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
60&issue=3&spage=285.

Palmieri:2008:BCE

[Pal08] Paolo Palmieri. Breaking the circle: the emergence of Archimedean mechanics in the late Renaissance. Archive for History of Exact Sciences, 62(3):301–346, May 2008. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=62&issue=3&spage=301.

Palmieri:2009:SCP

[Pal09] Paolo Palmieri. Superposition: on Cavalieri’s practice of mathematics. Archive for History of Exact Sciences, 63(5):471–495, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Papadopoulos:2017:NAT

[Pap17] Athanase Papadopoulos. Nicolas-Auguste Tissot: a link between cartography and quasiconformal theory. Archive for History of Exact Sciences, 71(4):319–336, July 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Parshall:1985:JHM

[Par85] Karen Hunger Parshall. Joseph H. M. Wedderburn and the structure theory of algebras. Archive for History of Exact Sciences, 32(3–4):223–349, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=32&issue=3&spage=223.

Parshall:1988:AFS

[Par88] Karen Hunger Parshall. America’s first school of mathematical research: James Joseph Sylvester at The Johns Hopkins University 1876–1883. Archive for History of Exact Sciences, 38(2): 153–196, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
38&issue=2&spage=153.

Provost:2006:RPT

[PB06] Jean-Pierre Provost and Christian Bracco. La Relativité de Poincaré de 1905 et les transformations actives. (French) [Poincaré’s Relativity of 1905 and active transformations]. Archive for History of Exact Sciences, 60(3):337–351, May 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=60&issue=3&spage=
337.

Pedersen:2000:WFT

[Ped00] Kurt Møller Pedersen. Water-filled telescopes and the prehistory of Fresnel’s ether dragging. Archive for History of Exact Sciences, 54(6):499–564, June 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=6&spage=499.

Pepper:1968:HCM

[Pep68] Jon V. Pepper. Harriot’s calculation of the meridional parts as logarithmic tangents. Archive for History of Exact Sciences, 4(5):359–413, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=4&issue=5&spage=359.

Perez:2009:EAT

[Pér09a] Enric Pérez. Ehrenfest’s adiabatic theory and the old quantum theory, 1916–1918. Archive for History of Exact Sciences, 63(1):81–125, January 2009. CODEN AH-ESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=63&issue=1&spage=81. See

erratum [Pér09b].

Perez:2009:EEA

[Pér09b] Enric Pérez. Erratum: Ehrenfest’s adiabatic hypothesis and the old quantum theory, 1916–1918. Archive for History of Exact Sciences, 63(1):127, January 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=1&spage=127. See [Pér09a].

Peters:1961:PGK

[Pet61] Wilhelm Servatius Peters. Das Parallelenproblem bei A. G. Kaestner Zur Parallelenforschung im 18. Jahrhundert. (German) [The parallel problem of A. G. Kaestner for parallel research in the 18^{th} century]. Archive for History of Exact Sciences, 1(5):480–487, October 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=5&spage=480. Zur Parallelen-

forschung im 18. Jahrhundert.

Peters:1963:BKK

[Pet63] Wilhelm Servatius Peters. Zum Begriff der Konstruierbarkeit bei I. Kant. (German) [On the term constructableness of I. Kant]. Archive for History of Exact Sciences, 2(2):153–167, August 1963. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
2&issue=2&spage=153.

Petrova:1987:HDS

[Pet87] Svetlana S. Petrova. Heaviside and the development of the symbolic calculus. Archive for History of Exact Sciences, 37(1): 1–23, March 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
37&issue=1&spage=1.

Petruccioli:2011:CBU

[Pet11] Sandro Petruccioli. Complementarity before uncertainty. Archive for History of Exact Sciences, 65(6):591–624, November 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=6&spage=591.

Petruccioli:2014:CPV

[Pet14] Sandro Petruccioli. Correspondence principle versus Planck-type theory of the atom. Archive for History of Exact Sciences, 68(5): 599–639, September 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-014-0137-5.

Pottage:1974:MQG

[Pot74] John Pottage. The mensuration of quadrilaterals and the generation of Pythagorean triads: a mathematical, heuristical and historical study with special reference to Brahmagupta’s rules. Archive for History of Exact Sciences, 12(4):299–354, August 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
12&issue=4&spage=299.

Pourciau:1992:NSO

[Pou92a] Bruce Pourciau. Newton’s solution of the one-body problem. Archive for History of Exact Sciences, 44(2):125–146, June 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
44&issue=2&spage=125.

Pourciau:1992:RP

[Pou92b] Bruce Pourciau. Radical principia. Archive for History of Exact Sciences, 44(4):331–363, December 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=4&spage=331.

Pourciau:1998:PML

[Pou98] Bruce Pourciau. The preliminary mathematical lemmas of Newton’s Principia. Archive for History of Exact Sciences, 52(3):279–295, February 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=52&issue=3&spage=279.

Pourciau:2001:ION

[Pou01] Bruce Pourciau. The integrability of ovals: Newton’s Lemma 28 and its counterexamples. Archive for History of Exact Sciences, 55(5):479–499, April 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=5&spage=479.

Pourciau:2003:NAP

[Pou03] Bruce Pourciau. Newton’s argument for Proposition 1 of the Principia. Archive for History of Exact Sciences, 57(4):267– 311, May 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=4&spage=267.

Pourciau:2004:IBE

[Pou04] Bruce Pourciau. The importance of being equivalent: Newton’s two models of one-body motion. Archive for History of Exact Sciences, 58(4):283–321, May 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=4&spage=283.

Pourciau:2006:NIN

[Pou06] Bruce Pourciau. Newton’s interpretation of Newton’s Second Law. Archive for History of Exact Sciences, 60(2):157–207, March 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
60&issue=2&spage=157.

Pourciau:2009:PIB

[Pou09] Bruce Pourciau. Proposition II (Book I) of Newton’s principia. Archive for History of Exact Sciences, 63(2):129–167, 2009. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Pourciau:2020:PSL

[Pou20] Bruce Pourciau. The Principia’s second law (as Newton understood it) from Galileo to Laplace. Archive for History of Exact Sciences, 74(3):183–242, May 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-019-00242-y.

Peruzzi:2011:QSU

[PR11] Giulio Peruzzi and Matteo Realdi. The quest for the size of the universe in early relativistic cosmology (1917–1930). Archive for History of Exact Sciences, 65(6):659–689, November 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=6&spage=659.

Peruzzi:2023:HGC

[PR23] Giulio Peruzzi and Valentina Roberti. Helmholtz and the geometry of color space: gestation and development of Helmholtz’s line element. Archive for History of Exact Sciences, 77(2): 201–220, March 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-023-00304-2.

Proverbio:1988:RGB

[Pro88] Edoardo Proverbio. R. G. Boscovich’s determination of instrumental errors in observation. Archive for History of Exact Sciences, 38(2):135–152, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=38&issue=2&spage=135.

Perez:2010:EQT

[PS10] Enric Pérez and Tilman Sauer. Einstein’s quantum theory of the monatomic ideal gas: non-statistical arguments for a new statistics. Archive for History of Exact Sciences, 64(5):561–612, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Palladino:1996:LEB

[PT96] Franco Palladino and Rossana Tazzioli. Le lettere di Eugenio Beltrami nella corrispondenza di Ernesto Cesàro. (Italian) [Letters of Eugene Beltrami in the correspondence of Ernesto Cesàro]. Archive for History of Exact Sciences, 49(4):321–353, December 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
49&issue=4&spage=321.

Pyenson:1977:HME

[Pye77] Lewis Pyenson. Hermann Minkowski and Einstein’s Special Theory of Relativity. Archive for History of Exact Sciences, 17(1): 71–95, May 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

17&issue=1&spage=71. With an appendix of Minkowski’s “Funktiontheorie” manuscript.

Pyenson:1979:PSM

[Pye79] Lewis Pyenson. Physics in the shadow of mathematics: The Göttingen electron-theory seminar of 1905. Archive for History of Exact Sciences, 21(1):55–89, March 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=1&spage=55.

Pyenson:1982:RLW

[Pye82] Lewis Pyenson. Relativity in late Wilhelmian Germany: The appeal to a preestablished harmony between mathematics and physics. Archive for History of Exact Sciences, 27(2):137– 155, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=2&spage=137.

Rabouin:2020:LSI

[RA20] David Rabouin and Richard T. W. Arthur. Leibniz’s syncategorematic infinitesimals II: their existence, their use and their role in the justification of the differential calculus. Archive for History of Exact Sciences, 74(5):401–443, September 2020. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Rabinovitch:1970:RLB

[Rab70] Nachum L. Rabinovitch. Rabbi Levi Ben Gershon and the origins of mathematical induction. Archive for History of Exact Sciences, 6(3):237–248, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=3&spage=237.

Rabinovitch:1974:EAE

[Rab74] Nachum L. Rabinovitch. Early antecedents of error theory. Archive for History of Exact Sciences, 13(4):348–358, December 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
13&issue=4&spage=348.

Rashed:1970:OGD

[Ras70] Roshdi Rashed. Optique Géometrique et doctrine optique chez Ibn Al Haytham. Archive for History of Exact Sciences, 6(4):271–298, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=4&spage=271.

Rasche:1971:GBS

[Ras71] G. Rasche. Zur Geschichte des Begriffes “Isospin”. (German) [On the history of the term “isospin”]. Archive for History of Exact Sciences, 7(4):257–276, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=7&issue=4&spage=257.

Rashed:1972:LMA

[Ras72] Roshdi Rashed. L’induction mathématique: al-Karaji, as-Samaw’al. (French) [Mathematical induction: al-Karaji, as-Samaw’al]. Archive for History of Exact Sciences, 9(1):1–21, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
9&issue=1&spage=1.

Rashed:1974:REN

[Rās74] Roshdī Rāshed. Résolution des équations numériques et algèbre: Saraf-al-din al-Tusi, Viète. (French) [Resolution of numerical equations and algebra: Šaraf-al-din al-Tūsī, Viète]. Archive for History of Exact Sciences, 12(3):244–290, September 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
12&issue=3&spage=244.

Rashed:1978:LRI

[Rās78] Roshdī Rāshed. L’extraction de la racine nième et l’invention des fractions Décimales (XI^{e}–XII^{e} siècles). (French) [The extraction of the n^{th} root and the invention of decimal fractions (11^{th}–12^{th} centuries)]. Archive for History of Exact Sciences, 18(3):191–243, September 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=18&issue=3&spage=191.

Rashed:1980:IAH

[Ras80] Roshdi Rashed. Ibn al-Haytham et le théorème de Wilson. (French) [Ibn al-Haytham and Wilson’s Theorem]. Archive for History of Exact Sciences, 22(4):305–321, December 1980. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=22&issue=4&spage=305.

Rashed:1983:NAP

[Ras83] R. Rashed. Nombres amiables, parties aliquotes et nombres figurés aux XIIIèm et XIVèm siècles. (French) [friendly numbers, aliquot parts and figured numbers in the 13^{th} and 14^{th} centuries]. Archive for History of Exact Sciences, 28(2):107– 147, June 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=2&spage=107.

Rawlins:1982:ESN

[Raw82] Dennis Rawlins. The Eratosthenes–Strabo Nile Map. Is it the earliest surviving instance of spherical cartography? Did it supply the 5000 stades arc for Eratosthenes’ experiment? Archive for History of Exact Sciences, 26(3):211–219, September 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
26&issue=3&spage=211.

Raynaud:2014:BSC

[Ray14] Dominique Raynaud. Building the stemma codicum from geometric diagrams. Archive for History of Exact Sciences, 68 (2):207–239, March 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-013-0134-0.

Raynaud:2020:MAT

[Ray20] Dominique Raynaud. Mathématiques et architecture: le tracé de l’entasis par Nicolas–François Blondel. Archive for History of Exact Sciences, 74(5):445–468, September 2020. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Rybski:2023:ALZ

[RC23] Diego Rybski and Antonio Ciccone. Auerbach, Lotka, and Zipf: pioneers of power-law city-size distributions. Archive for History of Exact Sciences, 77(6):601–613, November 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-023-00314-0.

Romo:1994:FIM

[RD94] José Romo and Manuel G. Doncel. Faraday’s initial mistake concerning the direction of induced currents, and the manuscript of series I of his researches. Archive for History of Exact Sciences, 47(4):291–385, December 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=47&issue=4&spage=291.

Rommevaux:2001:RHD

[RDV01] S. Rommevaux, A. Djebbar, and B. Vitrac. Remarques sur l’histoire du texte des Éléments d’Euclide. (French) [Remarks on the history of the text of Euclid’s Elements]. Archive for History of Exact Sciences, 55(3):221–295, January 2001. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=55&issue=3&spage=221.

Recasens:1994:JZC

[Rec94] E. Recasens. J. Zaragosa’s centrum minimum, an early version of barycentric geometry. Archive for History of Exact Sciences, 46(4):285–320, December 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=46&issue=4&spage=285.

Reich:1973:GDG

[Rei73] Karin Reich. Die Geschichte der Differentialgeometrie von Gauß bis Riemann (1828–1868). (German) [The history of differential geometry from Gauss to Riemann]. Archive for History of Exact Sciences, 11(4):273–376, December 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=11&issue=4&spage=273.

Reich:1989:EVD

[Rei89] Karin Reich. Das Eindringen des Vektorkalküls in die Differentialgeometrie. (German) [The penetration of vector calculus in differential geometry]. Archive for History of Exact Sciences, 40(3):275–303, September 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=40&issue=3&spage=275.

Reich:1992:LCP

[Rei92] Karin Reich. Levi-Civitasche Parallelverschiebung, affiner Zusammenhang, Übertragungsprinzip: 1916/17–1922/23. (German) [Levi-Civita translation, affine connection, transmission principle]. Archive for History of Exact Sciences, 44(1):77– 105, March 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
44&issue=1&spage=77.

Reich:2003:CGC

[Rei03] Karin Reich. Cauchy und Gauß. Cauchys Rezeption im Umfeld von Gauß. (German) [Cauchy and Gauss. Cauchy’s reception in the field surrounding Gauss]. Archive for History of Exact Sciences, 57(6):433–463, September 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=57&issue=6&spage=433.

Renn:1997:ECD

[Ren97] Jürgen Renn. Einstein’s controversy with Drude and the origin of statistical mechanics: a new glimpse from the “Love Letters”. Archive for History of Exact Sciences, 51(4):315–354, December 1997. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
51&issue=4&spage=315.

Raynaud:2019:ANC

[RGM19] Dominique Raynaud, Samuel Gessner, and Bernardo Mota. Andalò di Negro’s De compositione astrolabii: a critical edition with English translation and notes. Archive for History of Exact Sciences, 73(6):551–617, November 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Riddell:1978:LVM

[Rid78] R. C. Riddell. The latitudes of Venus and Mercury in the Almagest. Archive for History of Exact Sciences, 19(2):95– 111, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=2&spage=95.

Riddell:1979:EME

[Rid79] R. C. Riddell. Eudoxan mathematics and the Eudoxan spheres. Archive for History of Exact Sciences, 20(1):1–19, March 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
20&issue=1&spage=1.

Riddell:1980:PDP

[Rid80] R. C. Riddell. Parameter disposition in pre-Newtonian planetary theories. Archive for History of Exact Sciences, 23(2):87– 157, June 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=2&spage=87.

Rittberg:2015:HWC

[Rit15] Colin J. Rittberg. How Woodin changed his mind: new thoughts on the Continuum Hypothesis. Archive for History of Exact Sciences, 69(2):125–151, March 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-014-0142-8.

RileyKnoedler:2021:UMJ

[RKH^{+}21] Molly Riley Knoedler, Julianna C. Kostas, Caroline Mary Hogan, Harper Kerkhoff, and Chad M. Topaz. An unpublished manuscript of John von Neumann on shock waves in boostered detonations: historical context and mathematical analysis. Archive for History of Exact Sciences, 75(1):83–108, January 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s00407-020-00258-9.pdf.

Rocci:2021:BRV

[Roc21] Alessio Rocci. Back to the roots of vector and tensor calculus: Heaviside versus Gibbs. Archive for History of Exact Sciences, 75(4):369–413, July 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-020-00264-x.

Rodriguez:2015:FRE

[Rod15] Laura Rodríguez. Frigyes Riesz and the emergence of general topology. Archive for History of Exact Sciences, 69(1):55–102, January 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-014-0144-6.

Ronchi:1960:COM

[Ron60] Vasco Ronchi. Classical optics is a mathematical science. Archive for History of Exact Sciences, 1(2):160–171, December 1960. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=1&issue=2&spage=
160.

Roque:1992:MSN

[Roq92] Xavier Roqué. Møller scattering: a neglected application of early quantum electrodynamics. Archive for History of Exact Sciences, 44(3):197–264, September 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=44&issue=3&spage=197.

Roque:2011:STP

[Roq11] Tatiana Roque. Stability of trajectories from Poincaré to Birkhoff: approaching a qualitative definition. Archive for History of Exact Sciences, 65(3):295–342, May 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=3&spage=295.

Rosenfeld:1965:NLG

[Ros65] L. Rosenfeld. Newton and the Law of Gravitation. Archive for History of Exact Sciences, 2(5):365–386, August 1965. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=5&spage=
365.

Rosenfeld:1969:NVA

[Ros69] L. Rosenfeld. Newton’s views on aether and gravitation. Archive for History of Exact Sciences, 6(1):29–37, January 1969. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=6&issue=1&spage=
29.

Rosenfeld:1971:MIH

[Ros71] L. Rosenfeld. Men and ideas in the history of atomic theory. Archive for History of Exact Sciences, 7(2):69–90, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=7&issue=2&spage=
69.

Rosen:1981:DTB

[Ros81] Edward Rosen. In defense of Tycho Brahe. Archive for History of Exact Sciences, 24(4):257–265, December 1981. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=24&issue=4&spage=257.

Rosso:2015:FCW

[Ros15] Riccardo Rosso. Felice Casorati’s work on finite differences and its influence on Salvatore Pincherle. Archive for History of Exact Sciences, 69(2):153–215, March 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.

springer.com/article/10.1007/s00407-014-0143-7.

Rouxel:1980:LMD

[Rou80] Bernard Rouxel. L’OEuvre Mathématique d’Albert Ribaucour. (French) [The mathematical work of Albert Ribaucour]. Archive for History of Exact Sciences, 23(2):159–177, June 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=2&spage=159.

Rowe:2024:FKE

[Row24] David E. Rowe. Felix Klein’s early contributions to anschauliche Geometrie. Archive for History of Exact Sciences, 78(4):401– 477, July 2024. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL https://link.springer.com/
article/10.1007/s00407-024-00329-1.

Roberti:2023:HLC

[RP23] Valentina Roberti and Giulio Peruzzi. The Helmholtz legacy in color metrics: Schrödinger’s color theory. Archive for History of Exact Sciences, 77(6):615–635, November 2023. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-023-00317-x.

Rajagopal:1978:USM

[RR78] C. T. Rajagopal and M. S. Rangachari. On an untapped source of medieval Keralese mathematics. Archive for History of Exact Sciences, 18(2):89–102, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=18&issue=2&spage=89.

Rajagopal:1986:MKM

[RR86] C. T. Rajagopal and M. S. Rangachari. On medieval Kerala mathematics. Archive for History of Exact Sciences, 35(2):91– 99, June 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=2&spage=91.

Rivolo:1998:CDR

[RS98] M. T. Rivolo and A. Simi. Il calcolo delle radici quadrate e cubiche in Italia da Fibonacci a Bombelli. (Italian) [The calculation of square and cube roots in Italy from Fibonacci to Bombelli]. Archive for History of Exact Sciences, 52(2):161–193, February 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
52&issue=2&spage=161.

Roughton:2004:LBN

[RSW04] N. A. Roughton, J. M. Steele, and C. B. F. Walker. A late Babylonian normal and Ziqpu star text. Archive for History of Exact Sciences, 58(6):537–572, September 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=6&spage=537.

Ruffner:1963:RGN

[Ruf63] James A. Ruffner. Reinterpretation of the genesis of Newton’s “law of Cooling”. Archive for History of Exact Sciences, 2(2):138–152, August 1963. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=2&spage=138.

Ruffner:2000:NPC

[Ruf00] J. A. Ruffner. Newton’s propositions on comets: Steps in transition, 1681–84. Archive for History of Exact Sciences, 54(4):259– 277, January 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=4&spage=259.

Ruffner:2012:NGR

[Ruf12] J. A. Ruffner. Newton’s De gravitatione: a review and reassessment. Archive for History of Exact Sciences, 66(3):241– 264, May 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
66&issue=3&spage=241.

Ruffner:2013:SSN

[Ruf13] J. A. Ruffner. The snare of simplicity: the Newton–Flamsteed correspondence revisited. Archive for History of Exact Sciences, 67(4):415–455, July 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-013-0116-2.

Russo:1998:DFG

[Rus98] Lucio Russo. The definitions of fundamental geometric entities contained in Book I of Euclid’s Elements. Archive for History of Exact Sciences, 52(3):195–219, February 1998. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=52&issue=3&spage=195.

Redei:2012:HIP

[RW12] Miklós Rédei and Charlotte Werndl. On the history of the isomorphism problem of dynamical systems with special regard to von Neumann’s contribution. Archive for History of Exact Sciences, 66(1):71–93, January 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=1&spage=71.

Sabra:1982:IAH

[Sab82] A. I. Sabra. Ibn al-Haytham’s lemmas for solving “Alhazen’s problem”. Archive for History of Exact Sciences, 26(4):299–324, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
26&issue=4&spage=299.

Sakellariadis:1982:DEP

[Sak82] Spyros Sakellariadis. Descartes’ experimental proof of the infinite velocity of light and Huygens’ rejoinder. Archive for History of Exact Sciences, 26(1):1–12, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=26&issue=1&spage=1.

Salzer:1974:TLE

[Sal74] Herbert E. Salzer. Two letters from Einstein concerning his distant parallelism field theory. Archive for History of Exact Sciences, 12(1):89–96, March 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=12&issue=1&spage=89.

Sampson:1990:SGT

[Sam90] J. H. Sampson. Sophie Germain and the theory of numbers. Archive for History of Exact Sciences, 41(2):157–161, June 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=2&spage=157.

Sarkar:2012:CSN

[Sar12] Sahotra Sarkar. Complementarity and the selection of nature reserves: algorithms and the origins of conservation planning, 1980–1995. Archive for History of Exact Sciences, 66(4):397– 426, July 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
66&issue=4&spage=397.

Sauer:1999:RDH

[Sau99] Tilman Sauer. The relativity of discovery: Hilbert’s First Note on the Foundations of Physics. Archive for History of Exact Sciences, 53(6):529–575, January 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=53&issue=6&spage=529.

Sauer:2005:EEH

[Sau05] Tilman Sauer. Einstein equations and Hilbert action: What is missing on page 8 of the proofs for Hilbert’s First Communication on the Foundations of Physics? Archive for History of Exact Sciences, 59(6):577–590, October 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=6&spage=577.

Sauer:2007:EET

[Sau07] Tilman Sauer. Einstein and the early Theory of Superconductivity, 1919–1922. Archive for History of Exact Sciences, 61(2):159– 211, March 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=2&spage=159.

Sauer:2008:NGO

[Sau08] Tilman Sauer. Nova Geminorum 1912 and the origin of the idea of gravitational lensing. Archive for History of Exact Sciences, 62(1):1–22, January 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=1&spage=1.

Seidenberg:1980:ROB

[SC80] A. Seidenberg and J. Casey. The ritual origin of the balance. Archive for History of Exact Sciences, 23(3):179–226, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=3&spage=179.

Schneider:1968:MAM

[Sch68] Ivo Schneider. Der Mathematiker Abraham de Moivre (1667– 1754). (German) [The mathematician Abraham de Moivre (1667–1754)]. Archive for History of Exact Sciences, 5(3–4):177– 317, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
5&issue=3&spage=177.

Schneider:1971:DDF

[Sch71] Ivo Schneider. Descartes’ Diskussion der Fermatschen Extremwertmethode — ein Stück Ideengeschichte der Mathematik. (German) [Descartes’ discussion of Fermat’s extreme value method — a piece of the conceptual history of mathematics]. Archive for History of Exact Sciences, 7(5):354–374, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=7&issue=5&spage=
354.

Schroder:1973:WOJ

[Sch73] W. Schröder. Das Wirken von Otto Jesse für die Erforschung der Hochatmosphäre. (German) [Works of Otto Jesse on the study of the upper atmosphere]. Archive for History of Exact Sciences, 11(2–3):267–272, June 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=11&issue=2&spage=267.

Schneider:1974:CEA

[Sch74a] Ivo Schneider. Clausius’ erste Anwendung der Wahrscheinlichkeitsrechnung im Rahmen der atmosphärischen Lichtstreuung. (German) [Clausius’s first application of probability calculations in the context of atmospheric light scattering]. Archive for History of Exact Sciences, 14(2):143–158, June 1974. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=14&issue=2&spage=143.

Schutt:1974:PMV

[Sch74b] Hans-Werner Schütt. Das Problem der Mischkristallisation von Vitriolen vor der Entdeckung des Isomorphismus. (German) [The problem of the mixture crystallization of vitriol before the discovery of isomorphicism]. Archive for History of Exact Sciences, 12(2):186–198, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=12&issue=2&spage=186.

Schneider:1975:RCB

[Sch75] Ivo Schneider. Rudolph Clausius’ Beitrag zur Einführung wahrscheinlichkeitstheoretischer Methoden in die Physik der Gase nach 1856. (German) [Rudolph Clausius’s contribution to the introduction of probabilistic methods to the physics of gases after 1856]. Archive for History of Exact Sciences, 14(3):237–261, September 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=14&issue=3&spage=237.

Schlissel:1977:DAS

[Sch77] Arthur Schlissel. The development of asymptotic solutions of linear ordinary differential equations, 1817–1920. Archive for History of Exact Sciences, 16(4):307–378, December 1977. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=16&issue=4&spage=307.

Scharlau:1982:UAA

[Sch82a] Winfried Scharlau. Unveröffentlichte algebraische Arbeiten Richard Dedekinds aus seiner Göttinger Zeit 1855–1858. (German) [Unpublished algebraic work of Richard Dedekind from his Göttingen years 1855–1858]. Archive for History of Exact Sciences, 27(4):335–367, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=4&spage=335.

Scholz:1982:RFN

[Sch82b] E. Scholz. Riemanns frühe Notizen zum Mannigfaltigkeitsbegriff und zu den Grundlagen der Geometrie. (German) [Riemann’s early notes on the diversity term and on the basis of geometry]. Archive for History of Exact Sciences, 27(3):213–232, September 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=3&spage=213.

Schroder:1982:EWS

[Sch82c] Wilfried Schröder. Emil Wiechert und seine Bedeutung für die Entwicklung der Geophysik zur exakten Wissenschaft. (German) [Emil Wiechert and his significance for the development of geophysics as an exact science]. Archive for History of Exact Sciences, 27(4):369–389, December 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=27&issue=4&spage=369.

Schroder:1984:HAL

[Sch84] Wilfried Schröder. Hendrik Antoon Lorentz und Emil Wiechert (Briefwechsel und Verhältnis der beiden Physiker). (German) [Hendrik Antoon Lorentz and Emil Wiechert (exchange of letters and the relationship of the two physicists)]. Archive for History of Exact Sciences, 30(2):167–187, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=30&issue=2&spage=167.

Schroder:1985:ASE

[Sch85] Wilfried Schröder. Arnold Sommerfeld und Emil Wiechert. (German) [Arnold Sommerfeld and Emil Wiechert]. Archive for History of Exact Sciences, 32(1):77–93, March 1985. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=32&issue=1&spage=77.

Scharlau:1986:EAZ

[Sch86] Winfried Scharlau. Zur Entstehung der Algebraischen Zahlentheorie — ein Bericht eines Augenzeugen. (German) [On the emergence of the algebraic number theory —- a report of an eyewitness]. Archive for History of Exact Sciences, 36(1):63– 74, March 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
36&issue=1&spage=63.

Schroder:1988:AES

[Sch88] Wilfried Schröder. Albert Einstein in seinen Beziehungen zu Mitgliedern der Gesellschaft der Wissenschaften in Göttingen. (German) [Albert Einstein and his relations with members of the Göttingen Scientific Society]. Archive for History of Exact Sciences, 39(2):157–171, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=2&spage=157.

Schubring:2017:SOE

[Sch17] Gert Schubring. Searches for the origins of the epistemological concept of model in mathematics. Archive for History of Exact Sciences, 71(3):245–278, May 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Schironi:2023:ESR

[Sch23] Francesca Schironi. Eudoxus’ simultaneous risings and settings. Archive for History of Exact Sciences, 77(4):423–441, July 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00309-x.

Scriba:1961:LDP

[Scr61] Christoph J. Scriba. Zur Lösung des 2. Debeauneschen Problems durch Descartes Ein Abschnitt aus der Frühgeschichte der inversen Tangentenaufgaben. (German) [On the solution of the 2nd Debeaune problem of Descartes: a section from the early history of the inverse-tangent exercises]. Archive for History of Exact Sciences, 1(4):406–419, July 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=4&spage=406.

Scriba:1963:IMT

[Scr63] Christoph J. Scriba. The inverse method of tangents: a dialogue between Leibniz and Newton (1675–1677). Archive for History of Exact Sciences, 2(2):113–137, August 1963. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=2&spage=113.

Steele:2023:EST

[SdJ23] John Steele and Teije de Jong. An early system A-type scheme for Saturn from Babylon. Archive for History of Exact Sciences, 77(5):501–535, September 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-023-00311-3.

Segre:1980:REG

[Seg80] Michael Segre. The role of experiment in Galileo’s physics. Archive for History of Exact Sciences, 23(3):227–252, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=3&spage=227.

Segre:1994:PAT

[Seg94] Michael Segre. Peano’s axioms in their historical context. Archive for History of Exact Sciences, 48(3–4):201–342, September 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
48&issue=3&spage=201.

Seidenberg:1961:ROG

[Sei61] A. Seidenberg. The ritual origin of geometry. Archive for History of Exact Sciences, 1(5):488–527, October 1961. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?

genre=article&issn=0003-9519&volume=1&issue=5&spage=
488.

Seidenberg:1962:ROC

[Sei62] A. Seidenberg. The ritual origin of counting. Archive for History of Exact Sciences, 2(1):1–40, August 1962. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=1&spage=1.

Seidenberg:1965:SSS

[Sei65] A. Seidenberg. The sixty system of Sumer. Archive for History of Exact Sciences, 2(5):436–440, August 1965. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=5&spage=436.

Seidenberg:1966:RND

[Sei66] A. Seidenberg. Remarks on Nicomedes’ duplication. Archive for History of Exact Sciences, 3(2):97–101, January 1966. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=3&issue=2&spage=
97.

Seidenberg:1972:ASC

[Sei72] A. Seidenberg. On the area of a semi-circle. Archive for History of Exact Sciences, 9(3):171–211, January 1972. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=9&issue=3&spage=
171.

Seidenberg:1975:DEE

[Sei75] A. Seidenberg. Did Euclid’s Elements, Book I, develop geometry axiomatically? Archive for History of Exact Sciences, 14(4):263–295, December 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=14&issue=4&spage=263.

Seidenberg:1976:KWR

[Sei76] A. Seidenberg. km, a widespread root for ten. Archive for History of Exact Sciences, 16(1):1–16, March 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=16&issue=1&spage=1.

Seidenberg:1978:OM

[Sei78] A. Seidenberg. The origin of mathematics. Archive for History of Exact Sciences, 18(4):301–342, December 1978. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=18&issue=4&spage=301.

Seidenberg:1981:ROC

[Sei81] A. Seidenberg. The ritual origin of the circle and square. Archive for History of Exact Sciences, 25(4):269–327, December 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
25&issue=4&spage=269.

Seidenberg:1984:AMT

[Sei84] A. Seidenberg. ”on ancient mathematical terminology”. Archive for History of Exact Sciences, 31(1):1–13, March 1984. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=31&issue=1&spage=1.

Seidenberg:1988:VS

[Sei88] A. Seidenberg. On the volume of a sphere. Archive for History of Exact Sciences, 39(2):97–119, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=2&spage=97.

Seneta:1998:EIP

[Sen98] E. Seneta. Early influences on probability and statistics in the Russian Empire. Archive for History of Exact Sciences, 53(3–4):201–213, November 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=53&issue=3&spage=201.

Sesiano:1977:TEI

[Ses77] Jacques Sesiano. Le traitement des équations indéterminées dans le Bad i fi’l-Hisab d’Abu Bakr Al–Karaji. (French) [The treatment of indeterminate equations in the Bad i fi’l-Hisab d’Abu Bakr Al–Karaji]. Archive for History of Exact Sciences, 17(4):297–379, July 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=4&spage=297.

Sesiano:1985:ANS

[Ses85] Jacques Sesiano. The appearance of negative solutions in mediaeval mathematics. Archive for History of Exact Sciences, 32(2): 105–150, June 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=2&spage=105.

Sesiano:1991:TPN

[Ses91] J. Sesiano. Two problems of number theory in Islamic times. Archive for History of Exact Sciences, 41(3):235–238, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=3&spage=235.

Sesiano:1998:CMM

[Ses98] Jacques Sesiano. Les carrés magiques de Manuel Moschopoulos. (French) [The magics squares of Manuel Moschopoulos]. Archive for History of Exact Sciences, 53(5):377–397, November 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
53&issue=5&spage=377.

Sesiano:2003:CMS

[Ses03] Jacques Sesiano. Construction of magic squares using the knight’s move in Islamic mathematics. Archive for History of Exact Sciences, 58(1):1–20, November 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=1&spage=1.

Schreiber:2008:NLR

[SFS08] Peter Schreiber, Gisela Fischer, and Maria Luise Sternath. New light on the rediscovery of the Archimedean solids during the Renaissance. Archive for History of Exact Sciences, 62(4):457– 467, July 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
62&issue=4&spage=457.

Shapiro:1973:KOS

[Sha73] Alan E. Shapiro. Kinematic optics: a study of the wave theory of light in the seventeenth century. Archive for History of Exact Sciences, 11(2–3):134–266, June 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=11&issue=2&spage=134.

Shafer:1978:NAP

[Sha78] Glenn Shafer. Non-additive probabilities in the work of Bernoulli and Lambert. Archive for History of Exact Sciences, 19(4):309–370, December 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=19&issue=4&spage=309.

Shapiro:1979:NSD

[Sha79] Alan E. Shapiro. Newton’s “achromatic” dispersion law: Theoretical background and experimental evidence. Archive for History of Exact Sciences, 21(2):91–128, June 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=21&issue=2&spage=91.

Sheynin:1971:JHL

[She71a] O. B. Sheynin. J. H. Lambert’s work on probability. Archive for History of Exact Sciences, 7(3):244–256, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=7&issue=3&spage=

244. See corrections and notes [She83].

Sheynin:1971:NCT

[She71b] O. B. Sheynin. Newton and the classical theory of probability. Archive for History of Exact Sciences, 7(3):217–243, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

7&issue=3&spage=217. See corrections and notes [She83].

Sheynin:1972:MTO

[She72] O. B. Sheynin. On the mathematical treatment of observations by L. Euler. Archive for History of Exact Sciences, 9(1):45–56, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

9&issue=1&spage=45. See corrections and notes [She83].

Sheynin:1973:FRS

[She73a] O. B. Sheynin. Finite random sums (a historical essay). Archive for History of Exact Sciences, 9(4–5):275–305, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=9&issue=4&spage=

275. See corrections and notes [She83].

Sheynin:1973:MTA

[She73b] O. B. Sheynin. Mathematical treatment of astronomical observations (a historical essay). Archive for History of Exact Sciences, 11(2–3):97–126, June 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=11&issue=2&spage=97. See corrections and

notes [She83].

Sheynin:1973:RJB

[She73c] O. B. Sheynin. R. J. Boscovich’s work on probability. Archive for History of Exact Sciences, 9(4–5):306–324, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=9&issue=4&spage=

306. See corrections and notes [She83].

Sheynin:1974:PTP

[She74] O. B. Sheynin. On the prehistory of the theory of probability. Archive for History of Exact Sciences, 12(2):97–141, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

12&issue=2&spage=97. See corrections and notes [She83].

Sheynin:1976:PLW

[She76] O. B. Sheynin. P. S. Laplace’s work on probability. Archive for History of Exact Sciences, 16(2):137–187, June 1976. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=16&issue=2&spage=137. See

corrections and notes [She83].

Sheynin:1977:EHT

[She77a] O. B. Sheynin. Early history of the theory of probability. Archive for History of Exact Sciences, 17(3):201–259, September 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

17&issue=3&spage=201. See corrections and notes [She83].

Sheynin:1977:LTE

[She77b] O. B. Sheynin. Laplace’s theory of errors. Archive for History of Exact Sciences, 17(1):1–61, May 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=17&issue=1&spage=1. See corrections and

notes [She83].

Sheynin:1978:DPW

[She78] O. B. Sheynin. S. D. Poisson’s work in probability. Archive for History of Exact Sciences, 18(3):245–300, September 1978. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=18&issue=3&spage=245. See

corrections and notes [She83].

Sheynin:1979:CFG

[She79] O. B. Sheynin. C. F. Gauss and the theory of errors. Archive for History of Exact Sciences, 20(1):21–72, March 1979. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=20&issue=1&spage=21. See

corrections and notes [She83].

Sheynin:1980:HSM

[She80] O. B. Sheynin. On the history of the statistical method in biology. Archive for History of Exact Sciences, 22(4):323–371, December 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

22&issue=4&spage=323. See corrections and notes [She83].

Sheynin:1982:HMS

[She82] O. B. Sheynin. On the history of medical statistics. Archive for History of Exact Sciences, 26(3):241–286, September 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=

26&issue=3&spage=241. See corrections and notes [She83].

Sheynin:1983:CSN

[She83] O. B. Sheynin. Corrections and short notes on my papers. Archive for History of Exact Sciences, 28(2):171–180, June 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=2&spage=171. See [She82].

Sheynin:1984:HSMa

[She84a] O. B. Sheynin. On the history of the statistical method in astronomy. Archive for History of Exact Sciences, 29(2):151– 199, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
29&issue=2&spage=151.

Sheynin:1984:HSMb

[She84b] O. B. Sheynin. On the history of the statistical method in meteorology. Archive for History of Exact Sciences, 31(1):53– 95, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
31&issue=1&spage=53.

Sheynin:1985:HSM

[She85] O. B. Sheynin. On the history of the statistical method in physics. Archive for History of Exact Sciences, 33(4):351–382, December 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
33&issue=4&spage=351.

Sheynin:1986:QS

[She86] O. B. Sheynin. A. Quetelet as a statistician. Archive for History of Exact Sciences, 36(4):281–325, December 1986. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=36&issue=4&spage=281.

Shen:1988:HDC

[She88] Kang Sheng Shen. Historical development of the Chinese remainder theorem. Archive for History of Exact Sciences, 38(4):285–305, September 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=38&issue=4&spage=285.

Sheynin:1989:MWP

[She89a] O. B. Sheynin. A. A. Markov’s work on probability. Archive for History of Exact Sciences, 39(4):337–377, December 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
39&issue=4&spage=337. See erratum [She89b].

Sheynin:1989:ECM

[She89b] O. B. Sheynin. Errata for the contribution “A. A. Markov’s work on probability”. Archive for History of Exact Sciences, 40(4):387, December 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=40&issue=4&spage=387. See [She89a].

Sheynin:1991:HPW

[She91a] O. B. Sheynin. H. Poincaré’s work on probability. Archive for History of Exact Sciences, 42(2):137–171, June 1991. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=42&issue=2&spage=137.

Sheynin:1991:VYB

[She91b] O. B. Sheynin. On V. Ya. Buniakovsky’s work in the theory of probability. Archive for History of Exact Sciences, 43(3):199–223, September 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=43&issue=3&spage=199.

Sheynin:1993:HPL

[She93a] Oscar Sheynin. On the history of the principle of least squares. Archive for History of Exact Sciences, 46(1):39–54, March 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=1&spage=39.

Sheynin:1993:TOE

[She93b] Oscar Sheynin. The treatment of observations in early astronomy. Archive for History of Exact Sciences, 46(2):153– 192, June 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=2&spage=153.

Sheynin:1994:BWP

[She94a] Oscar Sheynin. Bertrand’s work on probability. Archive for History of Exact Sciences, 48(2):155–199, June 1994. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=48&issue=2&spage=155.

Sheynin:1994:CFG

[She94b] Oscar Sheynin. C. F. Gauss and geodetic observations. Archive for History of Exact Sciences, 46(3):253–283, February 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=3&spage=253.

Sheynin:1994:CLT

[She94c] Oscar Sheynin. Chebyshev’s lectures on the theory of probability. Archive for History of Exact Sciences, 46(4):321–340, December 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=4&spage=321.

Sheynin:1995:DCT

[She95a] Oscar Sheynin. Density curves in the theory of Errors. Archive for History of Exact Sciences, 49(2):163–196, June 1995. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=49&issue=2&spage=163.

Sheynin:1995:HWT

[She95b] Oscar Sheynin. Helmert’s work in the theory of errors. Archive for History of Exact Sciences, 49(1):73–104, March 1995. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=49&issue=1&spage=73.

Sheynin:1998:TPD

[She98] Oscar Sheynin. The Theory of Probability: Its definition and its relation to statistics. Archive for History of Exact Sciences, 52(2):99–108, February 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=52&issue=2&spage=99.

Sheynin:2003:NWP

[She03] Oscar B. Sheynin. Nekrasov’s work on probability: The background. Archive for History of Exact Sciences, 57(4):337– 353, May 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=4&spage=337.

Shi:2003:KAC

[Shi03] Yunli Shi. The Korean adaptation of the Chinese–Islamic astronomical tables. Archive for History of Exact Sciences, 57(1):25– 60, January 2003. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
57&issue=1&spage=25.

Shkolenok:1972:GCE

[Shk72] Galina A. Shkolenok. Geometrical constructions equivalent to non-linear algebraic transformations of the plane in Newton’s early papers. Archive for History of Exact Sciences, 9(1):22–44, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
9&issue=1&spage=22.

Shnider:2017:BTC

[Shn17] Steven Shnider. Britton’s theory of the creation of Column Φ in Babylonian System A lunar theory. Archive for History of Exact Sciences, 71(3):279–318, May 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Sidoli:2018:CGG

[Sid18a] Nathan Sidoli. The concept of given in Greek mathematics. Archive for History of Exact Sciences, 72(4):353–402, July 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Sidoli:2018:UCP

[Sid18b] Nathan Sidoli. Uses of construction in problems and theorems in Euclid’s Elements I–VI. Archive for History of Exact Sciences, 72(4):403–452, July 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Silvestri:1979:SGF

[Sil79] Richard Silvestri. Simple groups of finite order in the nineteenth century. Archive for History of Exact Sciences, 20 (3–4):313–356, September 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=20&issue=3&spage=313.

Steele:2021:SBR

[SM21] J. M. Steele and E. L. Meszaros. A study of Babylonian records of planetary stations. Archive for History of Exact Sciences, 75 (4):415–438, July 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-021-00272-5.

Smadja:2011:MIP

[Sma11] Ivahn Smadja. Des méthodes d’intégration par arcs de sections coniques aux échelles de modules. Legendre lecteur de Landen. (French) [Methods of integration by arcs of conic sections for module scales. Legendre teacher of Landen]. Archive for History of Exact Sciences, 65(4):343–395, July 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=????&volume=65&issue=4&spage=343.

Smeur:1970:VEA

[Sme70] A. J. E. M. Smeur. On the value equivalent to π in ancient mathematical texts. A new interpretation. Archive for History of Exact Sciences, 6(4):249–270, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=4&spage=249.

Smith:1977:WTC

[Smi77] Crosbie W. Smith. William Thomson and the creation of thermodynamics: 1840–1855. Archive for History of Exact Sciences, 16(3):231–288, September 1977. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=3&spage=231.

Smith:1978:MOA

[Smi78] Robin Smith. The mathematical origins of Aristotle’s syllogistic. Archive for History of Exact Sciences, 19(3):201–209, September 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=3&spage=201.

Smith:1981:SAA

[Smi81] A. Mark Smith. Saving the appearances of the appearances: The foundations of classical geometrical optics. Archive for History of Exact Sciences, 24(2):73–99, June 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=24&issue=2&spage=73.

Smith:1982:PSL

[Smi82] A. Mark Smith. Ptolemy’s search for a Law of Refraction: a case-study in the classical methodology of “saving the appearances” and its limitations. Archive for History of Exact Sciences, 26(3):221–240, September 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=26&issue=3&spage=221.

Smithies:1986:CCR

[Smi86] F. Smithies. Cauchy’s conception of rigour in analysis. Archive for History of Exact Sciences, 36(1):41–61, March 1986. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=36&issue=1&spage=41.

Spalt:1991:BLM

[Spa91a] Detlef D. Spalt. Bolzanos Lehre von den meßbaren Zahlen 1830–1989. (German) [Bolzano’s theory of measurable numbers 1830–1989]. Archive for History of Exact Sciences, 42(1):15– 70, March 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
42&issue=1&spage=15.

Spalt:1991:MPG

[Spa91b] Detlef D. Spalt. Die mathematischen und philosophischen Grundlagen des Weierstraßschen Zahlbegriffs zwischen Bolzano und Cantor. (German) [The mathematical and philosophical basis of the Weierstrass number concept between Bolzano and Cantor]. Archive for History of Exact Sciences, 41(4):311–362, December 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=4&spage=311.

Spalt:2002:CKH

[Spa02] Detlef D. Spalt. Cauchys Kontinuum: Eine historiografische Annäherung via Cauchys Summensatz. (German) [Cauchy’s continuum: a historigraphical approach via cauchy’s sum theorem]. Archive for History of Exact Sciences, 56(4):285–338, June 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=4&spage=285.

Spencer:1967:BTR

[Spe67] J. Brookes Spencer. Boscovich’s theory and its relation to Faraday’s researches: an analytic approach. Archive for History of Exact Sciences, 4(3):184–202, January 1967. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=4&issue=3&spage=184.

Speiser:1996:KPN

[Spe96] David Speiser. The Kepler Problem from Newton to Johann Bernoulli. Archive for History of Exact Sciences, 50(2):103– 116, August 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
50&issue=2&spage=103.

Seneta:2001:NCD

[SPJ01] Eugene Seneta, Karen Hunger Parshall, and François Jongmans. Nineteenth-Century developments in geometric probability: J. J. Sylvester, M. W. Crofton, J.-É. Barbier, and J. Bertrand. Archive for History of Exact Sciences, 55(6):501– 524, June 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=6&spage=501.

Siegmund-Schultze:1982:AFI

[SS82] Reinhard Siegmund-Schultze. Die Anfänge der Funktionalanalysis und ihr Platz im Umwälzungsprozeß der Mathematik um 1900. (German) [The beginnings of functional analysis and its place during the circulation process of mathematics around 1900]. Archive for History of Exact Sciences, 26(1):13–71, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
26&issue=1&spage=13.

Siegmund-Schultze:1986:BHS

[SS86] R. Siegmund-Schultze. Der Beweis des Hilbert–Schmidt-Theorems. (German) [The proof of the Hilbert–Schmidt Theorem]. Archive for History of Exact Sciences, 36(3):251–270, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
36&issue=3&spage=251.

Siegmund-Schultze:1998:EHM

[SS98] Reinhard Siegmund-Schultze. Eliakim Hastings Moore’s “General Analysis”. Archive for History of Exact Sciences, 52(1):51– 89, January 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
52&issue=1&spage=51.

Scarpello:2005:WTH

[SS05] Giovanni Mingari Scarpello and Aldo Scimone. The Work of Tschirnhaus, La Hire and Leibniz on catacaustics and the birth of the envelopes of lines in the 17th century. Archive for History of Exact Sciences, 59(3):223–250, March 2005. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=59&issue=3&spage=223.

Siegmund-Schultze:2006:PMP

[SS06] Reinhard Siegmund-Schultze. Probability in 1919/20: the von Mises–Pólya-controversy. Archive for History of Exact Sciences, 60(5):431–515, September 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=60&issue=5&spage=431.

Sidoli:2009:RGC

[SS09] Nathan Sidoli and Ken Saito. The role of geometrical construction in Theodosius’s Spherics. Archive for History of Exact Sciences, 63(6):581–609, November 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=6&spage=581.

Sauer:2021:EIP

[SS21] Tilman Sauer and Tobias Schütz. Einstein on involutions in projective geometry. Archive for History of Exact Sciences, 75(5): 523–555, September 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-020-00270-z.

Schroder:1994:ELV

[ST94] Wilfried Schröder and Hans-Jürgen Treder. Zu Einsteins letzter Vorlesung — Beobachtbarkeit, Realität und Vollständigkeit in Quanten- und Relativitätstheorie. (German) [On Einstein’s last lecture — observability, reality, and completeness in quantum and Relativity theory]. Archive for History of Exact Sciences, 48(2):149–154, June 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=48&issue=2&spage=149.

Stedall:2000:RGP

[Ste00a] Jacqueline A. Stedall. Robd́ of glories: The posthumous misfortunes of Thomas Harriot and his algebra. Archive for History of Exact Sciences, 54(6):455–497, June 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=54&issue=6&spage=455.

Steele:2000:UAT

[Ste00b] John M. Steele. A 3405: An unusual astronomical text from Uruk. Archive for History of Exact Sciences, 55(2):103–135, December 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=2&spage=103.

Steele:2000:EPM

[Ste00c] John M. Steele. Eclipse Prediction in Mesopotamia. Archive for History of Exact Sciences, 54(5):421–454, February 2000. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=54&issue=5&spage=421.

Stedall:2001:DWR

[Ste01] Jacqueline A. Stedall. The discovery of wonders: Reading between the lines of John Wallis’s Arithmetica infinitorum. Archive for History of Exact Sciences, 56(1):1–28, November 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=1&spage=1.

Steele:2006:MLT

[Ste06] J. M. Steele. Miscellaneous lunar tables from Babylon. Archive for History of Exact Sciences, 60(2):123–155, March 2006. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=60&issue=2&spage=123.

Stedall:2008:NMT

[Ste08] Jacqueline Stedall. Notes made by Thomas Harriot on the treatises of François Viète. Archive for History of Exact Sciences, 62(2):179–200, March 2008. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=2&spage=179.

Steele:2020:PNS

[Ste20] John Steele. A proto-Normal Star Almanac dating to the reign of Artaxerxes III: BM 65156. Archive for History of Exact Sciences, 74(3):243–253, May 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-019-00246-8.

Steele:2022:AIB

[Ste22] John Steele. An alternative interpretation of BM 76829: astrological schemes for length of life and parts of the body. Archive for History of Exact Sciences, 76(1):1–14, January 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-021-00279-y.

Steele:2024:NMI

[Ste24a] John Steele. The new moon interval NA and the beginning of the Babylonian month. Archive for History of Exact Sciences, 78(3):245–270, May 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-023-00325-x.

Stewart:2024:GSG

[Ste24b] Ian Stewart. Galois and the simple group of order 60. Archive for History of Exact Sciences, 78(1):1–28, January 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00319-9.

Stromholm:1968:FMM

[Str68] Per Strømholm. Fermat’s methods of maxima and minima and of tangents. A reconstruction. Archive for History of Exact Sciences, 5(1):47–69, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=5&issue=1&spage=47.

Straker:1981:KTO

[Str81] Stephen Straker. Kepler, Tycho, and the ‘optical part of astronomy’: the genesis of Kepler’s theory of pinhole images. Archive for History of Exact Sciences, 24(4):267–293, December 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
24&issue=4&spage=267.

Straumann:1992:SJT

[Str92] Norbert Straumann. Von der Stech–Jensen-Transformation zur universellen V-A-Wechselwirkung. (German) [On the Stech– Jensen transformation to the universal V-A-reciprocal effect]. Archive for History of Exact Sciences, 44(4):365–386, December 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
44&issue=4&spage=365.

Saccomandi:2024:ASP

[SV24] Giuseppe Saccomandi and Maurizio Stefano Vianello. Antonio Signorini and the proto-history of the non-linear theory of elasticity. Archive for History of Exact Sciences, 78(4):375–400, July 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-024-00328-2.

Swerdlow:1980:HDL

[Swe80] Noel M. Swerdlow. Hipparchus’s determination of the length of the tropical year and the rate of precession. Archive for History of Exact Sciences, 21(4):291–309, December 1980. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=21&issue=4&spage=291.

Swerdlow:1999:ARB

[Swe99] N. M. Swerdlow. Acronychal risings in Babylonian planetary theory. Archive for History of Exact Sciences, 54(1):49– 65, April 1999. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=1&spage=49.

Sylla:1971:MQQ

[Syl71] Edith Sylla. Medieval quantifications of qualities: The “Merton School”. Archive for History of Exact Sciences, 8(1–2):9–39, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
8&issue=1&spage=9.

Szabo:1960:AEA

[Sza60] Árpád Szabó. Anfänge des euklidischen Axiomensystems. (German) [Beginnings of the Euclidean axiom system]. Archive for History of Exact Sciences, 1(1):37–106, August 1960. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=1&issue=1&spage=

Szabo:1964:FPI

[Sza64] Árpád Szabó. Die frühgriechische Proportionenlehre im Spiegel ihrer Terminologie. (German) [The early Greek proportion teachings in the mirror of their terminology]. Archive for History of Exact Sciences, 2(3):197–270, December 1964. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=3&spage=
197.

Takacs:1980:PC

[Tak80] Lajos Takács. The problem of coincidences. Archive for History of Exact Sciences, 21(3):229–244, September 1980. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=21&issue=3&spage=229.

Tanacs:2009:GCD

[Tan09] János Tanács. Grasping the conceptual difference between János Bolyai and Lobachevskii’s notions of non-Euclidean parallelism. Archive for History of Exact Sciences, 63(5):537–552, 2009. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Taylor:1974:DNT

[Tay74] Angus E. Taylor. The differential: Nineteenth and twentieth century developments. Archive for History of Exact Sciences, 12(4):355–383, August 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=12&issue=4&spage=355.

Taylor:1982:SMF

[Tay82] Angus E. Taylor. A study of Maurice Fréchet: I. His early work on point set theory and the theory of functionals. Archive for History of Exact Sciences, 27(3):233–295, September 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=3&spage=233.

Taylor:1985:SMF

[Tay85] Angus E. Taylor. A study of Maurice Fréchet: II. Mainly about his work on general topology, 1909–1928. Archive for History of Exact Sciences, 34(4):279–380, December 1985. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=34&issue=4&spage=279.

Taylor:1987:SMF

[Tay87] Angus E. Taylor. A study of Maurice Fréchet: III. Fréchet as analyst, 1909–1930. Archive for History of Exact Sciences, 37(1): 25–76, March 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
37&issue=1&spage=25.

Tazzioli:1993:ETE

[Taz93] Rossana Tazzioli. Ether and theory of elasticity in Beltrami’s work. Archive for History of Exact Sciences, 46(1):1–37, March 1993. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=1&spage=1.

Thorup:1992:PET

[Tho92] Anders Thorup. A pre-Euclidean theory of proportions. Archive for History of Exact Sciences, 45(1):1–16, March 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=45&issue=1&spage=1.

Thomaidis:2005:EFD

[Tho05a] Yannis Thomaidis. Erratum: a Framework for Defining the Generality of Diophantos’ Methods in “arithmetica”. Archive for History of Exact Sciences, 59(6):641, October 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
59&issue=6&spage=641. See [Tho05b].

Thomaidis:2005:FDG

[Tho05b] Yannis Thomaidis. A framework for defining the generality of Diophantos’ methods in “Arithmetica”. Archive for History of Exact Sciences, 59(6):591–640, October 2005. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=59&issue=6&spage=591. See

erratum [Tho05a].

Thro:1994:LVS

[Thr94] E. Broydrick Thro. Leonardo da Vinci’s solution to the problem of the pinhole camera. Archive for History of Exact Sciences, 48(3–4):343–371, September 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=48&issue=3&spage=343.

Thurston:1992:GIP

[Thu92] Hugh Thurston. Greek and Indian planetary longitudes. Archive for History of Exact Sciences, 44(3):191–195, September 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
44&issue=3&spage=191.

Toepell:1986:ODH

[Toe86] Michael Toepell. On the origins of David Hilbert’s “Grundlagen der Geometrie” [Foundations of Geometry]. Archive for History of Exact Sciences, 35(4):329–344, December 1986. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=35&issue=4&spage=329.

Toomer:1974:HDS

[Too74] G. J. Toomer. Hipparchus on the distances of the sun and moon. Archive for History of Exact Sciences, 14(2):126–142, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
14&issue=2&spage=126.

Toomer:1985:GAA

[Too85] G. J. Toomer. Galen on the astronomers and astrologers. Archive for History of Exact Sciences, 32(3–4):193–206, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=3&spage=193.

Topper:1971:CMJ

[Top71] David R. Topper. Commitment to mechanism: J. J. Thomson, the early years. Archive for History of Exact Sciences, 7(5):393– 410, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
7&issue=5&spage=393.

Toth:1966:PIC

[Tót66] Imre Tóth. Das Parallelenproblem im Corpus Aristotelicum. (German) [The problem of parallels in Corpus Aristotelicum]. Archive for History of Exact Sciences, 3(4–5):249–422, January 1966. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
3&issue=4&spage=249.

Truesdell:1960:PTR

[Tru60] C. Truesdell. A program toward rediscovering the rational mechanics of the age of reason. Archive for History of Exact Sciences, 1(1):1–36, August 1960. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=1&spage=1.

Truesdell:1975:EKT

[Tru75] C. Truesdell. Early kinetic theories of gases. Archive for History of Exact Sciences, 15(1):1–66, March 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=15&issue=1&spage=1.

Truesdell:1979:ATC

[Tru79] C. Truesdell. Absolute temperatures as a consequence of Carnot’s general axiom. Archive for History of Exact Sciences, 20(3–4):357–380, September 1979. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=20&issue=3&spage=357.

Truesdell:1989:MGA

[Tru89] C. Truesdell. Maria Gaetana Agnesi. Archive for History of Exact Sciences, 40(2):113–142, June 1989. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=40&issue=2&spage=113. See corrections

and additions [Tru92].

Truesdell:1992:CAS

[Tru92] C. Truesdell. Corrections and additions for “Maria Gaetana Agnesi”. Archive for History of Exact Sciences, 43(4):385–386, December 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
43&issue=4&spage=385. See [Tru89].

Truesdell:1996:TFA

[Tru96] C. Truesdell. The thirty-fifth anniversary of this archive, by the editor. Archive for History of Exact Sciences, 50(1):1– 4, March 1996. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
50&issue=1&spage=1.

Tucciarone:1973:DTS

[Tuc73] John Tucciarone. The development of the theory of summable divergent series from 1880 to 1925. Archive for History of Exact Sciences, 10(1–2):1–40, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=10&issue=1&spage=1.

Tuman:1992:ADT

[Tum92] V. S. Tuman. An attempt to date text 3 of Enuma Anu Enlil, tablets 50–51 “tentative date December 2, 1878”. Archive for History of Exact Sciences, 45(2):95–103, June 1992. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=45&issue=2&spage=95.

Tupikova:2022:CSA

[Tup22] Irina Tupikova. A common-sense approach to the problem of the itinerary stadion. Archive for History of Exact Sciences, 76 (4):319–361, July 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00287-6.

Tweddle:1991:JMR

[Twe91a] Ian Tweddle. John Machin and Robert Simson on inversetangent series for π. Archive for History of Exact Sciences, 42(1): 1–14, March 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
42&issue=1&spage=1.

Tweddle:1991:SRC

[Twe91b] Ian Tweddle. Some results on conic sections in the correspondence between Colin MacLaurin and Robert Simson. Archive for History of Exact Sciences, 41(4):285–309, December 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=4&spage=285.

Tweddle:1992:JSE

[Twe92] Ian Tweddle. James Stirling’s early work on acceleration of convergence. Archive for History of Exact Sciences, 45(2):105– 125, June 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
45&issue=2&spage=105.

Ulivi:2015:MQC

[Uli15] Elisabetta Ulivi. Masters, questions and challenges in the abacus schools. Archive for History of Exact Sciences, 69(6):651– 670, November 2015. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-015-0160-1.

Ullmann:1984:CEE

[Ull84] Dieter Ullmann. Chladni und die Entwicklung der experimentellen Akustik um 1800. (German) [Chladni and the development of experimental acoustics in 1800]. Archive for History of Exact Sciences, 31(1):35–52, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=31&issue=1&spage=35.

Ullrich:1989:WVS

[Ull89] Peter Ullrich. Weierstraß’ Vorlesung zur “Einleitung in die Theorie der analytischen Funktionen”. (German) [Weierstrass’s lectures on “Introduction to the theory of analytical functions]. Archive for History of Exact Sciences, 40(2):143–172, June 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
40&issue=2&spage=143.

Ullrich:2000:PVT

[Ull00] Peter Ullrich. The Poincaré–Volterra Theorem: From hyperelliptic integrals to manifolds with countable topology. Archive for History of Exact Sciences, 54(5):375–402, February 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
54&issue=5&spage=375.

Unguru:1974:PTC

[Ung74] Sabetai Unguru. Pappus in the thirteenth century in the Latin West. Archive for History of Exact Sciences, 13(4):307–324, December 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
13&issue=4&spage=307.

Unguru:1975:NRH

[Ung75] Sabetai Unguru. On the need to rewrite the history of Greek mathematics. Archive for History of Exact Sciences, 15(1):67– 114, March 1975. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
15&issue=1&spage=67.

VandenEynde:1992:HEC

[Van92] Ria Vanden Eynde. Historical evolution of the concept of homotopic paths. Archive for History of Exact Sciences, 45(2):127– 188, June 1992. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
45&issue=2&spage=127.

VanBrummelen:2006:TLP

[Van06] Glen Van Brummelen. Taking latitude with Ptolemy: Jamshīd al-Kāshī’s novel geometric model of the motions of the inferior planets. Archive for History of Exact Sciences, 60(4):353– 377, July 2006. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
60&issue=4&spage=353.

VanBrummelen:2018:EEB

[Van18] Glen Van Brummelen. The end of an error: Bianchini, Regiomontanus, and the tabulation of stellar coordinates. Archive for History of Exact Sciences, 72(5):547–563, September 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

VanBrummelen:2021:BEE

[Van21] Glen Van Brummelen. Before the end of an error: Giovanni Bianchini’s original flawed treatise on the conversion of stellar coordinates. Archive for History of Exact Sciences, 75(1): 109–124, January 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00407-020-00261-0.pdf.

vanBrakel:1976:SRP

[vB76] J. van Brakel. Some remarks on the prehistory of the concept of statistical probability. Archive for History of Exact Sciences, 16(2):119–136, June 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=2&spage=119.

vanBesouw:2017:WVV

[vB17] Jip van Besouw. The wedge and the vis viva controversy: how concepts of force influenced the practice of early eighteenthcentury mechanics. Archive for History of Exact Sciences, 71 (2):109–156, March 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-016-0182-3.

vanDongen:2004:EMS

[vD04] Jeroen van Dongen. Einstein’s methodology, semivectors and the unification of electrons and protons. Archive for History of Exact Sciences, 58(3):219–254, March 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=58&issue=3&spage=219.

vanDongen:2009:RMM

[vD09] Jeroen van Dongen. On the role of the Michelson–Morley Experiment: Einstein in Chicago. Archive for History of Exact Sciences, 63(6):655–663, November 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=63&issue=6&spage=655.

vanderWaerden:1960:AMP

[vdW60] B. L. van der Waerden. Ausgleichspunkt, “Methode der Perser” und indische Planetenrechnung. (German) [Balance point, “method of the Persian” and Indian planet calculation]. Archive for History of Exact Sciences, 1(2):107–121, December 1960. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=1&issue=2&spage=
107.

vanderWaerden:1968:DIB

[vdW68] B. L. van der Waerden. The date of invention of Babylonian planetary theory. Archive for History of Exact Sciences, 5(1):70–78, January 1968. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
5&issue=1&spage=70.

vanderWaerden:1971:FAG

[vdW71] B. L. van der Waerden. The foundation of algebraic geometry from Severi to André Weil. Archive for History of Exact Sciences, 7(3):171–180, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=7&issue=3&spage=171.

vanderWaerden:1972:GTH

[vdW72a] B. L. van der Waerden. Die Galois–Theorie von Heinrich Weber bis Emil Artin. (German) [Galois theory from Heinrich Weber to Emil Artin]. Archive for History of Exact Sciences, 9(3):240–248, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=9&issue=3&spage=240.

vanderWaerden:1972:BZP

[vdW72b] B. L. van der Waerden. Ein Briefwechsel zwischen P. Finsler und B. L. van der Waerden über die Grundlegung der algebraischen Geometrie. (German) [Letters between P. Finsler and B. L. van der Waerden on the foundation of algebraic geometry]. Archive for History of Exact Sciences, 9(3):249–256, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=9&issue=3&spage=
249.

vanderWaerden:1976:DSP

[vdW76] B. L. van der Waerden. Defence of a “shocking” point of view. Archive for History of Exact Sciences, 15(3):199–210, September 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
15&issue=3&spage=199.

vanderWaerden:1978:PKF

[vdW78a] B. L. van der Waerden. Die Postulate und Konstruktionen in der frühgriechischen Geometrie. (German) [Postulates and constructions in early Greek geometry]. Archive for History of Exact Sciences, 18(4):343–357, December 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=18&issue=4&spage=343.

vanderWaerden:1978:GYG

[vdW78b] B. L. van der Waerden. The Great Year in Greek, Persian and Hindu astronomy. Archive for History of Exact Sciences, 18(4):359–383, December 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=18&issue=4&spage=359.

vanderWaerden:1980:PBMa

[vdW80a] B. L. van der Waerden. On pre-Babylonian mathematics I. Archive for History of Exact Sciences, 23(1):1–25, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=1&spage=1.

vanderWaerden:1980:PBMb

[vdW80b] B. L. van der Waerden. On pre-Babylonian mathematics II. Archive for History of Exact Sciences, 23(1):27–46, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
23&issue=1&spage=27.

vanderWaerden:1982:MVM

[vdW82] B. L. van der Waerden. The motion of Venus, Mercury and the Sun in early Greek astronomy. Archive for History of Exact Sciences, 26(2):99–113, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=26&issue=2&spage=99.

vanderWaerden:1984:GACa

[vdW84a] B. L. van der Waerden. Greek astronomical calendars I. The parapegma of Euctemon. Archive for History of Exact Sciences, 29(2):101–114, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=29&issue=2&spage=101.

vanderWaerden:1984:GACb

[vdW84b] B. L. van der Waerden. Greek astronomical calendars II. Callippos and his calendar. Archive for History of Exact Sciences, 29(2):115–124, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=29&issue=2&spage=115.

vanderWaerden:1984:GACc

[vdW84c] B. L. van der Waerden. Greek astronomical calendars III. The calendar of Dionysios. Archive for History of Exact Sciences, 29(2):125–130, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=29&issue=2&spage=125.

vanderWaerden:1985:GACa

[vdW85a] B. L. van der Waerden. Greek astronomical calendars IV. The parapegma of the Egyptians and their “perpetual tables”. Archive for History of Exact Sciences, 32(2):95–104, June 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
32&issue=2&spage=95.

vanderWaerden:1985:GACb

[vdW85b] B. L. van der Waerden. Greek astronomical calendars V. The motion of the Sun in the parapegma of Geminos and in the Romaka–Siddhānta. Archive for History of Exact Sciences, 34(3):231–239, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=34&issue=3&spage=231.

vanderWaerden:1988:DBW

[vdW88a] B. L. van der Waerden. The “day of Brahman” in the work of Āryabhata. Archive for History of Exact Sciences, 38(1):13– 22, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
38&issue=1&spage=13.

vanderWaerden:1988:RS

[vdW88b] B. L. van der Waerden. On the Romaka–Siddhānta. Archive for History of Exact Sciences, 38(1):1–11, March 1988. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=38&issue=1&spage=1.

vanderWaerden:1988:RGT

[vdW88c] B. L. van der Waerden. Reconstruction of a Greek table of chords. Archive for History of Exact Sciences, 38(1):23–38, March 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
38&issue=1&spage=23.

Verdun:2013:LEEa

[Ver13a] Andreas Verdun. Leonhard Euler’s early lunar theories 1725– 1752. Archive for History of Exact Sciences, 67(3):235–303, May 2013. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-012-0112-y.

Verdun:2013:LEEb

[Ver13b] Andreas Verdun. Leonhard Euler’s early lunar theories 1725– 1752. Archive for History of Exact Sciences, 67(5):477–551, September 2013. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL http://link.springer.com/
article/10.1007/s00407-013-0118-0.

Vergnerie:2020:CEP

[Ver20] Cédric Vergnerie. Une controverse entre Émile Picard et Leopold Kronecker. (French) [A controversy between Émile Picard and Leopold Kronecker]. Archive for History of Exact Sciences, 74 (2):131–164, March 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-019-00241-z.

Viertel:2021:DCU

[Vie21] Klaus Viertel. The development of the concept of uniform convergence in Karl Weierstrass’s lectures and publications between 1861 and 1886. Archive for History of Exact Sciences, 75(4):455–490, July 2021. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-020-00266-9.

Vignolles:2000:DEA

[Vig00] Hélène Vignolles. La distance des étoiles au dix-huitième siècle: L’échelle des magnitudes de John Michell. (French) [The distance of stars in the 18^{th} century: John Mitchell’s magnitude scale]. Archive for History of Exact Sciences, 55(1):77–101, August 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
55&issue=1&spage=77.

Vinel:2005:CMP

[Vin05] Nicolas Vinel. Un carré magique pythagoricien ? Jamblique précurseur des témoins Arabo–Byzantins. (French) [A Pythagorean magic square? Jamblique precursor of the arab– byzantine witnesses]. Archive for History of Exact Sciences, 59(6):545–562, October 2005. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=6&spage=545.

Vita:1982:PNT

[Vit82] Vincenzo Vita. Il punto nella terminologia matematica greca. (Italian) [The point in Greek mathematical terminology]. Archive for History of Exact Sciences, 27(2):101–114, June 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
27&issue=2&spage=101.

Vitrac:2002:NTP

[Vit02] Bernard Vitrac. Note textuelle sur un (problème de) lieu Géométrique dans les Météorologiques d’aristote (III. 5, 375 b 16–376 b 22). (French) [Textual note on the (problem of) geometric place in the meteorologies of Aristotle (III. 5, 375 b 16–376 b 22)]. Archive for History of Exact Sciences, 56(3):239– 283, May 2002. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=3&spage=239.

Vitrac:2004:PDA

[Vit04] Bernard Vitrac. À propos des Démonstrations alternatives et autres substitutions de preuves dans les Éléments dÉuclide. (French) [On alternative demonstrations and other substitutions of proofs in Euclid’s Elements]. Archive for History of Exact Sciences, 59(1):1–44, November 2004. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=59&issue=1&spage=1.

Voelke:2008:TFG

[Voe08] Jean-Daniel Voelke. Le théorème fondamental de la géométrie projective: évolution de sa preuve entre 1847 et 1900. (French) [The Fundamental Theorem of Projective Geometry: evolution of its proof between 1847 and 1900]. Archive for History of Exact Sciences, 62(3):243–296, May 2008. CODEN AHESAN.

ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=62&issue=3&spage=243.

Volkert:1987:GPF

[Vol87] Klaus Volkert. Die Geschichte der pathologischen Funktionen — Ein Beitrag zur Entstehung der mathematischen Methodologie. (German) [The history of pathological functions — a contribution for the emergence of the mathematical methodology]. Archive for History of Exact Sciences, 37(3):193–232, September 1987. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
37&issue=3&spage=193.

Volkert:1989:DSF

[Vol89] Klaus Volkert. Zur Differenzierbarkeit stetiger Funktionen — Ampère’s Beweis und seine Folgen. (German) [On the differentiability of constant functions — Ampère’s proof and its consequences]. Archive for History of Exact Sciences, 40(1):37– 112, March 1989. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
40&issue=1&spage=37.

vonPlato:1991:BEH

[vP91] Jan von Plato. Boltzmann’s ergodic hypothesis. Archive for History of Exact Sciences, 42(1):71–89, March 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=42&issue=1&spage=71.

vanRootselaar:1963:BTR

[vR63] B. van Rootselaar. Bolzano’s theory of real numbers. Archive for History of Exact Sciences, 2(2):168–180, August 1963. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=2&spage=
168.

Vidal:2018:CMD

[VT18a] J. Climent Vidal and J. Soliveres Tur. Correction to: The modernity of Dedekind’s anticipations contained in What are numbers and what are they good for? Archive for History of Exact Sciences, 72(2):143, March 2018. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s00407-018-0203-5.pdf. See [VT18b].

Vidal:2018:MDA

[VT18b] J. Climent Vidal and J. Soliveres Tur. The modernity of Dedekind’s anticipations contained in What are numbers and what are they good for? Archive for History of Exact Sciences, 72(2):99–141, March 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). See correction [VT18a].

Wagner:2010:NNA

[Wag10] Roy Wagner. The natures of numbers in and around Bombelli’s l’algebra. Archive for History of Exact Sciences, 64(5):485–523, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Warwick:1991:RFL

[War91] Andrew Warwick. On the role of the FitzGerald–Lorentz contraction hypothesis in the development of Joseph Larmor’s electronic theory of matter. Archive for History of Exact Sciences, 43(1):29–91, March 1991. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=43&issue=1&spage=29.

Waschkies:1971:NHE

[Was71] H.-J. Waschkies. Eine neue Hypothese zur Entdeckung der inkommensurablen Größen durch die Griechen. (German) [A new hypothesis on the discovery of incommensurable sizes by the Greeks]. Archive for History of Exact Sciences, 7(5):325– 353, January 1971. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
7&issue=5&spage=325.

Waterhouse:1972:DRS

[Wat72] William C. Waterhouse. The discovery of the regular solids. Archive for History of Exact Sciences, 9(3):212–221, January 1972. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
9&issue=3&spage=212.

Waterhouse:1978:NME

[Wat78] William C. Waterhouse. Note on a method of extracting roots in as-Samaw’al. Archive for History of Exact Sciences, 19(4):383–384, December 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=19&issue=4&spage=383.

Waterhouse:1980:EPS

[Wat80] William C. Waterhouse. The early proofs of Sylow’s theorem. Archive for History of Exact Sciences, 21(3):279–290, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
21&issue=3&spage=279.

Waterhouse:1990:GFA

[Wat90] William C. Waterhouse. Gauss’s first argument for least squares. Archive for History of Exact Sciences, 41(1):41–52, March 1990. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
41&issue=1&spage=41.

Webster:1965:DBL

[Web65] C. Webster. The discovery of Boyle’s Law, and the concept of the elasticity of air in the seventeenth century. Archive for History of Exact Sciences, 2(6):441–502, December 1965. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=2&issue=6&spage=
441.

Weil:1978:WBE

[Wei78] André Weil. Who betrayed Euclid? (extract from a letter to the editor). Archive for History of Exact Sciences, 19(2):91– 93, June 1978. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
19&issue=2&spage=91.

Weil:1979:RBB

[Wei79] André Weil. Riemann, Betti and the birth of topology. Archive for History of Exact Sciences, 20(2):91–96, June 1979. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=20&issue=2&spage=91. See

postscript [Wei80].

Weil:1980:PMA

[Wei80] André Weil. A postscript to my article “Riemann, Betti and the birth of topology”. Archive for History of Exact Sciences, 21(4):387, December 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=21&issue=4&spage=387. See [Wei79].

Weinstock:2000:ISO

[Wei00] Robert Weinstock. Inverse–square orbits in Newton’s Principia and twentieth-century commentary thereon. Archive for History of Exact Sciences, 55(2):137–162, December 2000. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=55&issue=2&spage=137.

Westfall:1964:INC

[Wes64] Richard S. Westfall. Isaac Newton’s coloured circles twixt two contiguous glasses. Archive for History of Exact Sciences, 2(3):181–196, December 1964. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=2&issue=3&spage=181.

Westfall:1986:NAG

[Wes86] Richard S. Westfall. Newton and the acceleration of gravity. Archive for History of Exact Sciences, 35(3):255–272, September 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
35&issue=3&spage=255.

Wilson:2014:CHM

[WH14] Curtis Wilson and William Harper. The coming-to-be of Hansen’s method. Archive for History of Exact Sciences, 68 (4):409–497, July 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-013-0128-y.

Wagner:2019:JBM

[WH19] Roy Wagner and Samuel Hunziker. Jost Bürgi’s methods of calculating sines, and possible transmission from India. Archive for History of Exact Sciences, 73(3):243–260, May 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Whiteside:1961:PMT

[Whi61] Derek Thomas Whiteside. Patterns of mathematical thought in the later seventeenth century. Archive for History of Exact Sciences, 1(3):179–388, March 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=1&issue=3&spage=179.

Wilson:1970:KLC

[Wil70] Curtis A. Wilson. From Kepler’s Laws, so-called, to universal gravitation: Empirical factors. Archive for History of Exact Sciences, 6(2):89–170, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=6&issue=2&spage=89.

Wilson:1980:PSTa

[Wil80a] Curtis A. Wilson. Perturbations and solar tables from Lacaille to Delambre: the rapprochement of observation and theory, part I. Archive for History of Exact Sciences, 22(1–2):53– 188, March 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
22&issue=1&spage=53.

Wilson:1980:PSTb

[Wil80b] Curtis A. Wilson. Perturbations and solar tables from Lacaille to Delambre: the rapprochement of observation and theory, part II. Archive for History of Exact Sciences, 22(3):189–304, September 1980. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
22&issue=3&spage=189.

Wilder:1982:MWR

[Wil82] R. L. Wilder. The mathematical work of R. L. Moore: Its background, nature and influence. Archive for History of Exact Sciences, 26(1):73–97, March 1982. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=26&issue=1&spage=73.

Wilson:1985:GIJ

[Wil85] Curtis Wilson. The great inequality of Jupiter and Saturn: from Kepler to Laplace. Archive for History of Exact Sciences, 33(1–3):15–290, March 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=33&issue=1&spage=15.

Wilson:1987:DVE

[Wil87] Curtis Wilson. D’Alembert versus Euler on the precession of the equinoxes and the mechanics of rigid bodies. Archive for History of Exact Sciences, 37(3):233–273, September 1987. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=37&issue=3&spage=233.

Wimsatt:2012:AGG

[Wim12] W. C. Wimsatt. The analytic geometry of genetics: part I: the structure, function, and early evolution of Punnett squares. Archive for History of Exact Sciences, 66(4):359– 396, July 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
66&issue=4&spage=359.

Wisan:1974:NSM

[Wis74] Winifred L. Wisan. The new science of motion: a study of Galileo’s De motu locali. Archive for History of Exact Sciences, 13(2–3):103–306, June 1974. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=13&issue=2&spage=103.

Wise:1981:FAE

[Wis81] M. Norton Wise. The flow analogy to electricity and magnetism, part I: William Thomson’s reformulation of action at a distance. Archive for History of Exact Sciences, 25(1):19– 70, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
25&issue=1&spage=19.

Wagenmakers:2023:HNJ

[WL23] Eric-Jan Wagenmakers and Alexander Ly. History and nature of the Jeffreys–Lindley paradox. Archive for History of Exact Sciences, 77(1):25–72, January 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00298-3. See

comments [GCWL23].

Wright:2000:GRP

[Wri00] M. T. Wright. Greek and Roman portable sundials: An ancient essay in approximation. Archive for History of Exact Sciences, 55(2):177–187, December 2000. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=55&issue=2&spage=177.

Yavetz:1998:HSE

[Yav98] Ido Yavetz. On the homocentric spheres of Eudoxus. Archive for History of Exact Sciences, 52(3):221–278, February 1998. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=52&issue=3&spage=221.

Yavetz:2001:NRH

[Yav01] Ido Yavetz. A new role for the Hippopede of Eudoxus. Archive for History of Exact Sciences, 56(1):69–93, November 2001. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
56&issue=1&spage=69.

Yazdi:2011:AKA

[Yaz11a] Hamid-Reza Giahi Yazdi. Al-Khwarizmi and annular solar eclipse. Archive for History of Exact Sciences, 65(5):499–517, September 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
65&issue=5&spage=499. See erratum [Yaz11b].

Yazdi:2011:EAK

[Yaz11b] Hamid-Reza Giahi Yazdi. Erratum to: Al-Khwarizmi and annular solar eclipse. Archive for History of Exact Sciences, 65(5):589, September 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=65&issue=5&spage=589. See [Yaz11a].

Yazdi:2018:MTL

[Yaz18] Hamid-Reza Giahi Yazdi. The mysterious table of lunar crescent visibility attributed to Al-Bīrunī and h.abash al-h.asib’s contribution. Archive for History of Exact Sciences, 72(1):89–98, January 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Yazdi:2020:JCE

[Yaz20] Hamid-Reza Giahi Yazdi. The Jalali Calendar: the enigma of its radix date. Archive for History of Exact Sciences, 74(2): 165–182, March 2020. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-019-00240-0.

Yeang:2011:TRB

[Yea11] Chen-Pang Yeang. Tubes, randomness, and Brownian motions: or, how engineers learned to start worrying about electronic noise. Archive for History of Exact Sciences, 65(4):437–470, July 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=????&volume=65&issue=4&spage=
437.

Yiwen:2020:QJW

[Yiw20] Zhu Yiwen. On Qin Jiushao’s writing system. Archive for History of Exact Sciences, 74(4):345–379, July 2020. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s00407-019-00243-x.

Yong:1969:EFY

[Yon69] Lam Lay Yong. On the existing fragments of Yang Hui’s Hsiang Chieh Suan Fa. Archive for History of Exact Sciences, 6(1):82– 88, January 1969. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
6&issue=1&spage=82.

Youschkevitch:1976:CFM

[You76] A. P. Youschkevitch. The concept of function up to the middle of the 19^{th} century. Archive for History of Exact Sciences, 16(1):37–85, March 1976. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=16&issue=1&spage=37.

Yilmaz:2018:AOH

[YT18] Mustafa Yilmaz and Ibrahim Tiryakioğlu. The astronomical orientation of the historical Grand mosques in Anatolia (Turkey). Archive for History of Exact Sciences, 72(6):565–590, November 2018. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Zabell:1988:BPL

[Zab88] S. L. Zabell. Buffon, Price, and Laplace: Scientific attribution in the 18th century. Archive for History of Exact Sciences, 39(2):173–181, June 1988. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=39&issue=2&spage=173.

Zaitsev:1994:IPL

[Zai94] Evgeny A. Zaitsev. An interpretation of Peano’s logic. Archive for History of Exact Sciences, 46(4):367–383, December 1994. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
46&issue=4&spage=367.

Zambelli:2010:CKD

[Zam10] Stefano Zambelli. Chemical kinetics and diffusion approach: the history of the Klein–Kramers equation. Archive for History of Exact Sciences, 64(4):395–428, 2010. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Zangwill:2013:HTF

[Zan13] Andrew Zangwill. Hartree and Thomas: the forefathers of density functional theory. Archive for History of Exact Sciences, 67(3):331–348, May 2013. CODEN AH-ESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/article/10.1007/s00407-
013-0114-4;   https://en.wikipedia.org/wiki/Douglas_
Hartree;   https://en.wikipedia.org/wiki/Llewellyn_Thomas.

Zangwill:2014:EWK

[Zan14] Andrew Zangwill. The education of Walter Kohn and the creation of density functional theory. Archive for History of Exact Sciences, 68(6):775–848, November 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-014-0140-x.

Zehe:1983:GNF

[Zeh83] Horst Zehe. Die Gravitationstheorie des Nicolas Fatio de Duillier. (German) [The gravitational theory of Nicolas Fatio de Duillier]. Archive for History of Exact Sciences, 28(1):1–23, March 1983. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
28&issue=1&spage=1.

Zeldes:2009:GRT

[Zel09] Nissan Zeldes. Giulio Racah and theoretical physics in Jerusalem. Archive for History of Exact Sciences, 63(3):289– 323, 2009. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Zik:2012:MHT

[ZH12] Yaakov Zik and Giora Hon. Magnification: how to turn a spyglass into an astronomical telescope. Archive for History of Exact Sciences, 66(4):439–464, July 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0003-9519&volume=66&issue=4&spage=439.

Zik:2017:HSS

[ZH17] Yaakov Zik and Giora Hon. History of science and science combined: solving a historical problem in optics — the case of Galileo and his telescope. Archive for History of Exact Sciences, 71(4):337–344, July 2017. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Zik:2024:FFP

[ZH24] Yaakov Zik and Giora Hon. Francesco Fontana (1580– 1656) from practice to rules of calculation of lens systems. Archive for History of Exact Sciences, 78(2):153–182, March 2024. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657

(electronic). URL https://link.springer.com/article/10.
1007/s00407-023-00321-1.